A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7600–7649

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

2 [150, 520]. 3 [290, 233, 237, 458, 406, 29, 223, 242, 520, 28, 473]. 4 [228]. ^{3} [455]. k [89]. L_{1} [435]. t [289]. Θ(n^{2}) [450].

-Approximation [150]. -Domatic [89]. -Normal [289].

1/WG [674]. 10th [698]. 11th [684]. 13th [699]. 14-3-3 [117]. 14th [697, 696, 683]. 160 [472]. 17th [682]. 18th [673]. 19th [675].

2012 [191, 198]. 20th [687]. 25th [701]. 26th [677]. 2nd [672].

31st [678].

4-Bit [465]. 4G [161]. 4th [686, 681, 688, 671].

5th [676].

6th [685, 702].

7th [674, 693, 694].

8th [680].

9th [689].

= [403].

Abdominal [21, 671]. Abnormal [281]. Abnormalities [489]. Abortable [155, 154]. ABS [551]. Abstract [74, 132, 120]. Abstracting [531]. Abstraction [444, 566]. Abstractions [454]. Accents [213]. Access [144, 638, 637, 39]. Accessing [636]. Accessory [401]. Account [628]. Accountable [609]. Accounting [124]. ACE [689]. Acquisition [226, 62, 18]. Action [305, 326]. Actions [442]. Activation [319]. Active [431, 105]. Actor [555]. Adaptation [421, 35, 576]. Adaptative [40]. Adapting [35]. Adaptive [318, 309, 14, 486, 537]. Add [245]. Add-Ons [245]. Additional [394]. Address [661]. Addressing [653]. Adjusting [136]. Administration [36, 642]. Adults [344]. Advances [693, 684, 680, 689, 699]. Advancing [638]. Advantage [473]. Adversarie [612]. Adversary [467]. Advertising [403]. Advice [248]. AdvoCATE [180]. Advocatus [58]. Aerial [513]. Aerospace [545, 531]. AES [466]. Affect [527]. Affective [508]. Affinity [597]. Again [149]. against [621, 649, 467]. Agent [642]. Agents [139, 640, 701]. Aggregated [667]. Aggregation [66]. Agile [373]. Agreement [624]. AI [699, 629, 701]. AICOL [701]. AICOL-III [701]. Aided [19, 18, 15]. Aircraft [37]. Airstic [393]. Airworthiness [159]. AKAP [410]. Alberti [229]. Algebraic [451]. Algorithm [491, 293, 141, 150, 511, 591, 433, 483, 592, 484, 271, 552, 500]. Algorithmic [89]. Algorithms [465, 604, 57, 143, 87]. Alignment [485, 211]. Allowing [53]. Alloy [39]. Almost [109]. Alphabets [86]. Always [355]. am [701]. Ambiguous [274]. American [699]. among [344, 577]. Amorphous [73]. Amsterdam [674]. Analogical [312, 205]. Analysis [632, 410, 183, 289, 290, 413, 360, 361, 45, 219, 236, 166, 377, 610, 684, 29, 654, 668, 131, 554, 394, 614, 496, 583, 129, 650, 375, 161, 284, 285, 493, 633, 118, 532, 193]. Analyzer [218]. Anatomical [29]. Ancient [226, 238]. Android [463, 331, 281]. Annihilators [474]. Annotated [107, 7]. Annotation [643]. Anomalous [301]. Anonymity [661]. Anonymous [659]. Answer [444]. Anthropomorphic [442]. Anthropomorphism [345]. Antipatterns [565]. API [377]. Applaud [389]. Application [318, 132, 642, 428, 16, 122, 372, 18, 496, 494, 268, 544, 24, 386, 17, 464, 281]. Applications [457, 123, 246, 551, 34, 2, 376, 676, 385, 671, 170, 670].

Applied [293, 485]. Applying [621, 546, 162]. Approach [383, 183, 171, 108, 643, 604, 45, 41, 581, 25, 506, 30, 21, 660, 369, 534, 213, 32, 35, 181, 489, 525, 576, 185, 255, 193]. Approachable [327]. Approaches [231, 352, 353, 2, 701]. Approximate [94, 436]. Approximation [411, 150, 148, 595]. Approximative [413]. April [687]. Arabic [218]. Arc [107]. Arc-Annotated [107]. Archaeological [235, 234, 240, 223, 225]. Archaeology [234]. Architectural [242, 35]. Architecture [280, 567, 456, 537]. Architectures [229]. Areas [513]. Argumentation [701, 55, 641]. Art [19, 3]. Article [8]. Articles [108]. Articulated [450]. Artifact [566]. Artifact-Centric [566]. Artifacts [502]. Artificial [408, 699]. ASCoMS [679, 191]. Asia [696]. Asia-Pacific [696]. ASMONIA [161]. Assemble [491]. Assembly [291, 157, 417]. Assess [339]. Assessing [382, 189, 256, 577, 384]. Assessment [159, 617, 492, 620, 176, 174, 185]. Assistant [339, 341]. Assisted [406, 492]. Assistive [338]. Association [593]. Associations [491, 300]. Associative [254]. Assumption [656, 505]. Assurance [180, 581, 181, 185]. Astronomy [239]. ATL [559]. AToM [455]. Atomic [548]. Attack [612, 613, 475, 649, 268]. Attack-Defense [613]. Attacker [614]. Attacker-Defender [614]. Attacks [257, 267, 474]. Attending [407]. Attestation [651]. Attitudes [304]. Attribute [435, 359, 358, 278, 276]. Attribute-Based [278, 276]. Attributes [70, 580]. Audio [262]. Audit [609]. Augmentation [391]. Augmented [240, 128]. August [701]. Authenticated [467]. Authentication [359, 662, 358, 199, 664, 663]. author [532]. Autism [332]. Autistic [333]. Auto [33]. Automata [75, 84, 91, 176]. Automated [246, 228, 650, 344, 227]. Automatic [524, 400, 57, 223, 623, 560, 230, 213, 27]. Automaticity [327]. Automation [168, 180, 587]. Automotive [132, 167, 163, 196, 128]. Automotive/Infotainment [132]. Autonomous [195, 692, 446]. Available [255]. Average [78]. Average-Case [78]. Avoidance [283]. Aware [139, 68, 551, 572, 576]. Awareness [653, 269]. Axial [436]. Axiom [59]. Axioms [636].

B [561]. Babel [368]. Back [9, 50, 260, 324, 419, 479, 625, 647]. Backbone [143, 256]. Background [655]. Bacterial [493]. Balancing [632]. Ballots [665]. Bangalore [531]. Banking [244]. Base [323, 210]. Baseband [651]. Based [280, 42, 410, 439, 628, 309, 5, 482, 383, 566, 183, 644, 642, 643, 293, 112, 619, 431, 45, 379, 262, 234, 515, 33, 134, 428, 279, 39, 433, 662, 8, 158, 392, 581, 463, 223, 25, 164, 483, 60, 208, 437, 665, 278, 520, 21, 660, 37, 492, 122, 649, 516, 213, 583, 528, 53, 63, 209, 181, 210, 205, 657, 526, 415, 446, 518, 525, 473, 573, 664, 43, 464, 572, 276, 595, 281, 184]. Baselines [605]. Bases [59, 53]. Basis [504]. Bathcratch [392]. Bathtub [392]. Battista [229]. Bayesian [311]. Be [83, 374]. Behavior [258, 270, 321, 327, 448, 446]. Behavioral [344]. Behaviors [333]. Behaviour [449, 527, 332]. Behind [340]. Beijing [695]. Benchmark [632, 458]. Benefits [95]. Better [587, 286]. between [431, 173, 430, 432]. Beyond [70, 422]. Bibliographic [530]. Binary [582, 306]. Binding [485]. Binomial [301]. Biochemical [413]. Biochemistry [415]. Bioinformatics [490, 694]. Biological [412, 297, 597]. Biology [140, 690]. Biomarker [500]. Biomedical [100]. biometric [199]. Bioscientific [116]. Bipartite [148]. Bird [659]. Bisimulation [81, 90]. Bisimulations [79]. Bit [465]. Bits [607]. Black [172, 610]. Black-Box [172]. Block [475]. Blocking [615]. Blocks [548]. Blueprint [32]. BM25MF [70]. Board [396, 174]. Boardroom [665]. Bond [497]. Bond-Orientational [497]. Bone [412]. Bonolo [528]. Boomerang [472]. Botnet [270]. Boundary [640, 23]. Bounded [142]. Bounded-Contention [142]. Bounds [612, 140, 151, 147]. Box [172, 229]. Brain [489]. Branching [90]. Brazil [677]. Brazilian [644, 492]. Breaches [78]. Breaking [318, 152]. Bringing [347]. Broadcast [153]. Browsing [534]. Brushing [391]. BSG [128]. BSG-E [128]. Bubble [106]. Budapest [700]. Build [635]. Building [215]. Buried [235]. Bus [296]. Business [381, 578]. Byzantine [153].

C [458]. Caches [110]. CafeOBJ [129]. CalcWcet167 [131]. Call [568]. Cambridge [687]. Camellia [475]. camera [224]. Can [258, 527, 83, 332, 335, 374]. Cancer [312, 498, 30, 599, 27]. Caption [407]. Carbohydrates [485]. Care [336, 162]. Carlo [587]. Cars [195]. Cartagena [675, 699]. Cascading [495]. Case [412, 644, 78, 180, 157, 676, 181]. Case-Based [181]. Cases [633, 185]. Casino [355, 354]. Castle [234]. Categories [377]. Cathedral [242]. CBTree [136]. CCL [448]. Cell [690]. Centered [670, 688]. Centred [369]. Centric [566, 182, 534]. Cervical [599]. Chain [184]. Challenge [124, 645]. Challenges [195, 360, 361, 701, 630]. Chandy [552]. Change [569]. Changing [304]. Channel [144, 152, 83, 516]. Character [208]. Character-Based [208]. Characterisation [86]. Checking [449, 552, 583, 416]. Cheer [389]. Chemometrics [492]. Chengdu [686]. Children [332, 335, 333]. China [683, 686, 695, 698, 702]. Chinese [535, 203]. Chinese-Style [535]. Choices [371]. Chopstick [341]. Chopstick-Equipped [341]. Choreographies [581]. Choreography [579]. Chosen [656]. Cinderella [77]. Cine [22]. Cine-MRI [22]. Cipher [471, 475, 473]. Ciphertext [656, 468, 470]. Citation [531]. Citations [529]. City [228, 673]. Classical [318, 321]. Classification [611, 299, 598, 533, 603, 519, 41]. Classifier [598]. Classifiers [294, 427, 19]. Cleaning [13]. Cleansing [65]. Clinical [12, 671]. Clothed [507]. Cloud [631, 224, 551, 582, 624, 275, 32, 259, 571, 573, 572]. Clouds [32, 513]. Cluster [591, 592]. Clustering [97, 293, 509, 604, 426, 593, 214, 591, 592, 437, 660, 302, 600, 574, 487, 118]. Clustering-Based [437, 660]. Co [532, 486]. Co-author [532]. Co-identification [486]. Coalgebras [79]. Code [47]. Codes [418]. Coding [142, 33]. Cognitive [589, 327, 332]. Coherence [628]. Coherence-Based [628]. Coherent [310]. Coinductive [252]. coli [309]. Collaboration [335].

Collaborative [202, 140, 539, 397]. COLLADA [399]. Collected [256]. Collecting [139, 334, 385]. Collection [102, 224]. Collections [528]. Collective [550, 639]. Colombia [675, 699]. Colon [13]. Colonography [11, 16, 19, 18, 14, 15, 12]. Colonoscopy [23]. Color [505]. Colorization [508]. Combination [1]. Combinations [639]. Combinatorial [266]. Commercial [52]. Commitment [478]. Common [107, 432]. Commons [645]. Communication [168, 167, 207, 169]. Communications [683]. Community [481]. Compact [602, 468]. Comparable [206]. Comparative [19, 250]. Compared [319]. Comparing [2]. Comparison [459, 336]. Compensation [22]. Competitions [247]. Compilation [74]. Complementary [561]. Complete [521]. Complex [412, 449, 461, 155, 298, 415, 306, 481]. Complexity [429, 430, 84, 89, 530, 629, 701, 384]. Compliance [627, 184]. Component [567, 493]. Components [163]. Composable [549]. Composed [575]. Composing [456]. Composite [11]. Composition [381, 583, 572]. Compositional [166, 423, 370]. Compressed [98, 103, 112]. Compression [293, 262, 502, 472]. Compromised [620]. Computability [73]. Computation [104, 428, 34, 145]. Computational [690, 671, 695]. Computations [477]. Compute [568]. Computer [674, 134, 693, 406, 19, 18, 15, 689, 618, 446, 678, 679]. Computer-Aided [19, 18, 15]. Computers [631]. Computing [677, 101, 611, 430, 655, 698, 582, 275, 73]. Computting [631]. Concept [292, 380]. Concepts [339]. Conceptual [314, 36, 60, 469, 3, 542]. Concerns [45]. Concurrency [549, 547]. Concurrent [136, 449, 476, 548]. Conditional [499, 24]. Conference [672, 697, 674, 696, 683, 686, 700, 695, 681, 680, 529, 682, 698, 689, 692, 678, 699, 694, 688, 702, 673]. Confidence [185]. Confidentiality [613]. Configurations [99, 328]. Configuring [32]. Confinement [558]. Conflict [381]. Conflicting [371]. Conformance [463]. Congress [701]. Conjunction [671]. Connected [244, 153]. Connecting [322]. Consensus [99]. Conservation [231, 236]. Consideration [504]. Considerations [123]. Consistent [628, 138, 69]. Constant [277]. Constant-Round [277]. Constrained [302, 580, 514, 467, 465]. Constraining [375]. Constraint [527, 581, 126]. Constraint-Based [581]. Constraints [6, 58, 130, 563, 125, 196, 127, 562, 418]. Construction [478]. Consumer [646]. Consuming [398]. Contact [450, 334]. Content [394, 532]. Contention [142, 151]. Context [318, 80, 209, 387, 397, 378, 43, 576]. Context-Aware [576]. Context-Based [43]. Context-Free [80]. Contexts [445, 296, 301]. Contextual [315]. Continual [272]. Continuous [291, 453, 436]. Continuum [454]. Contracts [629]. Contractual [624]. Contradiction [586]. Contrast [515, 27, 518, 24]. Contributors [34]. Control [549, 33, 134, 39, 158, 623, 366, 251, 454, 578, 650, 446]. Control-Flow [251]. Controller [175, 392]. Controllers [456]. Conversion [273, 216]. Cooperative [194]. Coopetitive [280]. Coordinated [575]. core [188].

Corpora [206, 211]. Correct [560]. Correction [94, 524, 297]. Correlation [435]. Corresponding [258, 661]. Cost [619, 223, 227]. Cost-Based [619]. Costume [535]. Countermeasures [268]. Course [289]. Court [633]. Courts [632]. Cover [148]. CP [584]. CP-Nets [584]. Cracking [254]. Creating [375]. Creative [645]. Creativity [386]. Credentials [255]. Crete [676]. Crimes [629]. Criminal [641]. Criminal Process [644]. Critical [653, 172, 163, 372, 466, 227, 200, 162]. Criticality [615]. Crohn [21]. Cross [166, 221, 329]. Cross-Cultural [329]. Cross-Language [221]. Cross-Level [166]. Crosscutting [45]. Crossing [107, 425]. Crossing-Free [107]. Crowds [54]. Crowdsourcing [202, 535]. Cryptanalysis [265, 471, 473, 266]. Cryptography [263, 608, 464]. Cryptosystem [266]. CT [20, 11, 16, 19, 18, 14, 15, 12, 27, 17]. CT-Scans [27]. CTA [28]. CTC [17]. Cultural [229, 238, 681, 387, 329, 239]. Culture [329]. Curve [297]. Curves [464]. Customization [402]. Customized [503]. CVM [695]. Cyber [248, 247]. Cybercriminal [610]. Cybernics [443]. Cyprus [681].

D [290, 233, 237, 458, 406, 29, 228, 223, 242, 520, 28, 499, 473, 486]. D-Based [473]. D-Line-Drawing-Based [520]. Dada [293]. DAnIEL [208]. Dao [523]. Dao-Fa [523]. Daoism [523]. Data [97, 672, 286, 5, 95, 289, 609, 291, 7, 262, 593, 670, 68, 684, 167, 668, 221, 297, 116, 660, 301, 492, 667, 230, 259, 594, 596, 284, 285, 597, 49, 603, 486, 56, 121]. Data- [670]. Data-Compression [262]. Data-Discovery [7]. Database [7, 234, 670]. Database- [670]. Datasets [65]. DC [665]. DC-Net [665]. DDH [656]. Dealing [292]. Debugging [253]. Deceptive [616]. Decidability [650]. Decision [311, 44, 582, 599, 641, 216, 700]. Decisions [317]. Decomposition [163, 437, 90, 504, 505]. Decompression [512]. Deep [220, 434]. Defect [231]. Defender [614]. Defending [621]. Defense [613]. Defenses [619]. Defined [20]. Definition [209, 370]. Deformable [119]. Delivery [570]. Dementia [378]. Demonstrations [239]. Denial [268]. Denial-of-Service [268]. Denoising [511, 533]. Dense [450, 146, 513]. Dependable [192]. Dependence [553]. Dependency [43]. Dependent [318, 329]. Deploy [457]. Deploying [460]. Deposited [600]. Derivation [579]. Description [399, 128]. DESEC4LCCI [679, 198]. Design [383, 399, 141, 457, 345, 236, 371, 624, 375, 574, 117, 340, 368, 575]. Designated [658]. Designed [244]. Designers [384]. Designing [401, 404, 397]. Desktop [239]. Details [511]. Detect [21, 603]. Detecting [667, 599, 489]. Detection [94, 65, 291, 515, 270, 19, 30, 649, 565, 18, 14, 15, 23, 586, 232, 518, 666, 281, 481]. Determining [257]. Deterministic [84, 85]. Developing [644, 561]. Development [132, 382, 379, 38, 163, 373, 372, 248, 387, 397, 174]. Deviations [296]. Devices [237]. DHT [5]. DHT-Based [5]. Diaboli [58]. Diagnose [498]. Diagnosis [642]. Diagnostics [44]. Diagram [45]. Diagrams [582]. Dialects [418]. Diamond [600]. Did [418]. Difference [265]. Different [425]. Differential [267, 451, 601]. Differential-Algebraic [451]. Difficulty [394]. Diffusion [20, 28]. Digital [535, 696, 523, 539, 230, 528, 502, 537]. Digitally [328]. Digitally-Extracted [328]. Dimensional [498]. Dimensionality [295]. Dimensions [55]. Direct [466]. Directed [106]. Direction [608]. Directly [656]. Disabled [342]. Disambiguation [668]. DISC [677]. Discontinuity [25]. Discourse [63]. Discovery [383, 7, 312, 52, 56]. Discrete [305, 434, 438, 474]. Discretization [595]. Discriminant [495]. Discrimination [27]. Discriminative [295]. Discussers [105]. Discussion [349, 363, 355, 357, 353, 347, 361, 359, 351, 365]. Disease [21]. Dissemination [637, 147]. Dissertations [529]. Dissipation [310]. Distance [425, 435, 327]. Distances [430]. Distant [108]. Distinguishers [472]. Distributed [172, 144, 150, 149, 143, 87, 552, 145, 477, 677]. Distribution [622, 296]. Distributions [173]. Divergence [429]. DJ [392]. DL [120]. Do [568, 96]. Docking [496]. Doctrine [628]. Document [635, 109, 604, 96, 108]. Documentation [231, 384]. Documents [6, 57, 60]. Domain [352, 353, 457, 646, 273, 122, 120, 532]. Domain-Models [120]. Domain-Specific [352, 353, 457, 532]. Domains [653]. Domatic [89]. Dorothy [337]. Dose [16, 15]. Dots [322]. Double [502]. Drawing [520]. Drift [292]. Driven [75, 132, 381, 134, 32]. Driver [115]. Drug [589, 52]. Drum [393]. Drums [393]. Drumstick [393]. DSLs [38]. DSpace [540, 526]. Dual [20, 25, 14]. Dual-Energy [20, 14]. Duality [322]. due [235]. Duplication [251]. Duration [296]. during [536]. Dutch [633]. Dynamic [280, 567, 450, 147, 578, 580, 146, 145, 488, 272, 576]. Dynamical [413]. Dynamics [318, 309].

E-Indexing [531]. E-Voting [360, 361]. Early [375]. Economic [572]. Ecosystem [418]. Ecosystems [569]. Edit [425, 435]. editing [202]. Edwards [464]. Effect [307, 257]. Effective [70]. Effectiveness [619]. Efficient [103, 104, 106, 152, 137, 662, 476, 615, 119, 154, 13, 510, 477]. Effort [369]. EFSMs [170]. Eigenimages [438]. EKAW [673]. Elder [336]. Eldercare [340]. Electricity [622]. Electronic [350, 351]. Elements [161, 64]. Elite [362, 363]. Ellsberg [308]. Embedded [33, 166, 560, 196]. Embedding [435]. Embodied [442]. Embodiment [332]. Emergence [313]. Emergent [315, 408]. Emotional [336, 328]. Emotions [331]. Empathic [401]. Empirical [249]. Enabling [7, 576]. Encryption [470, 278, 469, 468]. Encyclopedia [210]. End [379, 34, 53, 580]. End-to-End [580]. End-Users [34]. Ending [586]. Endosymbiosis [418]. Enemy [621]. Energy [20, 322, 236, 14, 117]. Enforcing [624, 350, 351]. Engine [110, 175, 228]. Engineering [672, 697, 195, 529, 182, 583, 368, 186, 386, 688, 673]. Engineers [531]. English [211]. Enhanced [200, 27]. Enhancement [279, 28, 516]. Enhancing [263, 294, 649, 394, 596]. Enjoyment [538]. Enrichment [59, 58, 60]. Ensemble [437]. Entangled [315]. Entanglement [314, 320, 319]. Enterprise [574, 48]. Entertainment [689]. Entities [314, 315]. Entity [102, 70, 62]. Entrain [418]. Enumeration [106]. Environment [379, 123, 406, 460, 659]. Environments [404, 192]. Equation [504]. Equilibria [611]. Equipment [162]. Equipped [341]. ERCIM [679]. ERCIM/EWICS [679]. Errors [175]. Escherichia [309]. Estimating [661]. Estimation [311, 591, 507, 422, 459, 369, 605, 421]. Ethical [701]. Eunomos [635]. EuroMed [681]. Evaluated [426]. Evaluating [619, 2, 187, 540]. Evaluation [100, 458, 376, 343, 28, 271, 605, 199, 447, 341, 663]. Event [579, 132, 561]. Event-B [561]. Event-Driven [132]. Eventually [355]. Everyday [444]. Everything [257]. Evidence [184]. Evolution [8, 408, 49, 418, 601]. Evolutionary [293, 604, 428, 484, 184]. EWICS [679]. Examination [249]. Example [167]. Exchange [274, 467]. Exciting [596]. Exclusion [154]. Executable [122]. Execution [132]. Executions [92]. Exemplar [26]. Exemplified [158]. Exercises [247]. Exergame [403]. Exhibiting [371]. Experience [404, 373]. Experiences [36]. Experiments [297, 247, 113]. Expert [646, 670, 369]. Expert-Centred [369]. Expert-Systems [670]. Expertise [373]. Experts [54]. Explaining [296]. Exploitation [232]. Exploiting [210, 526]. Exploratory [58, 541]. Exploring [588, 233, 530, 241, 387]. Exponent [101]. Expression [486]. Expressions [86, 328, 329]. Extended [602]. Extending [574]. Extensible [399, 490, 456]. Extension [586, 584]. Extensions [567]. External [12]. Extract [485]. Extracted [328]. Extracting [61, 170]. Extraction [206, 218, 221, 63, 209, 210, 48, 434, 517]. Eye [659].

Fa [523]. Facial [328, 329]. Facilitate [335]. Factorization [103]. Factors [368]. Fail [355, 244]. Failure [172, 173]. Fair [274]. Fast [141, 450, 145, 67, 474]. Faster [286]. Fault [165, 44, 187, 537, 193]. Fault-Propagation [193]. Fault-Tolerant [537]. Faults [173, 558, 192]. Favourite [389]. FDIR [176]. Feature [575, 521, 517]. Features [18, 287]. Fecal [17]. Federated [102, 32]. Feedback [113]. Feel [339]. Fetch [137]. Fetch-and-Increment [137]. Few [300]. Field [323]. Fieldbus [164, 169]. Fields [499]. File [399, 528, 659]. File-Based [528]. File-Sharing [659]. Films [600]. Filter [649]. Filtering [269]. Filters [28]. Finders [559]. Finding [149, 296]. Findings [336]. Finds [240, 223]. Finite [566, 449]. Finite-State [449]. Finland [684]. Firewall [269]. First [557, 80, 695]. First-Order [557, 80]. Fish [368]. Fitness [604]. Fittest [633]. Fitting [283, 297, 217]. Flashback [405]. Flattening [13]. Flexible [452, 126]. Flipbook [390]. FlipIt [621]. Flow [30, 366, 554, 251]. FMEA [171]. Follow [545]. Food [403]. Foods [341]. Forensic [167]. Forensics [502]. Forest [482]. Form [74, 411]. Formal [697, 545, 38, 133, 130, 550, 87, 676, 129, 47, 174, 544, 170, 697]. Formalizing [298]. Forward [467, 276]. Foundation [290]. Fourier [474]. FPGA [271]. Fractals [310]. Fragments [68]. Frame [644, 64, 644]. Frame-Based [644]. FrameNet [643]. FrameNet-Based [643]. Frames [63]. Framework [202, 439, 634, 503, 428, 617, 490, 187, 484, 182, 583, 460, 486]. France [672, 685, 688, 671]. Frankfurt [701]. FRBR [542]. Free [101, 107, 80, 85]. French [213]. Frequent [298]. Freshening [67]. Friction [451]. Friendly [259]. Front [10, 31, 480, 501, 522, 543, 564, 585, 606, 626, 648, 669, 51, 72, 93, 114, 135, 156, 177, 178, 190, 197, 201, 222, 243, 261, 282, 303, 325, 346, 367, 388, 409, 420, 441, 462]. Fujian [702]. Fukuoka [693]. Full [33, 472]. Fun [389, 403]. Function [604, 615, 472, 473]. Functional [289, 194, 415]. Functions [482]. Functors [79]. Fus [523]. Fusion [230, 443, 601]. Fuzzy [589, 603, 525].

Gait [408]. Gallbladder [29]. Galway [673]. Game [389, 700, 617, 610, 406, 615, 506, 614, 538, 396, 618, 397]. Game-Theoretic [617, 614]. Game-Theoretical [506]. Games [611, 623, 613, 620]. GameSec [700]. Gaming [587]. Gates [165, 466]. GE [228]. Gene [491, 484, 598]. Gene-Protein-Reaction [491]. General [528]. Generalized [141, 450, 428, 127]. Generating [400]. Generation [535, 57, 228, 164, 623, 592, 560, 329]. Generative [495]. Generic [478]. Genetic [604, 57, 483, 500, 284, 418]. Genome [104, 491]. Genome-Scale [491]. Genomics [289]. Geo [233]. Geo-Referred [233]. Geometric [430, 422, 442]. Germany [678, 679, 701, 242, 336]. Gesture [370]. GF [217]. Global [504, 397, 518, 505]. Globalized [696]. Go [587]. Goal [583]. Goal-Oriented [583]. Good [373]. Governance [362, 363, 609]. GPU [512]. Gradient [510]. Grammar [103, 112, 217]. Grammar-Based [112]. Grammarware [6]. Grand [645]. Granny [338]. Graph [429, 431, 425, 435, 394, 53, 432, 521, 113]. Graph-Based [53]. Graphs [106, 424, 69, 89, 553, 469, 48]. Grassmann [487]. Gravitational [591, 592]. Greece [676]. Grid [653, 623]. Ground [372]. Group [270, 398]. GSM [566]. GSM-Based [566]. GUC [478]. GUI [374]. Guidance [69]. Guide [386]. Guided [160]. Guidelines [371]. Guppy [307].

H.264 [663]. H.264/SVC [663]. Hamburg [228]. Handling [341]. Handoff [662]. Hard [127]. Hardware [173, 271]. Harnessing [54]. HAS-160 [472]. Hash [271, 473]. Having [389]. Hazard [160]. HCSE [688]. Healthcare [124, 123, 343, 122, 121]. Hearing [407]. Heart [334]. Held [701, 671]. Help [332]. Helsinki [684]. Her [338]. Heraklion [676]. Heritage [237, 238, 681, 230, 241]. Heterogeneous [97, 48, 56]. HFCVD [600]. Hide [607]. Hierarchical [439, 311, 433, 423, 598]. High [334, 142, 511, 297, 498, 118, 486]. High-D [486]. High-Resolution [511]. High-Throughput [118]. Highly [450, 255]. Hijacking [251]. Hindi [211]. Hiroshima [691]. Historic [228]. Historical [236, 667, 525]. HIV [589]. Hole [217]. Home [338]. Honesty [624]. Honeypot [618]. Hong [683]. Hosts [661]. Housing [236]. HRI [343]. Hui [523]. Hui-Yuan [523]. Human [445, 345, 527, 507, 34, 443, 284, 368, 688]. Human-Centered [688]. Human-Robot [345]. Humanoid [332, 329]. Hungary [700]. Hybrid [376, 561, 571, 67, 255]. Hydrophobic [496]. Hypergraphs [440]. Hyperproperties [252].

I2P [659]. IaaS [582]. IAPR [691, 694]. IBERAMIA [699]. Ibero [699]. ICA [521]. ICADL [696]. ICFEM [697]. ICICS [683]. ICSOC [698]. ICSR [686]. IDA [684]. Identification [235, 45, 279, 160, 486]. Identify [258]. Identifying [654, 105, 301, 502, 594, 340]. Identity [279, 662, 313]. Identity-Based [279, 662]. If [444]. IFIP [674]. II [676]. III [701]. Illumination [504]. Image [439, 509, 433, 507, 226, 223, 506, 540, 502, 227, 508, 505, 118]. Image-Based [223]. Imagery [232, 342]. Images [524, 20, 503, 230, 504, 489]. Imaging [671]. Imbalanced [603]. Impact [202, 110, 173, 264, 175, 373, 597]. Impaired [407]. Imperial [242]. Implementation [141, 2, 376, 466, 374]. Implemented [392]. Implications [140]. Implicit [266]. Improve [421, 211]. Improved [112, 267, 663]. Improvement [496]. Improving [62, 369, 488, 533, 603, 300]. Incentives [622]. Increment [137]. Incremental [6, 519]. Independent [208]. Indeterminacy [317]. Index [235, 20]. Indexer [533]. Indexes [112]. Indexing [531]. Indian [387]. Indias [675, 699]. Indication [641]. Indiscernability [323]. Individual [356, 357, 313]. Individuals [41]. Inductive [46]. Industrial [293, 44, 162]. Industrial-Strength [162]. Industry [545]. Inertia [450]. Inference [316, 76]. Infinite [86]. Influence [327, 577]. Influencers [105]. Influences [320]. Inform [340]. Informatics [637]. Information [100, 139, 74, 675, 638, 234, 425, 147, 424, 693, 221, 298, 366, 554, 40, 538, 209, 443, 536, 595, 683]. Infotainment [132]. Infrared [334]. Infrastructure [52, 200]. Infrastructures [249]. Initializations [288]. Injection [187]. Input [75]. Input-Driven [75]. Inputs [173]. Inspired [304]. Instability [313]. Installation [403]. Installations [404]. Institutes [638]. Integer [273]. Integrating [38]. Integration [393, 453, 121]. Integrity [6, 161]. Intelligence [699, 630, 330]. Intelligent [133, 684, 284]. Intensity [11]. Interaction [445, 685, 345, 323, 496, 536, 398, 481]. Interactions [401, 344]. Interactive [390, 404, 405, 69, 366, 372]. Interesting [296]. Interestingness [298]. Interface [402, 381, 91, 371, 387]. Interfaces [35, 556]. Interference [307, 189]. Interfering [315]. Interlinking [65]. International [672, 677, 697, 674, 685, 675, 696, 683, 687, 686, 691, 700, 693, 684, 695, 681, 680, 698, 676, 689, 692, 678, 701, 694, 688, 702, 671, 673]. Internet [348, 349, 503, 256]. Interpersonal [327]. Interpretation [628, 492]. Interpreting [589]. Intrinsic [505]. Introducing [496]. Introduction [191, 347, 198, 179]. Intruder [611]. Intrusion [649]. intrusive [502]. Intuitive [641]. Invariant [263]. Invasive [599]. Inversion [450]. Investigating [64]. Investigation [173]. Investigations [167]. IP [649]. IR [60]. Ireland [673]. Irrational [316]. Isoform [117]. Isoform-Selective [117]. ISoLA [676]. Isomorphisms [657]. Issue [320, 670, 690]. Issues [360, 361, 364, 365, 73]. IT-Forensic [167]. Italian [632]. Iterative [453]. IVR [701]. IWDE [679]. IWSEC [693].

J [499]. Jam [397]. Japan [697, 691, 693, 680, 692, 694]. Japanese [203, 217]. JapTAL [680]. Java [458]. Jensen [429, 440]. Jet [175]. Joint [691]. JPEG [502]. Judgment [634]. June [685]. jViz.RNA [488].

K-Means [288]. Kaiserdom [242]. Kanazawa [680]. Karlskrona [682]. Kathmandu [689]. Kernel [439, 440, 427, 654, 668]. Kernel-Based [439]. Kernels [425]. Key [470, 267, 664, 266, 467]. Key-Based [664]. Keyphrase [533]. Keyphrases [533]. Keys [65]. Keyword [275]. Kitchen [257]. KNN [426]. Knowledge [635, 59, 277, 646, 670, 52, 62, 53, 63, 574, 210, 56, 673, 673]. Knowledge-Centered [670]. Kong [683]. Königslutter [242]. KPIs [577]. KRK [478]. Kyoto [697]. Kyushu [529].

label [294, 299, 605, 519]. Labelling [220, 432]. Lamport [552]. Landmark [11, 25, 328]. Landmark-Based [25]. Landscapes [225]. Language [218, 644, 290, 457, 61, 406, 680, 221, 455, 208, 645, 701, 128]. Languages [352, 353, 85, 86, 374, 415]. Laplacian [438]. Large [36, 59, 233, 670, 373, 372, 540]. Large-Scale [670]. Laser [242, 230]. Latent [43]. Law [318, 631, 629]. Laymen [646]. LBlock [267]. Lead [312]. Leaders [105]. Leakage [272]. Leakage-Resilient [272]. Learn [332]. Learning [482, 382, 233, 258, 427, 406, 21, 492, 605, 421, 586, 332, 598, 397, 489, 48, 521, 519]. Least [305]. Legal [644, 635, 638, 643, 637, 645, 701, 630]. Legislation [636]. LeGo [294]. Leisure [541]. Lemma [218]. Lemonizing [610]. Leon [229]. Lesions [14, 27]. Less [327]. Let [338]. Level [166, 359, 358, 394, 205, 256, 268, 289]. Levels [47]. Leveraging [676]. Lexically [215]. Lexicon [644]. Libraries [696, 138]. Library [535, 539, 528, 537]. LiDAR [513]. Life [347, 408]. Lifecycle [544]. Ligand [117]. Light [401]. Lighting [401]. Lightweight [157, 471]. Like [309, 316, 108, 321]. Likeness [345]. Limassol [681]. Limitations [95]. Limited [373]. Line [4, 520, 233]. Linear [101, 80, 615, 498]. Linguistic [213, 63]. Link [615, 652]. Linked [68]. Linked-Data [68]. Links [256]. Linux [115]. List [649]. List-Based [649]. Literatures [532]. Liver [29, 28, 26, 24, 27]. Liver-Gallbladder [29]. Living [447]. Load [327]. Local [294, 652, 387, 505, 595]. Localizing [397]. Locks [155]. Logic [557, 76, 563, 88, 126, 466, 552, 562]. Logics [639]. Long [211]. Longest [107]. Loop [355, 354]. Loosely [5, 153]. Loss [615]. Lossless [262]. Low [223, 16, 15, 227, 268]. Low-Cost [223, 227]. Low-Rate [268]. Lower [140, 147]. LTS [82]. LZ78 [103].

Machina [308]. Machine [482, 203, 204, 44, 498, 492, 494, 205, 443, 489, 212]. Machines [449]. Magdeburg [678, 679]. Main [701]. Make [339, 328, 645]. Making [641]. malleability [470]. Malware [666]. Management [5, 609, 635, 381, 151, 189, 578, 673, 537]. Managing [36]. Manga [534]. Manifold [487]. Manipulating [326]. Manipulation [444]. Manipulative [258]. Manuscript [524]. Manuscripts [226]. Many [312]. Mapping [231, 646, 344, 205, 514]. Mappings [295]. MapReduce [278]. Maps [380, 589, 600]. Margin [124]. Markets [610]. Markov [411]. Markup [374]. Marts [49]. MAS [639]. Masked [118]. Mass [322]. Massive [503]. Matching [431, 30, 525, 150]. Matrix [111, 450]. Matter [9, 50, 260, 324, 419, 479, 625, 647, 10, 31, 480, 501, 522, 543, 564, 585, 606, 626, 648, 669, 51, 72, 93, 114, 135, 156, 177, 178, 190, 197, 201, 222, 243, 261, 282, 303, 325, 346, 367, 388, 409, 420, 441, 462]. Maximal [101, 104]. Maximal-Exponent [101]. Maximize [207]. May [340]. MCU [465]. MCut [299]. MDE [37]. MDE-Based [37]. Meal [341]. Mean [426, 428, 323, 562]. Mean-Payoff [562]. Meaning [322]. Means [288]. Measure [605, 516]. Measuring [615]. Mechanism [619, 649]. Mechanisms [609, 78]. MEDI [672]. Media [695, 55]. Median [428, 432]. Mediation [646]. Medical [162]. Medieval [234]. Medoids [287]. Meet [475]. Meet-in-the-Middle [475]. Meeting [188]. Meets [125]. Memory [140]. Mental [326]. Merging [588]. Mesh [517, 512]. Meshes [511]. Meta [38, 186]. Meta-model [186]. Meta-modeling [38]. Metabolic [491, 418]. Metabolism [309]. Metabolomics [492]. Metadata [535, 57, 534]. Metadata-Centric [534]. Metastasis [30]. Method [11, 158, 339, 160, 655, 250, 129, 205, 577, 590, 447]. Methodological [290, 368]. Methodologies [561]. Methodology [71, 157, 550]. Methods [697, 36, 545, 133, 377, 676, 544, 481]. Metric [287]. Metrics [356, 357, 270, 194]. MICCAI [671]. Micro [231]. Micro-morphology [231]. Microbloggers [105]. Microblogging [105]. Middle [475]. Military [348, 349]. Minimal [414, 14]. Minimal-Preparation [14]. Minimization [422, 623, 117]. Minimum [591]. Mining [286, 292, 500, 597]. Minkowski [288, 287]. Minority [99]. Mission [172]. Mitigation [617, 160]. Mixed [549]. Mixture [301]. Mob [362, 363]. Mobile [245, 280, 139, 389, 379, 237, 301, 538, 536, 651, 663]. Mobility [219]. Modal [91, 639, 494]. Mode [549, 173, 426]. Model [314, 308, 315, 410, 309, 183, 289, 642, 311, 559, 70, 491, 638, 283, 382, 449, 33, 134, 636, 463, 143, 228, 242, 164, 376, 615, 652, 613, 552, 496, 583, 32, 618, 47, 370, 478, 272, 451, 43, 548, 467, 416, 542, 184, 186, 672]. Model-Based [642, 33, 463, 184]. Model-Checking [449]. Model-Driven [134, 32]. Modeling [315, 458, 551, 34, 455, 116, 448, 122, 182, 692, 128, 639, 555, 573, 513, 38]. Modelling [352, 353, 634, 319, 589, 223, 304, 196, 374, 3, 620]. Models [42, 411, 559, 319, 453, 37, 301, 417, 701, 690, 35, 26, 495, 24, 575, 120, 517, 572]. Modifying [189]. Modular [82, 40, 556]. Modulating [391]. Monadic [76]. Money [545]. Monitor [579]. Monitoring [344]. Monte [587]. Monte-Carlo [587]. Monuments [232]. morphology [231]. MORSE [461]. Motion [290, 456, 23, 26, 22]. Motivation [446]. Motivation-Based [446].

Movement [341]. Moving [619, 660]. MR [21]. MRI [489, 22]. Multi [265, 289, 642, 108, 70, 152, 41, 294, 224, 188, 470, 490, 88, 299, 613, 605, 516, 199, 578, 494, 517, 519]. Multi-Agent [642]. Multi-biometric [199]. Multi-camera [224]. Multi-Channel [152, 516]. Multi-ciphertext [470]. Multi-core [188]. Multi-document [108]. Multi-label [294, 299, 605, 519]. Multi-level [289]. Multi-modal [494]. Multi-Prime [265]. Multi-scale [517]. Multi-stage [490]. Multi-step [613]. Multi-tenanted [578]. Multi-threaded [88]. Multi-valued [70]. Multi-viewpoints [41]. Multicore [187]. Multidimensional [5]. Multilingual [61, 210]. Multimedia [229]. Multimodal [206]. Multiobjective [484, 117]. Multiple [412, 144, 6, 257, 485, 418]. Multiplication [464]. Multispectral [226]. Multistate [451]. Mutex [155]. Mutual [97, 154].

Naked [507]. Named [62]. Names [525]. Nanny [337]. Nash [611]. Natural [61, 680]. Natural-Language [61]. Navigation [541]. Needs [381, 646, 539, 538, 340, 368]. Negative [58]. Negotiation [582]. Neighborhood [434]. Nepal [689]. Net [665]. Netherlands [674]. Nets [584]. Network [696, 509, 617, 615, 60, 649, 161, 633, 532, 702]. Networked [630, 666, 121]. Networks [622, 411, 142, 616, 152, 319, 151, 147, 662, 105, 143, 652, 153, 618, 146, 145, 663, 601, 481]. Neural [291, 601]. Never [586]. Never-Ending [586]. News [208]. NFC [280]. Nice [671]. Nichesourcing [54]. NLP [680]. NMR [492]. NMR-Based [492]. No [109, 148, 138, 155]. Nodes [431, 455]. Noise [20, 603]. Noisy [275]. Nominal [86]. Non [334, 470, 502, 639, 24, 73, 663]. Non-computability [73]. Non-contact [334]. Non-contrast [24]. Non-intrusive [502]. Non-malleability [470]. Non-normal [639]. Non-stationary [663]. Nonlinear [578]. Nordic [682]. NordSec [682]. Normal [289, 639, 27]. Normative [636, 627, 639]. Novel [503, 489]. Novelty [100]. November [697, 696, 691, 700, 693, 695, 681, 682, 698, 689, 692, 699, 694, 702]. NSS [702]. Nuclear [158]. Number [661]. Numbered [273]. Numbers [251, 546].

Object [509, 515, 423, 520, 660]. Objects [452, 654, 223, 640]. Oblivious [476]. Observation [433]. Observational [7]. Obtained [604]. October [672, 677, 675, 683, 686, 684, 681, 680, 682, 676, 688, 671, 673]. ODM [42]. ODM-Based [42]. Offs [371]. Offshore [160]. Offshoring [387]. Older [339, 344]. Olfactory [401]. On-board [174]. On-line [233]. One [467]. One-Round [467]. Online [172, 144, 244, 292, 592, 541]. Ons [245]. Ontological [171, 41, 40]. Ontologies [646, 58, 701]. Ontology [71, 634, 60, 586, 598, 48, 590, 525]. Ontology-Based [525]. OO [556]. OODA [355, 354]. OOV [209]. Open [638, 81, 637, 82, 490, 227, 52]. Open-Source [227]. Opening [656]. Operating [668]. Operations [160]. Operators [588]. Opinions [402]. Opponents [396]. Opportunities [237]. Optical [652]. Optimistic [274]. Optimization [422, 582, 271, 117, 269]. Optimized [579]. Order [557, 76, 80, 497]. Ordered [79, 582]. Ordering [80]. organising [304]. Organization [593, 373]. Organizational [382]. Organizations [382]. Orientation [227]. Orientational [497]. Oriented [583, 698]. OTS [129]. OTS/CafeOBJ [129]. Our [355]. Out-of-Plane [22]. Outreach [696]. Outsourcing [278, 624]. Ovarian [30]. Over-Complete [521]. Over-Fitting [283]. Overlap [101]. Overlap-Free [101]. Overlay [141]. Overseas [348, 349]. Overview [320, 227]. OWL [42, 59, 120]. OWL-DL [120].

P2P [270]. Pacific [696]. Packet [649, 256, 269]. Packet-Level [256]. Pairing [464]. Pairing-Based [464]. Pairs [211]. Pairwise [431]. Pancreatic [498]. Papers [529, 685, 687, 701]. Paradox [317]. Paradoxes [308]. Parallel [206, 286, 90, 211, 512]. Parameter [459, 481]. Parameters [311, 497, 18]. Paranga [390]. Paraphrase [215]. Paris [685]. Parsing [214, 220]. Part [676, 701]. Participatory [575]. Particle [322]. Particle-Wave [322]. Partition [84, 89, 287]. Partners [445]. Password [356, 357]. Passwords [254]. Path [466]. Pathologies [232]. Paths [312]. Pathway [496, 416]. Pattern [691, 158, 298, 205, 587, 694]. Pattern-Based [158, 205]. Patterns [294, 425, 61, 531, 8, 485, 599, 536, 386]. Payoff [562]. PCA [289]. PDGF [416]. Pedestrian [460]. Peg [217]. People [407, 342]. PerCAS [576]. Perceived [263]. Perception [527, 192, 328]. Perfect [274]. Performance [110, 527, 458, 62, 19, 271, 578, 596, 488, 664, 663]. Performances [407]. Performers [395]. Perfume [401]. Period [94]. Permission [281]. Permission-Based [281]. Permissions [547]. Persian [214]. Personality [383]. Personalized [660, 576]. Perspective [524, 134, 442, 630]. Perspectives [631, 538, 387]. PHACTS [52]. Phase [23]. Phones [379]. Photogrammetric [224]. Photogrammetry [234]. Photorealism [241]. Photos [398]. PHP [246]. Phrase [205]. Phrase-Level [205]. Physical [326]. Piecewise [20]. PKI [264]. Place [525]. Plane [22]. Planetarium [239]. Planners [456]. Platform [123, 492, 534, 394, 247]. Platforms [245]. Play [333]. Player [402]. Playful [404]. Point [334, 224, 513]. Points [248]. Poitiers [672]. Policy [637]. Polynomial [476, 473, 266]. Polynomial-Advantage [473]. Polynomials [657]. Population [410, 590]. Populations [597]. Portability [488]. Post [202]. Post-editing [202]. Potentially [371]. Power [139, 158, 54]. Power-Aware [139]. PRACSYS [456]. Practical [136, 204, 544]. Practice [125]. Practitioners [250]. Preaerobic [418]. pRebeca [555]. Precise [277]. Precision [334]. Predicates [557, 61]. Predicting [589, 499, 26]. Prediction [172, 4, 496, 209, 494, 495]. Predictive [4, 605]. Preference [66, 583]. Preliminary [187]. Premalignant [498]. Preparation [14]. Preprocessing [212]. Presence [558]. Present [396]. Presentation [407]. Presenting [530]. Preservation [681, 230]. Preserving [257, 273, 25, 569, 660, 275, 661, 477]. Preventing [251]. PRIB [694]. Prices [264]. Primal [427, 25]. Primal-Dual [25]. Prime [265]. Primitive [477]. Principal [493]. Principles [305]. Prior [298]. Priority [165].

Priority-OR [165]. Privacy [78, 257, 317, 273, 660, 366, 275, 259, 477]. Privacy-Friendly [259]. Privacy-Preserving [273, 275]. Private [398]. Probabilistic [76, 2, 555, 416]. Probabilistically [8]. Probabilities [431]. Probability [318]. Problem [99, 107, 150, 377, 558, 494, 568]. Problems [273]. Proceedings [676, 672, 677, 697, 674, 675, 696, 683, 686, 691, 700, 693, 684, 695, 681, 680, 682, 698, 689, 692, 678, 679, 699, 694, 688, 702, 671, 673]. Process [292, 69, 569, 372, 371, 600, 578]. Processes [81, 134, 544, 121, 690]. Processing [675, 680, 215, 116]. Processors [188]. Producing [398]. Product [411]. Production [158, 534]. Products [403]. Profile [168, 164]. Profiles [385]. Profiling [445]. Program [553]. Programming [604, 563, 692, 571, 548]. Programs [549, 88, 556]. Progress [681]. Project [229, 234]. Projected [221]. Prone [11, 12]. Proofs [92]. Propagating [49]. Propagation [569, 193]. Properties [84]. Propolis [492]. Prospecting [225]. Protecting [200]. Protection [168, 161]. Protein [410, 491, 497, 496, 494, 117, 487, 495, 481]. Proteomic [493]. Protocol [348, 349, 362, 363, 129]. Protocols [355, 347, 687, 664]. Prototype [424, 592, 603, 602]. Prototypes [542]. Provable [609]. Provably [560]. Provenance [66]. Proving [554]. Provisions [636]. Proxemic [60]. Proximity [301]. Proxy [657]. Pruning [593]. Pseudo [65, 255, 113]. Pseudo-Keys [65]. Pseudo-SSO [255]. Psychology [320]. Public [36, 642, 52, 664]. Publication [638, 660]. Publish [141]. Publish/Subscribe [141]. Putting [378]. Puzzle [379].

QI [685]. QMod [314, 315]. QoS [580, 572]. QoS-Aware [572]. Quadruped [408]. Qualitative [583]. Quality [36, 263, 8, 581, 207, 194, 538, 256]. Quantification [165]. Quantifiers [203]. Quantify [298]. Quantitative [632, 554, 199, 553]. Quantum [314, 308, 315, 309, 316, 311, 81, 305, 320, 321, 323, 304, 436, 685]. Quantum-Like [309, 316, 321]. Quasi [411]. Queries [579, 7, 646, 1, 4, 285, 67]. Query [53, 43]. Question [444]. Queuing [144].

Race [400]. Radio [152, 151]. Railway [168]. RAKAPOSHI [471]. Random [482, 499, 145, 113]. Randomization [78]. Randomized [140, 591, 154]. Rank [289]. Ranked [109]. Ranking [102, 66, 8]. Rate [334, 268]. Rates [415]. Ratio [529]. Rational [608]. Rationality [588]. Ray [27]. RDF [66, 61]. RDFS [69]. RDIS [455]. Re [41]. Re-classification [41]. Reachability [413]. Reaction [411, 491, 414]. Reactive [82]. Reader [244, 155]. Reader-Writer [155]. Real [130, 403, 273, 551, 393, 188, 23, 460, 306]. Real-Numbered [273]. Real-Time [130, 551, 188, 23]. Reality [240, 238]. Really [339]. Reasoning [444, 636, 2, 181]. Reasons [340]. Recognition [691, 485, 62, 520, 694, 546, 587]. Recombination [499]. Recommendation [300]. Recompression [503]. Reconfigurable [174]. Reconstruction [491, 167]. Reconstructions [239]. Recording [223]. Recovery [266].

Reduced [289, 267, 475, 582]. Reduced-Rank [289]. Reduced-Round [267]. Reduction [20, 295]. Referencing [68]. Referred [233]. Refinement [591]. Reflections [375]. Regeneration [236]. Regime [142]. Region [515, 518]. Regionalization [110]. Regions [235]. Registration [11, 25, 12]. Regression [482]. Regular [86]. Regularized [521]. Reha [124]. Reha-Sports [124]. Related [168, 189, 267, 196]. Related-Key [267]. Relation [509, 432]. Relational [439]. Relationship [173, 569]. Relationship-Preserving [569]. Relationships [530, 594, 577]. Relaxed [24]. Relay [616]. Relevance [494, 113]. Relevant [252]. Reliability [166, 678, 679]. Reliable [207]. Remains [235]. Remediation [666]. Remodelling [412]. Repeats [101, 104]. Repetitive [98]. Replicated [289]. Report [248]. Representation [316, 497, 63, 96]. Representations [423]. Representing [642, 60]. Requirement [163]. Requirements [380, 188, 250, 583, 386, 170]. Research [631, 531, 52, 342]. Researcher [219]. Residential [513]. Resilience [666]. Resilient [272]. Resistance [589]. Resolution [511, 417, 568, 481]. Resource [696, 551, 215, 217]. Resource-Aware [551]. Response [593]. Response-Style [593]. Restoration [213]. Result [110, 650]. Resulting [391]. Results [597]. Retooling [253]. Retrieval [100, 109, 675, 70, 312, 29, 119, 525]. Reuse [181]. Revealed [331]. Reversibility [279]. Review [258]. Revised [685, 687, 701]. Revisited [286, 317]. Reviving [238]. Revocation [264]. Rewriting [552, 546]. Rich [218, 399, 215]. Richer [418]. Rights [632]. Risk [609, 159, 160, 623, 161]. RMR [154]. RMR-Efficient [154]. RNA [488]. Roads [629]. Robot [445, 345, 338, 459, 448, 343, 453, 328, 454, 332, 335, 340, 329, 330, 333, 341]. Robotic [457, 461, 550, 339, 446, 337]. Robotics [452, 686, 336, 630, 342]. RobotML [457]. Robots [450, 345, 327, 339, 692, 629, 460]. Robotubby [337]. Robust [511, 662, 119, 434, 481]. Robustness [615]. Role [39, 220, 122]. Role-Based [122]. Rome [238]. ROS [455]. Rough [603, 602]. Round [277, 267, 467, 217]. Route [167]. Routing [616]. RSA [265]. Rule [593, 415]. Rule-Based [415]. Rules [296, 213]. Run [654]. Run-Time [654]. Runtime [35]. Rusks [403].

Safe [158]. SAFECOMP [678, 679]. Safer [195]. Safety [183, 168, 157, 189, 163, 164, 196, 37, 182, 251, 678, 679, 194, 176, 650, 181, 630, 184]. Safety-Critical [163]. Safety-Related [168, 189, 196]. Salience [516]. Salient [515, 518, 517]. Salvador [677]. Sampling [459]. SanjigenJiten [406]. Sassur [679, 179]. SAT [302]. Satellite [372, 176, 174]. Satisfaction [538]. Scaffold [410]. Scalable [280, 652, 537]. Scalar [464]. Scale [491, 233, 433, 670, 483, 372, 434, 438, 517]. Scale-Based [483]. Scaled [319]. Scales [305]. Scanner [230]. Scanning [246, 242]. Scans [27]. Scenarios [352, 353, 237, 461, 613]. Scenes [233]. Scheduling [570]. Schema [535]. Schemas [36]. Scheme [263, 148, 470, 665, 657, 663]. Schemes [68, 468]. School [539]. Science [674, 3]. Scientific [57]. Scientists [531]. Scoring [482]. Search [136, 95, 102, 110, 140, 527, 40, 275, 587]. Searching [526]. Second [76]. Secondary [539, 497, 494, 495]. Secrecy [467]. Secret [665, 608, 272]. Secrets [523]. Secure [348, 349, 478, 169, 276, 477, 255, 682]. Securing [253]. Security [245, 622, 168, 621, 352, 353, 159, 683, 279, 249, 700, 617, 693, 654, 668, 476, 623, 656, 250, 252, 199, 364, 365, 678, 679, 620, 248, 247, 702, 687]. Security-Relevant [252]. Seeking [426, 536]. Segmentation [433, 506, 30, 24, 118, 512]. Segmenting [211]. Segments [372]. Select [250]. Selected [195, 685, 687, 701]. Selection [102, 283, 424, 44, 484, 618, 580, 603, 602, 584, 573, 601]. Selective [320, 656, 117]. Self [136, 350, 351, 304, 417, 537]. Self-Adjusting [136]. Self-assembly [417]. Self-enforcing [350, 351]. Self-management [537]. Self-organising [304]. Semantic [643, 509, 219, 636, 319, 220, 52, 525, 306, 96, 56]. Semantically [7, 215]. Semantically-Annotated [7]. Semantically-Rich [215]. Semantics [549, 33, 359, 358, 90, 64, 526]. Semi [97, 150, 27]. Semi-automatic [27]. Semi-matching [150]. Semi-supervision [97]. Sensemaking [380]. Sensitive [30]. Sensor [334, 193]. Sentence [211, 641]. Sentencing [641]. Separation [557]. September [674, 678, 679]. Sequence [45]. Sequences [104, 107]. Sequentially [138]. Serial [418]. Series [297, 595]. Servers [578]. Service [567, 238, 264, 570, 123, 581, 343, 583, 580, 574, 577, 268, 584, 573, 575, 572, 698]. Service-Oriented [698]. Services [280, 635, 531, 571, 576]. Set [36, 667]. Sets [298, 596, 602]. Setting [149]. Shallow [212]. Shanghai [698]. Shannon [429, 440]. Shape [20, 29, 437, 516, 119, 24]. Shapes [507]. Shared [570, 398, 519]. Sharing [637, 608, 538, 659, 272]. Shawbak [234]. Shift [426]. Shifted [502]. Short [297, 466]. Show [138]. SIEM [200]. Sign [290]. Signaling [613, 416]. Signature [465, 657]. Signatures [276]. Signcryption [658]. Similar [515]. Similarity [532]. SIMPAR [692]. Simple [126, 122, 645, 496, 500]. Simulate [457]. Simulated [453]. Simulating [461]. Simulation [449, 450, 452, 16, 476, 483, 614, 454, 17, 692]. Simulator [459, 460]. Single [74, 334, 507]. Single-Point [334]. Sink [257]. SINR [143]. Situational [653]. Size [263, 591]. Skew [289]. Skew- [289]. Skewed [596]. Skyline [1]. SLA [582]. Slide [471]. Slope [595]. SM3 [271]. Small [265, 124, 149]. Smart [653, 236, 623, 192, 447, 193]. Smartphone [376]. SME [249]. Smile [328]. Smiles [403]. Smooth [505]. SMOTE [603]. Snapshot [552]. Snippet [219]. SNR [142]. SOA [565]. Social [236, 327, 339, 301, 344, 332, 55, 568, 335, 397, 536, 329, 398, 330, 333, 342, 686]. Socially [396]. Societies [304]. Socio [608]. Socio-Rational [608]. Soft [175]. Software [697, 195, 382, 33, 189, 490, 373, 640, 369, 454, 701, 47, 161, 488, 368, 544, 688, 384]. Solid [510]. Solids [510]. Solutions [227, 375]. SOM [600]. Some [588]. Sound [399, 392]. Sound-Based [392]. Sounds [391]. Source [47, 227]. Sources [56]. Space [104, 109, 435, 434, 438, 447]. Space-Efficient [104]. Spammers [667]. SPARQL [67]. Sparse [427].

Spatio [39]. Spatio-Temporal [39]. Special [670, 690]. Specific [352, 353, 457, 340, 532]. Specification [130, 39, 164, 565, 552, 170]. Specifications [138]. Spectra [474]. Spectral [235, 233]. Spectroscopic [492]. Speech [546]. Spike [291]. SPIRE [675]. Split [272]. Split-State [272]. Spoofing [649]. Sports [389, 124]. Square [217]. SSO [255]. SSPR&SPR [691]. Stable [192]. Stack [75]. Stacks [651]. stage [490]. Standard [467]. Standardized [465]. Standards [184]. State [449, 84, 19, 272]. State-of-the-Art [19]. State-Partition [84]. Stateful [468]. States [326, 310]. Static [74]. stationary [663]. Statistical [356, 357, 214, 691, 24, 269, 212]. Statistics [435]. Staying [67]. Steganalysis [262]. Step [176, 613]. Stepmother [77]. Stochastic [144]. Storage [259, 255]. Storytelling [405]. Strategic [618]. Strategies [459, 666]. Strategy [299]. Stream [471]. Street [359, 358]. Street-Level [359, 358]. Strength [356, 357, 162]. Stress [235]. String [99, 101, 675]. Strings [428]. Strong [467]. Structural [231, 20, 377, 582, 489, 691]. Structure [29, 143, 423, 497, 494, 495, 434]. Structured [5, 469]. Structures [43]. Students [539, 250]. Studies [676]. Study [100, 412, 531, 250, 541, 597, 329, 333]. Stuffed [395]. Style [535, 593]. Subdivision [514]. Subgraphs [149, 146]. Subgroup [383]. Subject [57, 526]. Subject-Based [526]. Subjective [298]. Sublogarithmic [148]. Sublogarithmic-Time [148]. Subscribe [141]. Subsequence [107]. Subset [484]. Subspace [519]. Suffix [98]. Suits [395]. Summarization [108]. Super [476]. Super-Polynomial-Simulation [476]. Supermaximal [104]. Superpixels [518]. Supervised [258, 21, 533]. Supervision [108, 97]. Supine [11, 12]. Support [280, 482, 7, 668, 540, 596, 386]. Supporting [69, 395, 193]. Supreme [633]. Surface [13]. Surfaces [514]. Surveillance [208]. Survey [402]. Survival [599, 633]. Suspects [258]. SVC [663]. SVM [596]. SW [171]. SW-FMEA [171]. Sweden [682]. SwingScape [404]. SWRL [42]. Symmetric [470]. Symmetric-Key [470]. Symmetries [436]. Symmetry [152]. Symposium [677, 685, 675, 684, 676]. Syntactic [588, 691, 220, 212]. Synthesis [37, 182, 510]. Synthesis-Centric [182]. System [5, 635, 621, 400, 234, 167, 406, 668, 207, 62, 407, 199, 395, 586, 128, 528, 174, 702]. Systematization [171]. Systemic [253]. Systems [567, 318, 172, 412, 566, 132, 642, 413, 293, 414, 133, 321, 134, 224, 82, 130, 550, 166, 670, 158, 83, 682, 223, 187, 560, 196, 182, 372, 421, 194, 701, 690, 598, 627, 577, 561, 555, 186, 193].

Tactile [332]. Tag [300]. Tagging [17]. Tagset [218, 216]. Taipei [696]. Taiwan [696]. Target [619]. Task [130]. Tasks [444]. Taste [557]. TC [674]. TCS [674]. Team [389]. Technique [16, 437, 516, 17]. Techniques [292, 262, 422, 386, 162]. Technology [257, 200]. Temperature [334]. Templates [23, 119]. Temporal [165, 39, 563, 221, 88, 562, 548]. tenanted [578]. Term [209, 546]. Terminology [210]. Terrestrial [242]. TESLA [129]. Test [164, 269]. Testability [163]. Tested [83]. Testing [71, 463, 162, 542]. Testing- [71]. Text [103, 68, 533]. Texts [98, 206, 643, 213]. Texture [514, 510]. Thailand [539]. THAPS [246]. Theatrical [407]. Their [24]. Theorem [554]. Theoretic [617, 424, 614, 618]. Theoretical [610, 506, 674]. Theory [314, 315, 80, 700, 2, 125, 63]. Therapies [312]. There [251, 341]. Thermal [232]. Third [700, 692]. threaded [88]. Three [319, 629]. Thresholding [299]. Throughput [297, 118]. Tickets [570]. Time [101, 289, 450, 130, 148, 563, 654, 551, 188, 297, 23, 436, 595]. Timed [129]. Timing [126, 125, 196, 192, 128]. Tissue [27]. Tokyo [694]. Tolerant [537]. Tool [71, 131, 384]. Toolset [180]. Toothbrush [391]. Topic [43]. Topological [13]. Torts [629]. Touch [379, 392]. Touch-Based [379]. Toulouse [688]. Trace [449]. Traces [256]. Track [400, 259]. Tracking [219]. Trade [371]. Trade-Offs [371]. Traditional [403]. Traffic [4, 269]. Train [291]. Trajectory [290, 660]. Transactional [549, 594]. Transactions [670, 690]. Transcript [349, 363, 355, 357, 353, 347, 361, 359, 351, 365]. Transfer [476, 655]. Transformation [42, 559, 39]. Transformations [559]. Transitions [418]. Translating [368]. Translation [202, 203, 204, 209, 205, 212]. Translocon [483]. Transportation [133]. Tree [136, 485, 44, 587]. Trees [98, 612, 165, 430, 599, 216]. Trend [410, 532]. Trend-Based [410]. Tri [149]. Trial [16, 333]. Triangles [149]. Trip [296]. Trust [359, 358, 639]. Trustworthy [186]. TSO [138]. Tsukuba [692]. Tumor [30]. Type [317, 209].

Ubiquitous [655]. UK [687, 360, 361, 336, 387]. Ultra [465, 16, 15]. Ultra-constrained [465]. Ultra-Low-Dose [16, 15]. UML [46]. Unattributed [424]. Uncertain [580]. Unconditionally [665]. Uncovering [523]. Undecidable [80]. Understand [333]. Understanding [439]. Unidirectional [83]. Unified [186, 486]. Unifying [134]. Union [85]. Union-Free [85]. Unique [399, 90]. Units [568]. Universal [59, 658]. University [529]. Unknown [621]. Unmet [340]. Unsupervised [590]. Unwinding [252]. Updates [6, 69]. Upper [612]. Urban [404, 513]. URI [68]. Usability [377, 376, 364, 365]. Usable [366, 372]. Usage [95, 650]. USB [244]. USB-Connected [244]. Use [531, 204, 421, 194, 384]. Used [374]. User [379, 381, 646, 339, 373, 371, 541, 387, 35, 368, 333, 542]. Users [257, 34, 339, 53, 385]. Using [383, 334, 108, 559, 509, 604, 258, 380, 523, 219, 425, 403, 11, 57, 296, 455, 498, 484, 613, 499, 301, 369, 649, 302, 497, 554, 23, 600, 578, 574, 26, 117, 269, 555, 333, 43, 510, 287, 113, 300]. Utility [612].

V [499]. V&V [134]. Validation [42, 6, 449, 453, 676, 12]. Value [192]. valued [70]. Values [632]. Variability [573]. Variable [547, 417, 447]. Variation [297]. Various [341]. Vector [482, 23, 494]. Vectors [596, 306]. Vegetation [235]. Vehicles [195, 194]. Verifiability [468]. Verification [412, 566, 115, 559, 130, 550, 164, 87, 547, 676, 649, 47, 555, 556, 542]. Verifier [658]. Versions [475]. versus [77, 327]. Vertex [148]. Very [466].

Vessel [28]. VI [670]. via [66, 450, 667, 200, 145, 418, 601]. Video [541]. Videogames [394]. Videos [23]. View [387, 659]. viewpoints [41]. Violations [579]. Virtual [238, 403, 393, 242, 16, 241, 17, 239]. Virtualized [551]. Visual [263, 449, 379, 695, 455, 366, 13]. Visual-Based [379]. Visual-Trace [449]. Visualisation [228]. Visualization [593, 600, 488]. Volumes [21]. Volumetric [14]. Voters [348, 349]. Voting [348, 349, 360, 361, 350, 351, 665]. vs [458]. Vulnerability [617, 246].

Walks [436, 145, 113]. Warcraft [402]. Warehouse [40, 49]. Warning [248]. Watershed [118]. Wave [322]. Wavelet [111]. WCET [131]. WCP [584]. WCP-Nets [584]. Weak [79, 90]. Weakly [127]. Weakly-Hard [127]. Weakness [138]. Wearable [401]. Web [65, 95, 110, 604, 219, 8, 385, 210, 590, 584, 170]. Web-Based [210]. Websites [258]. Weighted [428, 584, 288]. Weighting [287]. Weights [485]. Well [339]. WG [674]. Where [607]. Which [176]. while [67, 398, 661]. Who [568]. Wicked [77]. Wikification [96]. Wikinews [108]. Wikipedia [57]. Wireless [142, 662, 143, 652, 663]. within [158, 406, 398, 342]. without [220]. Women [599]. Word [322, 319, 214, 211, 508]. Workflows [122, 120]. Workload [173]. Workshop [687, 691, 693, 701, 671]. Workshops [679]. World [402]. Worst [78]. Worst- [78]. Would [338]. Writer [155]. WSN [200]. Wuyishan [702].

X [27]. X-Ray [27]. XI [684]. XIV [690]. XML [6]. XX [687].

Yuan [523].

Zero [277]. Zero-Knowledge [277]. Zipf [108]. Zipf-Like [108]. Zones [200].

References

Guo:2012:CSQ

[1] Xi Guo, Chuan Xiao, and Yoshiharu Ishikawa. Combination skyline queries. Lecture Notes in Computer Science, 7600:1–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

     link.springer.com/chapter/10.1007/978-3-642-34179-3_1/.

Kern-Isberner:2012:CEA

[2] Gabriele Kern-Isberner and Christoph Beierle. Comparing and evaluating approaches to probabilistic reasoning: Theory, implementation, and applications. Lecture Notes in Computer Science, 7600:31–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

     link.springer.com/chapter/10.1007/978-3-642-34179-3_2/.

Thalheim:2012:SAC

[3] Bernhard Thalheim. The science and art of conceptual modelling. Lecture Notes in Computer Science, 7600:76–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34179-3_3/.

Heendaliya:2012:PLQ

[4] Lasanthi Heendaliya, Dan Lin, and Ali Hurson. Predictive line queries for traffic prediction. Lecture Notes in Computer Science, 7600:106– 133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34179-3_4/.

Asiki:2012:DBS

[5] Athanasia Asiki and Dimitrios Tsoumakos. A DHT-based system for the management of loosely structured, multidimensional data. Lecture Notes in Computer Science, 7600:134–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34179-3_5/.

Bouchou:2012:GIV

[6] Béatrice Bouchou and Mirian Halfeld-Ferrari. A grammarware for the incremental validation of integrity constraints on XML documents under multiple updates. Lecture Notes in Computer Science, 7600:167– 197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34179-3_6/.

Cao:2012:DSE

[7] Huiping Cao and Shawn Bowers. Database support for enabling datadiscovery queries over semantically-annotated observational data. Lecture Notes in Computer Science, 7600:198–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34179-3_7/.

Han:2012:PRW

[8] Jingyu Han, Kejia Chen, and Dawei Jiang. Probabilistically ranking Web article quality based on evolution patterns. Lecture Notes in Computer Science, 7600:229–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34179-3_8/.

Anonymous:2012:BMa

[9] Anonymous. Back matter. Lecture Notes in Computer Science, 7600: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34179-3/1.

Anonymous:2012:FMa

[10] Anonymous. Front matter. Lecture Notes in Computer Science, 7600: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34179-3/1.

Hampshire:2012:PSC

[11] Thomas E. Hampshire, Holger R. Roth, Darren J. Boone, Greg Slabaugh, and Steve Halligan. Prone to supine CT colonography registration using a landmark and intensity composite method. Lecture Notes in Computer Science, 7601:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33612-6_1/.

Roth:2012:ECV

[12] Holger R. Roth, Darren J. Boone, Steve Halligan, and Thomas E. Hampshire. External clinical validation of prone and supine CT colonography registration. Lecture Notes in Computer Science, 7601:10–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33612-6_2/.

Shi:2012:ETC

[13] Rui Shi, Wei Zeng, Jerome Zhengrong Liang, and Xianfeng David Gu. Efficient topological cleaning for visual colon surface flattening. Lecture Notes in Computer Science, 7601:20–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33612-6_3/.

Nappi:2012:AVD

[14] Janne J. Näppi, Se Hyung Kim, and Hiroyuki Yoshida. Adaptive volumetric detection of lesions for minimal-preparation dual-energy CT colonography. Lecture Notes in Computer Science, 7601:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33612-6_4/.

Nappi:2012:CAD

[15] Janne J. Näppi, Masanori Imuta, Yasuyuki Yamashita, and Hiroyuki Yoshida. Computer-aided detection for ultra-low-dose CT colonography. Lecture Notes in Computer Science, 7601:40–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33612-6_5/.

Kim:2012:ACS

[16] Chang Won Kim and Jong Hyo Kim. Application of CT simulation technique for virtual ultra-low-dose trial in CT colonography. Lecture Notes in Computer Science, 7601:49–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33612-6_6/.

Yang:2012:ACS

[17] Zepa Yang, Hyeong min Jin, and Jong Hyo Kim. Application of CT simulation technique for virtual fecal tagging in CTC. Lecture Notes in Computer Science, 7601:58–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33612-6_7/.

Nappi:2012:ACA

[18] Janne J. Näppi, Don Rockey, Daniele Regge, and Hiroyuki Yoshida. Application of CT acquisition parameters as features in computer-aided detection for CT colonography. Lecture Notes in Computer Science, 7601: 69–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33612-6_8/.

Lee:2012:CPS

[19] Sang Ho Lee, Janne J. Näppi, and Hiroyuki Yoshida. Comparative performance of state-of-the-art classifiers in computer-aided detection for CT colonography. Lecture Notes in Computer Science, 7601:78– 87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33612-6_9/.

Cai:2012:PSD

[20] Wenli Cai, June-Goo Lee, Da Zhang, Christina Piel, and Hiroyuki Yoshida. Piecewise structural diffusion defined on shape index for noise reduction in dual-energy CT images. Lecture Notes in Computer Science, 7601:88–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33612-6_10/.

Mahapatra:2012:SLB

[21] Dwarikanath Mahapatra, Peter Schueffler, Jeroen A. W. Tielbeek, and Joachim M. Buhmann. A supervised learning based approach to detect Crohn’s disease in abdominal MR volumes. Lecture Notes in Computer Science, 7601:97–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33612-6_11/.

Yigitsoy:2012:PMC

[22] Mehmet Yigitsoy, Sonja Kirchhoff, Maximilian F. Reiser, and Nassir Navab. Out-of-plane motion compensation in cine-MRI. Lecture Notes in Computer Science, 7601:107–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33612-6_12/.

Nawarathna:2012:RTP

[23] Ruwan Nawarathna, JungHwan Oh, Jayantha Muthukudage, Wallapak Tavanapong, and Johnny Wong. Real-time phase boundary detection for colonoscopy videos using motion vector templates. Lecture Notes in Computer Science, 7601:116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33612-6_13/.

Tomoshige:2012:RCS

[24] Sho Tomoshige, Elco Oost, Akinobu Shimizu, Hidefumi Watanabe, and Hidefumi Kobatake. Relaxed conditional statistical shape models and their application to non-contrast liver segmentation. Lecture Notes in Computer Science, 7601:126–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33612-6_14/.

Kiriyanthan:2012:LBP

[25] Silja Kiriyanthan, Ketut Fundana, Tahir Majeed, and Philippe C. Cattin. A landmark-based primal-dual approach for discontinuity preserving registration. Lecture Notes in Computer Science, 7601:137–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33612-6_15/.

Samei:2012:PLM

[26] Golnoosh Samei, Christine Tanner, and Gabor Székely. Predicting liver motion using exemplar models. Lecture Notes in Computer Science, 7601: 147–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33612-6_16/.

Upadhyay:2012:SAD

[27] Sanat Upadhyay, Manos Papadakis, Saurabh Jain, and Gregory Gladish. Semi-automatic discrimination of normal tissue and liver cancer lesions in contrast enhanced X-ray CT-scans. Lecture Notes in Computer Science, 7601:158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33612-6_17/.

Luu:2012:EDF

[28] Ha Manh Luu, Adriaan Moelker, Camiel Klink, Adrienne Mendrik, and Wiro Niessen. Evaluation of diffusion filters for 3D CTA liver vessel enhancement. Lecture Notes in Computer Science, 7601:168–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33612-6_18/.

Huang:2012:SAL

[29] Weimin Huang, Wei Xiong, Jiayin Zhou, Jing Zhang, Tao Yang, Jiang Liu, and Yi Su. 3D shape analysis for liver-gallbladder anatomical structure retrieval. Lecture Notes in Computer Science, 7601:178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33612-6_19/.

Liu:2012:TSM

[30] Jianfei Liu, Shijun Wang, Marius G. Linguraru, and Ronald M. Summers. Tumor sensitive matching flow: An approach for ovarian cancer metastasis detection and segmentation. Lecture Notes in Computer Science, 7601: 188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33612-6_20/.

Anonymous:2012:FMb

[31] Anonymous. Front matter. Lecture Notes in Computer Science, 7601: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33612-6/1.

Papazoglou:2012:CBM

[32] Mike P. Papazoglou. Cloud blueprint: a model-driven approach to configuring federated clouds. Lecture Notes in Computer Science, 7602:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33609-6_1.

Feron:2012:MBA

[33] Eric Féron. Model-based auto coding of embedded control software with full semantics. Lecture Notes in Computer Science, 7602:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33609-6_2.

Karam:2012:MEU

[34] Roula Karam, Piero Fraternali, Alessandro Bozzon, and Luca Galli. Modeling end-users as contributors in human computation applications. Lecture Notes in Computer Science, 7602:3–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33609-6_3/.

Rodriguez-Gracia:2012:RAA

[35] Diego Rodríguez-Gracia, Javier Criado, Luis Iribarne, and Nicolás Padilla. Runtime adaptation of architectural models: An approach for adapting user interfaces. Lecture Notes in Computer Science, 7602:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33609-6_4/.

Batini:2012:MQL

[36] Carlo Batini, Marco Comerio, and Gianluigi Viscusi. Managing quality of large set of conceptual schemas in public administration: Methods and experiences. Lecture Notes in Computer Science, 7602:31–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33609-6_5/.

Maitrehenry:2012:MBS

[37] Sébastien Maitrehenry, Sylvain Metge, Yamine Aït-Ameur, and Pierre Bieber. An MDE-based synthesis of aircraft safety models. Lecture Notes in Computer Science, 7602:43–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_6/.

Djeddai:2012:IFD

[38] Selma Djeddai, Martin Strecker, and Mohamed Mezghiche. Integrating a formal development for DSLs into meta-modeling. Lecture Notes in Computer Science, 7602:55–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_7/.

Geepalla:2012:TST

[39] Emsaieb Geepalla, Behzad Bordbar, and Joel Last. Transformation of spatio-temporal role based access control specification to alloy. Lecture Notes in Computer Science, 7602:67–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33609-6_8/.

Mustapha:2012:MOW

[40] Nesrine Ben Mustapha, Marie-Aude Aufaure, Hajer Baazaoui Zghal, and Henda Ben Ghezala. Modular ontological warehouse for adaptative information search. Lecture Notes in Computer Science, 7602: 79–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33609-6_9/.

Djezzar:2012:ORC

[41] Meriem Djezzar, Mounir Hemam, and Zizette Boufaida. Ontological reclassification of individuals: a multi-viewpoints approach. Lecture Notes in Computer Science, 7602:91–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_10/.

Almendros-Jimenez:2012:TVS

[42] Jesús M. Almendros-Jiménez and Luis Iribarne. Transformation and validation with SWRL and OWL of ODM-based models. Lecture Notes in Computer Science, 7602:103–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_11/.

Yanagisawa:2012:CBQ

[43] Takashi Yanagisawa and Takao Miura. Context-based query using dependency structures based on latent topic model. Lecture Notes in Computer Science, 7602:116–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33609-6_12/.

Karabadji:2012:DTS

[44] Nour El Islem Karabadji, Hassina Seridi, Ilyes Khelf, and Lakhdar Laouar. Decision tree selection in an industrial machine fault diagnostics. Lecture Notes in Computer Science, 7602:129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33609-6_13/.

Dahi:2012:CCI

[45] Fairouz Dahi and Nora Bounour. Crosscutting concerns identification approach based on the sequence diagram analysis. Lecture Notes in Computer Science, 7602:141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_14/.

Barbier:2012:IU

[46] Franck Barbier and Eric Cariou. Inductive UML. Lecture Notes in Computer Science, 7602:153–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_15/.

Pires:2012:FSV

[47] Anthony Fernandes Pires, Thomas Polacsek, and Stéphane Duprat. Formal software verification at model and at source code levels. Lecture Notes in Computer Science, 7602:162–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33609-6_16/.

Soussi:2012:EOL

[48] Rania Soussi and Marie-Aude Aufaure. Enterprise ontology learning for heterogeneous graphs extraction. Lecture Notes in Computer Science, 7602:170–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33609-6_17/.

Taktak:2012:TPE

[49] Saïd Taktak and Jamel Feki. Toward propagating the evolution of data warehouse on data marts. Lecture Notes in Computer Science, 7602: 178–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33609-6_18/.

Anonymous:2012:BMb

[50] Anonymous. Back matter. Lecture Notes in Computer Science, 7602: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33609-6/1.

Anonymous:2012:FMc

[51] Anonymous. Front matter. Lecture Notes in Computer Science, 7602: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33609-6/1.

Harland:2012:OPS

[52] Lee Harland. Open PHACTS: a semantic knowledge infrastructure for public and commercial drug discovery research. Lecture Notes in Computer Science, 7603:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33876-2_1/.

Pradel:2012:AEU

[53] Camille Pradel. Allowing end users to query graph-based knowledge bases. Lecture Notes in Computer Science, 7603:8–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33876-2_2/.

deBoer:2012:NHP

[54] Victor de Boer, Michiel Hildebrand, Lora Aroyo, Pieter De Leenheer, and Chris Dijkshoorn. Nichesourcing: Harnessing the power of crowds of experts. Lecture Notes in Computer Science, 7603:16–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33876-2_3/.

Schneider:2012:DAS

[55] Jodi Schneider, Brian Davis, and Adam Wyner. Dimensions of argumentation in social media. Lecture Notes in Computer Science, 7603: 21–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33876-2_4/.

dAmato:2012:SKD

[56] Claudia d’Amato, Volha Bryl, and Luciano Serafini. Semantic knowledge discovery from heterogeneous data sources. Lecture Notes in Computer Science, 7603:26–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33876-2_5/.

Joorabchi:2012:ASM

[57] Arash Joorabchi and Abdulhussain E. Mahdi. Automatic subject metadata generation for scientific documents using Wikipedia and genetic algorithms. Lecture Notes in Computer Science, 7603:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33876-2_6/.

Ferre:2012:ADE

[58] Sébastien Ferré and Sebastian Rudolph. Advocatus Diaboli — exploratory enrichment of ontologies with negative constraints. Lecture Notes in Computer Science, 7603:42–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33876-2_7/.

Buhmann:2012:UOA

[59] Lorenz Bühmann and Jens Lehmann. Universal OWL axiom enrichment for large knowledge bases. Lecture Notes in Computer Science, 7603: 57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33876-2_8/.

Latiri:2012:PCN

[60] Chiraz Latiri, Lamia Ben Ghezaiel, and Mohamed Ben Ahmed. Proxemic conceptual network based on ontology enrichment for representing documents in IR. Lecture Notes in Computer Science, 7603:72– 86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33876-2_9/.

Gerber:2012:EMN

[61] Daniel Gerber and Axel-Cyrille Ngonga Ngomo. Extracting multilingual natural-language patterns for RDF predicates. Lecture Notes in Computer Science, 7603:87–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33876-2_10/.

Kim:2012:IPN

[62] Myung Hee Kim and Paul Compton. Improving the performance of a named entity recognition system with knowledge acquisition. Lecture Notes in Computer Science, 7603:97–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33876-2_11/.

Presutti:2012:KEB

[63] Valentina Presutti, Francesco Draicchio, and Aldo Gangemi. Knowledge extraction based on discourse representation theory and linguistic frames. Lecture Notes in Computer Science, 7603:114–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33876-2_12/.

Tonelli:2012:ISF

[64] Sara Tonelli, Volha Bryl, Claudio Giuliano, and Luciano Serafini. Investigating the semantics of frame elements. Lecture Notes in Computer Science, 7603:130–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33876-2_13/.

Atencia:2012:KPK

[65] Manuel Atencia, Jérôme David, and François Scharffe. Keys and pseudokeys detection for Web datasets cleansing and interlinking. Lecture Notes in Computer Science, 7603:144–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33876-2_14/.

Dividino:2012:RRP

[66] Renata Dividino, Gerd Gröner, Stefan Scheglmann, and Matthias Thimm. Ranking RDF with provenance via preference aggregation. Lecture Notes in Computer Science, 7603:154–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33876-2_15/.

Umbrich:2012:FWS

[67] Jürgen Umbrich, Marcel Karnstedt, Aidan Hogan, and Josiane Xavier Parreira. Freshening up while staying fast: Towards hybrid SPARQL queries. Lecture Notes in Computer Science, 7603:164–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33876-2_16/.

Hellmann:2012:LDA

[68] Sebastian Hellmann, Jens Lehmann, and Sören Auer. Linked-data aware URI schemes for referencing text fragments. Lecture Notes in Computer Science, 7603:175–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33876-2_17/.

Hermann:2012:IGP

[69] Alice Hermann, Sébastien Ferré, and Mireille Ducassé. An interactive guidance process supporting consistent updates of RDFS graphs. Lecture Notes in Computer Science, 7603:185–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33876-2_18/.

Campinas:2012:ERM

[70] Stéphane Campinas, Renaud Delbru, and Giovanni Tummarello. Effective retrieval model for entity with multi-valued attributes: BM25MF and beyond. Lecture Notes in Computer Science, 7603:200–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33876-2_19/.

Blomqvist:2012:OTM

[71] Eva Blomqvist, Azam Seil Sepour, and Valentina Presutti. Ontology testing- methodology and tool. Lecture Notes in Computer Science, 7603: 216–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33876-2_20/.

Anonymous:2012:FMd

[72] Anonymous. Front matter. Lecture Notes in Computer Science, 7603: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33876-2/1.

Wiedermann:2012:CNC

[73] Jiří Wiedermann. Computability and non-computability issues in amorphous computing. Lecture Notes in Computer Science, 7604:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33475-7_1/.

Ancona:2012:SSI

[74] Davide Ancona and Giovanni Lagorio. Static single information form for abstract compilation. Lecture Notes in Computer Science, 7604: 10–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33475-7_2/.

Bensch:2012:IDS

[75] Suna Bensch, Markus Holzer, Martin Kutrib, and Andreas Malcher. Input-driven stack automata. Lecture Notes in Computer Science, 7604: 28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33475-7_3/.

Bodlaender:2012:PIM

[76] Marijke Hans L. Bodlaender. Probabilistic inference and monadic second order logic. Lecture Notes in Computer Science, 7604:43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33475-7_4/.

Bodlaender:2012:CVW

[77] Marijke Hans L. Bodlaender and Cor A. J. Hurkens. Cinderella versus the wicked stepmother. Lecture Notes in Computer Science, 7604: 57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33475-7_5/.

Boreale:2012:WAC

[78] Michele Boreale and Michela Paolini. Worst- and average-case privacy breaches in randomization mechanisms. Lecture Notes in Computer Science, 7604:72–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33475-7_6/.

Brengos:2012:WBC

[79] Tomasz Brengos. Weak bisimulations for coalgebras over ordered functors. Lecture Notes in Computer Science, 7604:87–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33475-7_7/.

Carayol:2012:CFL

[80] Arnaud Carayol and Zoltán Ésik. A context-free linear ordering with an undecidable first-order theory. Lecture Notes in Computer Science, 7604: 104–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33475-7_8/.

Deng:2012:OBQ

[81] Yuxin Deng and Yuan Feng. Open bisimulation for quantum processes. Lecture Notes in Computer Science, 7604:119–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33475-7_9/.

Gadducci:2012:MLO

[82] Fabio Gadducci, Giacoma Valentina Monreale, and Ugo Montanari. A modular LTS for open reactive systems. Lecture Notes in Computer Science, 7604:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33475-7_10/.

Jancar:2012:UCS

[83] Petr Jančar, Prateek Karandikar, and Philippe Schnoebelen. Unidirectional channel systems can be tested. Lecture Notes in Computer Science, 7604:149–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33475-7_11/.

Jiraskova:2012:PSC

[84] Galina Jirásková and Tomáš Masopust. On properties and state complexity of deterministic state-partition automata. Lecture Notes in Computer Science, 7604:164–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33475-7_12/.

Jiraskova:2012:UFD

[85] Galina Jirásková and Benedek Nagy. On union-free and deterministic union-free languages. Lecture Notes in Computer Science, 7604: 179–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33475-7_13/.

Kurz:2012:CLI

[86] Alexander Kurz, Tomoyuki Suzuki, and Emilio Tuosto. A characterisation of languages on infinite alphabets with nominal regular expressions. Lecture Notes in Computer Science, 7604:193–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33475-7_14/.

Kufner:2012:FVD

[87] Philipp Küfner, Uwe Nestmann, and Christina Rickmann. Formal verification of distributed algorithms. Lecture Notes in Computer Science, 7604: 209–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33475-7_15/.

LaTorre:2012:TLM

[88] Salvatore La Torre and Margherita Napoli. A temporal logic for multithreaded programs. Lecture Notes in Computer Science, 7604:225–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33475-7_16/.

Liang:2012:ACD

[89] Hongyu Liang. The algorithmic complexity of k-domatic partition of graphs. Lecture Notes in Computer Science, 7604:240–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33475-7_17/.

Luttik:2012:UPD

[90] Bas Luttik. Unique parallel decomposition in branching and weak bisimulation semantics. Lecture Notes in Computer Science, 7604:250–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33475-7_18/.

Luttgen:2012:MIA

[91] Gerald Lüttgen and Walter Vogler. Modal interface automata. Lecture Notes in Computer Science, 7604:265–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33475-7_19/.

Beffara:2012:PE

[92] Emmanuel Beffara and Virgile Mogbil. Proofs as executions. Lecture Notes in Computer Science, 7604:280–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33475-7_20/.

Anonymous:2012:FMe

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33475-7/1.

Amir:2012:APD

[94] Amihood Amir and Avivit Levy. Approximate period detection and correction. Lecture Notes in Computer Science, 7608:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34109-0_1/.

Baeza-Yates:2012:UDW

[95] Ricardo Baeza-Yates and Yoelle Maarek. Usage data in Web search: Benefits and limitations. Lecture Notes in Computer Science, 7608:16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34109-0_2.

Witten:2012:SDR

[96] Ian Witten. Semantic document representation: Do it with wikification. Lecture Notes in Computer Science, 7608:17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-34109-0_3.

Abdullin:2012:CHD

[97] Artur Abdullin and Olfa Nasraoui. Clustering heterogeneous data with mutual semi-supervision. Lecture Notes in Computer Science, 7608: 18–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34109-0_4/.

Abeliuk:2012:CST

[98] Andrés Abeliuk and Gonzalo Navarro. Compressed suffix trees for repetitive texts. Lecture Notes in Computer Science, 7608:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34109-0_5/.

Amir:2012:CMS

[99] Amihood Amir, Haim Paryenty, and Liam Roditty. Configurations and minority in the string consensus problem. Lecture Notes in Computer Science, 7608:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34109-0_6/.

An:2012:SNE

[100] Xiangdong An, Nick Cercone, and Hai Wang. A study on novelty evaluation in biomedical information retrieval. Lecture Notes in Computer Science, 7608:54–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34109-0_7/.

Badkobeh:2012:CME

[101] Golnaz Badkobeh and Maxime Crochemore. Computing the maximalexponent repeats of an overlap-free string in linear time. Lecture Notes in Computer Science, 7608:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34109-0_8/.

Balog:2012:CRS

[102] Krisztian Balog and Robert Neumayer. Collection ranking and selection for federated entity search. Lecture Notes in Computer Science, 7608: 73–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34109-0_9/.

Bannai:2012:ELF

[103] Hideo Bannai and Shunsuke Inenaga. Efficient LZ78 factorization of grammar compressed text. Lecture Notes in Computer Science, 7608: 86–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34109-0_10/.

Beller:2012:SEC

[104] Timo Beller and Katharina Berger. Space-efficient computation of maximal and supermaximal repeats in genome sequences. Lecture Notes in Computer Science, 7608:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34109-0_11/.

Jabeur:2012:AMI

[105] Lamjed Ben Jabeur and Lynda Tamine. Active microbloggers: Identifying influencers, leaders and discussers in microblogging networks. Lecture Notes in Computer Science, 7608:111–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34109-0_12/.

Birmele:2012:EBE

[106] Etienne Birmelé and Pierluigi Crescenzi. Efficient bubble enumeration in directed graphs. Lecture Notes in Computer Science, 7608: 118–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34109-0_13/.

Blin:2012:LCS

[107] Guillaume Blin and Minghui Jiang. The longest common subsequence problem with crossing-free arc-annotated sequences. Lecture Notes in Computer Science, 7608:130–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34109-0_14/.

Bravo-Marquez:2012:ZLD

[108] Felipe Bravo-Marquez and Manuel Manriquez. A Zipf-like distant supervision approach for multi-document summarization using Wikinews articles. Lecture Notes in Computer Science, 7608:143–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34109-0_15/.

Brisaboa:2012:RDR

[109] Nieves R. Brisaboa and Ana Cerdeira-Pena. Ranked document retrieval in (almost) no space. Lecture Notes in Computer Science, 7608: 155–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34109-0_16/.

Cambazoglu:2012:IRP

[110] B. Barla Cambazoglu. Impact of regionalization on performance of Web search engine result caches. Lecture Notes in Computer Science, 7608: 161–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34109-0_17/.

Claude:2012:WM

[111] Francisco Claude and Gonzalo Navarro. The wavelet matrix. Lecture Notes in Computer Science, 7608:167–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34109-0_18/.

Claude:2012:IGB

[112] Francisco Claude and Gonzalo Navarro. Improved grammar-based compressed indexes. Lecture Notes in Computer Science, 7608:180–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34109-0_19/.

deGroc:2012:EPR

[113] Clément de Groc and Xavier Tannier. Experiments on pseudo relevance feedback using graph random walks. Lecture Notes in Computer Science, 7608:193–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34109-0_20/.

Anonymous:2012:FMf

[114] Anonymous. Front matter. Lecture Notes in Computer Science, 7608: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34109-0/1.

Beyer:2012:LDV

[115] Dirk Beyer and Alexander K. Petrenko. Linux driver verification. Lecture Notes in Computer Science, 7610:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34032-1_1/.

Kok:2012:BDP

[116] Joost Kok, Anna-Lena Lamprecht, Fons J. Verbeek, and Mark D. Wilkinson. Bioscientific data processing and modeling. Lecture Notes in Computer Science, 7610:7–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_2/.

Sanchez-Faddeev:2012:UMO

[117] Hernando Sanchez-Faddeev, Michael T. M. Emmerich, Fons J. Verbeek, and Andrew H. Henry. Using multiobjective optimization and energy minimization to design an isoform-selective ligand of the 14-3-3 protein. Lecture Notes in Computer Science, 7610:12–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34032-1_3/.

Yan:2012:SHT

[118] Kuan Yan and Fons J. Verbeek. Segmentation for high-throughput image analysis: Watershed masked clustering. Lecture Notes in Computer Science, 7610:25–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34032-1_4/.

Nezhinsky:2012:ERS

[119] Alexander E. Nezhinsky and Fons J. Verbeek. Efficient and robust shape retrieval from deformable templates. Lecture Notes in Computer Science, 7610:42–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34032-1_5/.

Wood:2012:ODD

[120] Ian Wood, Ben Vandervalk, Luke McCarthy, and Mark D. Wilkinson. OWL-DL domain-models as abstract workflows. Lecture Notes in Computer Science, 7610:56–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_6/.

vonReinersdorff:2012:PDI

[121] Andrea Braun von Reinersdorff, Tiziana Margaria, and Christoph Rasche. Processes and data integration in the networked healthcare. Lecture Notes in Computer Science, 7610:67–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_7/.

Margaria:2012:SME

[122] Tiziana Margaria, Steve Boßelmann, and Bertold Kujath. Simple modeling of executable role-based workflows: An application in the healthcare domain. Lecture Notes in Computer Science, 7610:70–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34032-1_8/.

Holubek:2012:CHA

[123] Andreas Holubek and Christian Metzger. Considerations for healthcare applications in a platform as a service environment. Lecture Notes in Computer Science, 7610:73–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-34032-1_9.

Doedt:2012:RSC

[124] Markus Doedt, Thomas Göke, Jan Pardo, and Bernhard Steffen. Reha-Sports: The challenge of small margin healthcare accounting. Lecture Notes in Computer Science, 7610:75–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34032-1_10/.

Lisper:2012:TCT

[125] Björn Lisper, Johan Nordlander, and Sophie Quinton. Timing constraints: Theory meets practice. Lecture Notes in Computer Science, 7610:78–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34032-1_11.

Lisper:2012:SFT

[126] Björn Lisper and Johan Nordlander. A simple and flexible timing constraint logic. Lecture Notes in Computer Science, 7610:80–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34032-1_12/.

Quinton:2012:GWH

[127] Sophie Quinton and Rolf Ernst. Generalized weakly-hard constraints. Lecture Notes in Computer Science, 7610:96–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34032-1_13/.

Peraldi-Frati:2012:MBA

[128] Marie-Agnès Peraldi-Frati, Arda Goknil, Morayo Adedjouma, and Pierre Yves Gueguen. Modeling a BSG-E automotive system with the timing augmented description language. Lecture Notes in Computer Science, 7610:111–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34032-1_14/.

Ouranos:2012:FAT

[129] Iakovos Ouranos, Kazuhiro Ogata, and Petros Stefaneas. Formal analysis of TESLA protocol in the timed OTS/CafeOBJ method. Lecture Notes in Computer Science, 7610:126–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_15/.

Ge:2012:FSV

[130] Ning Ge, Marc Pantel, and Xavier Crégut. Formal specification and verification of task time constraints for real-time systems. Lecture Notes in Computer Science, 7610:143–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_16/.

Kirner:2012:WAT

[131] Raimund Kirner. The WCET Analysis Tool CalcWcet167. Lecture Notes in Computer Science, 7610:158–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_17/.

Birken:2012:AEE

[132] Klaus Birken. Abstract execution for event-driven systems — an application from automotive/infotainment development. Lecture Notes in Computer Science, 7610:173–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34032-1_18/.

Fantechi:2012:FMI

[133] Alessandro Fantechi, Francesco Flammini, and Stefania Gnesi. Formal methods for intelligent transportation systems. Lecture Notes in Computer Science, 7610:187–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34032-1_19/.

Flammini:2012:MDV

[134] Francesco Flammini, Stefano Marrone, Nicola Mazzocca, and Roberto Nardone. Model-driven V&V processes for computer based control systems: a unifying perspective. Lecture Notes in Computer Science, 7610: 190–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34032-1_20/.

Anonymous:2012:FMg

[135] Anonymous. Front matter. Lecture Notes in Computer Science, 7610: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34032-1/1.

Afek:2012:CPC

[136] Yehuda Afek, Haim Kaplan, Boris Korenfeld, Adam Morrison, and Robert E. Tarjan. CBTree: a practical concurrent self-adjusting search tree. Lecture Notes in Computer Science, 7611:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33651-5_1/.

Ellen:2012:EFI

[137] Faith Ellen, Vijaya Ramachandran, and Philipp Woelfel. Efficient fetch-and-increment. Lecture Notes in Computer Science, 7611:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33651-5_2/.

Gotsman:2012:SNW

[138] Alexey Gotsman, Madanlal Musuvathi, and Hongseok Yang. Show no weakness: Sequentially consistent specifications of TSO libraries. Lecture Notes in Computer Science, 7611:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33651-5_3/.

Anaya:2012:CIP

[139] Julian Anaya, Jérémie Chalopin, Jurek Czyzowicz, Arnaud Labourel, and Andrzej Pelc. Collecting information by power-aware mobile agents. Lecture Notes in Computer Science, 7611:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33651-5_4/.

Feinerman:2012:MLB

[140] Ofer Feinerman and Amos Korman. Memory lower bounds for randomized collaborative search and implications for biology. Lecture Notes in Computer Science, 7611:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33651-5_5/.

Chen:2012:GAP

[141] Chen Chen, Roman Vitenberg, and Hans-Arno Jacobsen. A generalized algorithm for publish/subscribe overlay design and its fast implementation. Lecture Notes in Computer Science, 7611:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33651-5_6/.

Censor-Hillel:2012:BCC

[142] Keren Censor-Hillel, Bernhard Haeupler, Nancy Lynch, and Muriel Médard. Bounded-contention coding for wireless networks in the high SNR regime. Lecture Notes in Computer Science, 7611:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33651-5_7/.

Jurdzinski:2012:DBS

[143] Tomasz Jurdzinski and Dariusz R. Kowalski. Distributed backbone structure for algorithms in the SINR model of wireless networks. Lecture Notes in Computer Science, 7611:106–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33651-5_8/.

Bienkowski:2012:DOS

[144] Marcin Bienkowski, Tomasz Jurdzinski, Miroslaw Korzeniowski, and Dariusz R. Kowalski. Distributed online and stochastic queuing on a multiple access channel. Lecture Notes in Computer Science, 7611:121– 135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33651-5_9/.

Sarma:2012:FDC

[145] Atish Das Sarma, Anisur Rahaman Molla, and Gopal Pandurangan. Fast distributed computation in dynamic networks via random walks. Lecture Notes in Computer Science, 7611:136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33651-5_10/.

Sarma:2012:DSD

[146] Atish Das Sarma, Ashwin Lall, Danupon Nanongkai, and Amitabh Trehan. Dense subgraphs on dynamic networks. Lecture Notes in Computer Science, 7611:151–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33651-5_11/.

Haeupler:2012:LBI

[147] Bernhard Haeupler and Fabian Kuhn. Lower bounds on information dissemination in dynamic networks. Lecture Notes in Computer Science, 7611:166–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33651-5_12/.

Goos:2012:NST

[148] Mika Göös and Jukka Suomela. No sublogarithmic-time approximation scheme for bipartite vertex cover. Lecture Notes in Computer Science, 7611:181–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33651-5_13/.

Dolev:2012:TTA

[149] Danny Dolev, Christoph Lenzen, and Shir Peled. “tri, tri again”: Finding triangles and small subgraphs in a distributed setting. Lecture Notes in Computer Science, 7611:195–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33651-5_14/.

Czygrinow:2012:DAA

[150] Andrzej Czygrinow, Michal Hanćkowiak, Edyta Szymańska, and Wojciech Wawrzyniak. Distributed 2-approximation algorithm for the semimatching problem. Lecture Notes in Computer Science, 7611:210–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33651-5_15/.

Ghaffari:2012:BCM

[151] Mohsen Ghaffari, Bernhard Haeupler, Nancy Lynch, and Calvin Newport. Bounds on contention management in radio networks. Lecture Notes in Computer Science, 7611:223–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33651-5_16/.

Daum:2012:ESB

[152] Sebastian Daum, Fabian Kuhn, and Calvin Newport. Efficient symmetry breaking in multi-channel radio networks. Lecture Notes in Computer Science, 7611:238–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33651-5_17/.

Maurer:2012:BBL

[153] Alexandre Maurer and Sébastien Tixeuil. On Byzantine broadcast in loosely connected networks. Lecture Notes in Computer Science, 7611: 253–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33651-5_18/.

Pareek:2012:RER

[154] Abhijeet Pareek and Philipp Woelfel. RMR-efficient randomized abortable mutual exclusion. Lecture Notes in Computer Science, 7611:267–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33651-5_19/.

Jayanti:2012:ARW

[155] Prasad Jayanti and Zhiyu Liu. Abortable reader-writer locks are no more complex than abortable mutex locks. Lecture Notes in Computer Science, 7611:282–296, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33651-5_20/.

Anonymous:2012:FMh

[156] Anonymous. Front matter. Lecture Notes in Computer Science, 7611: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33651-5/1.

Denney:2012:LMS

[157] Ewen Denney and Ganesh Pai. A lightweight methodology for safety case assembly. Lecture Notes in Computer Science, 7612:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33678-2_1/.

Hauge:2012:PBM

[158] André Alexandersen Hauge and Ketil Stølen. A pattern-based method for safe control systems exemplified within nuclear power production. Lecture Notes in Computer Science, 7612:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33678-2_2/.

Casals:2012:RAA

[159] Silvia Gil Casals, Philippe Owezarski, and Gilles Descargues. Risk assessment for airworthiness security. Lecture Notes in Computer Science, 7612:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33678-2_3/.

Lasche:2012:MGH

[160] Christoph Läsche, Eckard Böde, and Thomas Peikenkamp. A method for guided hazard identification and risk mitigation for offshore operations. Lecture Notes in Computer Science, 7612:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33678-2_4/.

Schafer:2012:RAS

[161] Manfred Schäfer. Risk analysis and software integrity protection for 4G network elements in ASMONIA. Lecture Notes in Computer Science, 7612:49–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33678-2_5/.

Woskowski:2012:AIS

[162] Christoph Woskowski. Applying industrial-strength testing techniques to critical care medical equipment. Lecture Notes in Computer Science, 7612:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33678-2_6/.

Izosimov:2012:RDT

[163] Viacheslav Izosimov, Urban Ingelsson, and Andreas Wallin. Requirement decomposition and testability in development of safety-critical automotive components. Lecture Notes in Computer Science, 7612:74– 86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33678-2_7/.

Krause:2012:MBS

[164] Jan Krause, Elke Hintze, Stephan Magnus, and Christian Diedrich. Model based specification, verification, and test generation for a safety Fieldbus profile. Lecture Notes in Computer Science, 7612:87–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33678-2_8/.

Edifor:2012:QPG

[165] Ernest Edifor, Martin Walker, and Neil Gordon. Quantification of priority-OR gates in temporal fault trees. Lecture Notes in Computer Science, 7612:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33678-2_9/.

Glass:2012:CLC

[166] Michael Glaß, Heng Yu, Felix Reimann, and Jürgen Teich. Cross-level compositional reliability analysis for embedded systems. Lecture Notes in Computer Science, 7612:111–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33678-2_10/.

Hoppe:2012:IFA

[167] Tobias Hoppe, Sven Kuhlmann, Stefan Kiltz, and Jana Dittmann. ITforensic automotive investigations on the example of route reconstruction on automotive system and communication data. Lecture Notes in Computer Science, 7612:125–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33678-2_11/.

Bock:2012:TIS

[168] Hans-Hermann Bock, Jens Braband, Birgit Milius, and Hendrik Schäbe. Towards an IT security protection profile for safety-related communication in railway automation. Lecture Notes in Computer Science, 7612: 137–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33678-2_12/.

Wieczorek:2012:TSF

[169] Felix Wieczorek, Christoph Krauß, Frank Schiller, and Claudia Eckert. Towards secure Fieldbus communication. Lecture Notes in Computer Science, 7612:149–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33678-2_13/.

Zakonov:2012:EEW

[170] Andrey Zakonov and Anatoly Shalyto. Extracting EFSMs of Web applications for formal requirements specification. Lecture Notes in Computer Science, 7612:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33678-2_14/.

Bicchierai:2012:OAS

[171] Irene Bicchierai, Giacomo Bucci, Carlo Nocentini, and Enrico Vicario. An ontological approach to systematization of SW-FMEA. Lecture Notes in Computer Science, 7612:173–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33678-2_15/.

Baldoni:2012:OBB

[172] Roberto Baldoni, Giorgia Lodi, Luca Montanari, Guido Mariotta, and Marco Rizzuto. Online black-box failure prediction for mission critical distributed systems. Lecture Notes in Computer Science, 7612: 185–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33678-2_16/.

DiLeo:2012:IHF

[173] Domenico Di Leo, Fatemeh Ayatolahi, Behrooz Sangchoolie, and Johan Karlsson. On the impact of hardware faults — an investigation of the relationship between workload inputs and failure mode distributions. Lecture Notes in Computer Science, 7612:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33678-2_17/.

Tarasyuk:2012:FDA

[174] Anton Tarasyuk, Inna Pereverzeva, Elena Troubitsyna, Timo Latvala, and Laura Nummila. Formal development and assessment of a reconfigurable on-board satellite system. Lecture Notes in Computer Science, 7612:210–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33678-2_18/.

Hannius:2012:ISE

[175] Olof Hannius and Johan Karlsson. Impact of soft errors in a jet engine controller. Lecture Notes in Computer Science, 7612:223–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33678-2_19/.

Pintard:2012:WAW

[176] Ludovic Pintard, Christel Seguin, and Jean-Paul Blanquart. Which automata for which safety assessment step of satellite FDIR? Lecture Notes in Computer Science, 7612:235–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33678-2_20/.

Anonymous:2012:FMi

[177] Anonymous. Front matter. Lecture Notes in Computer Science, 7612: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33678-2/1.

Anonymous:2012:FMj

[178] Anonymous. Front matter. Lecture Notes in Computer Science, 7613: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33675-1/1/1.

Ruiz:2012:IS

[179] Alejandra Ruiz, Tim P. Kelly, Mehrdad Sabetzadeh, and Didier Van Den Abeele. Introduction to sassur 2012. Lecture Notes in Computer Science, 7613:3–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_1/.

Denney:2012:AAC

[180] Ewen Denney, Ganesh Pai, and Josef Pohl. AdvoCATE: An assurance case automation toolset. Lecture Notes in Computer Science, 7613: 8–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_2/.

Ruiz:2012:TCB

[181] Alejandra Ruiz, Ibrahim Habli, and Huáscar Espinoza. Towards a casebased reasoning approach for safety assurance reuse. Lecture Notes in Computer Science, 7613:22–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33675-1_3/.

Markovski:2012:MSS

[182] Jasen Markovski and J. M. van de Mortel-Fronczak. Modeling for safety in a synthesis-centric systems engineering framework. Lecture Notes in Computer Science, 7613:36–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33675-1_4/.

Belmonte:2012:MBA

[183] Fabien Belmonte and Elie Soubiran. A model based approach for safety analysis. Lecture Notes in Computer Science, 7613:50–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33675-1_5/.

delaVara:2012:TMB

[184] Jose Luis de la Vara, Sunil Nair, Eric Verhulst, Janusz Studzizba, and Piotr Pepek. Towards a model-based evolutionary chain of evidence for compliance with safety standards. Lecture Notes in Computer Science, 7613:64–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_6/.

Zhao:2012:NAA

[185] Xingyu Zhao, Dajian Zhang, Minyan Lu, and Fuping Zeng. A new approach to assessment of confidence in assurance cases. Lecture Notes in Computer Science, 7613:79–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33675-1_7/.

Verhulst:2012:UMM

[186] Eric Verhulst and Bernhard H. C. Sputh. An unified meta-model for trustworthy systems engineering. Lecture Notes in Computer Science, 7613:92–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_8/.

Lanzaro:2012:PFI

[187] Anna Lanzaro, Antonio Pecchia, Marcello Cinque, Domenico Cotroneo, and Ricardo Barbosa. A preliminary fault injection framework for evaluating multicore systems. Lecture Notes in Computer Science, 7613: 106–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_9/.

Kastner:2012:MRT

[188] Daniel Kästner, Marc Schlickling, Markus Pister, Christoph Cullmann, and Gernot Gebhard. Meeting real-time requirements with multi-core processors. Lecture Notes in Computer Science, 7613:117–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33675-1_10/.

Graydon:2012:ASI

[189] Patrick J. Graydon and Tim P. Kelly. Assessing software interference management when modifying safety-related software. Lecture Notes in Computer Science, 7613:132–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33675-1_11/.

Anonymous:2012:FMk

[190] Anonymous. Front matter. Lecture Notes in Computer Science, 7613: 147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-33675-1/2/1.

Casimiro:2012:IA

[191] António Casimiro and Jörg Kaiser. Introduction to ASCoMS 2012. Lecture Notes in Computer Science, 7613:149–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-33675-1_12.

Marques:2012:TDS

[192] Luís Marques and António Casimiro. Towards dependable and stable perception in smart environments with timing and value faults. Lecture Notes in Computer Science, 7613:151–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33675-1_13/.

Zug:2012:ASF

[193] Sebastian Zug, Tino Brade, Jörg Kaiser, and Sasanka Potluri. An approach supporting fault-propagation analysis for smart sensor systems. Lecture Notes in Computer Science, 7613:162–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33675-1_14/.

Ostberg:2012:UQM

[194] Kenneth Östberg and Rolf Johansson. Use of quality metrics for functional safety in systems of cooperative vehicles. Lecture Notes in Computer Science, 7613:174–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33675-1_15/.

Berger:2012:AVS

[195] Christian Berger. From autonomous vehicles to safer cars: Selected challenges for the software engineering. Lecture Notes in Computer Science, 7613:180–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33675-1_16/.

Ljungkrantz:2012:MSR

[196] Oscar Ljungkrantz, Henrik Lönn, Hans Blom, Cecilia Ekelin, and Daniel Karlsson. Modelling of safety-related timing constraints for automotive embedded systems. Lecture Notes in Computer Science, 7613: 190–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_17/.

Anonymous:2012:FMl

[197] Anonymous. Front matter. Lecture Notes in Computer Science, 7613: 203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/content/pdf/bfm:978-
3-642-33675-1/3/1.

Esposito:2012:ID

[198] Christian Esposito, Marco Platania, and Francesco Brancati. Introduction to DESEC4LCCI 2012. Lecture Notes in Computer Science, 7613: 205–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_18/.

Montecchi:2012:QSE

[199] Leonardo Montecchi, Paolo Lollini, Andrea Bondavalli, and Ernesto La Mattina. Quantitative security evaluation of a multi-biometric authentication system. Lecture Notes in Computer Science, 7613:209–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33675-1_19/.

Romano:2012:PWZ

[200] Luigi Romano, Salvatore D’Antonio, Valerio Formicola, and Luigi Coppolino. Protecting the WSN zones of a critical infrastructure via enhanced SIEM technology. Lecture Notes in Computer Science, 7613: 222–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33675-1_20/.

Anonymous:2012:FMm

[201] Anonymous. Front matter. Lecture Notes in Computer Science, 7613: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33675-1/1.

Aikawa:2012:ICP

[202] Takako Aikawa, Kentaro Yamamoto, and Hitoshi Isahara. The impact of crowdsourcing post-editing with the collaborative translation framework. Lecture Notes in Computer Science, 7614:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33983-7_1/.

Chen:2012:TQJ

[203] Shaoyu Chen and Tadahiro Matsumoto. Translation of quantifiers in Japanese–Chinese machine translation. Lecture Notes in Computer Science, 7614:11–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33983-7_2/.

Isahara:2012:TPU

[204] Hitoshi Isahara. Toward practical use of machine translation. Lecture Notes in Computer Science, 7614:23–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33983-7_3/.

Sakata:2012:PLP

[205] Jun Sakata, Masato Tokuhisa, and Jin’ichi Murakami. Phrase-level pattern-based machine translation based on analogical mapping method. Lecture Notes in Computer Science, 7614:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33983-7_4/.

Afli:2012:PTE

[206] Haithem Afli, Loïc Barrault, and Holger Schwenk. Parallel texts extraction from multimodal comparable corpora. Lecture Notes in Computer Science, 7614:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33983-7_5/.

Jin:2012:RCS

[207] Gan Jin and Natallia Khatseyeva. A reliable communication system to maximize the communication quality. Lecture Notes in Computer Science, 7614:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33983-7_6/.

Lejeune:2012:DLI

[208] Gaël Lejeune, Romain Brixtel, Antoine Doucet, and Nadine Lucas. DAnIEL: Language independent character-based news surveillance. Lecture Notes in Computer Science, 7614:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33983-7_7/.

Qu:2012:OTT

[209] Jian Qu, Akira Shimazu, and Minh Le Nguyen. OOV term translation, context information and definition extraction based on OOV term type prediction. Lecture Notes in Computer Science, 7614:76– 87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33983-7_8/.

Sadat:2012:EWB

[210] Fatiha Sadat. Exploiting a Web-based encyclopedia as a knowledge base for the extraction of multilingual terminology. Lecture Notes in Computer Science, 7614:88–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33983-7_9/.

Srivastava:2012:SLS

[211] Jyoti Srivastava and Sudip Sanyal. Segmenting long sentence pairs to improve word alignment in English–Hindi parallel corpora. Lecture Notes in Computer Science, 7614:97–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33983-7_10/.

Vuong:2012:SSP

[212] Hoai-Thu Vuong, Dao Ngoc Tu, Minh Le Nguyen, and Vinh Van Nguyen. Shallow syntactic preprocessing for statistical machine translation. Lecture Notes in Computer Science, 7614:108–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33983-7_11/.

Njonko:2012:LRB

[213] Paul Brillant Feuto Njonko, Sylviane Cardey-Greenfield, and Peter Greenfield. Linguistic rules based approach for automatic restoration of accents on French texts. Lecture Notes in Computer Science, 7614:118– 125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33983-7_12/.

Ghayoomi:2012:WCP

[214] Masood Ghayoomi. Word clustering for Persian statistical parsing. Lecture Notes in Computer Science, 7614:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33983-7_13/.

Kampeera:2012:BLS

[215] Wannachai Kampeera and Sylviane Cardey-Greenfield. Building a lexically and semantically-rich resource for paraphrase processing. Lecture Notes in Computer Science, 7614:138–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33983-7_14/.

Zaborowski:2012:TCD

[216] Bartosz Zaborowski and Adam Przepiórkowski. Tagset conversion with decision trees. Lecture Notes in Computer Science, 7614:144–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33983-7_15/.

Zimina:2012:FRP

[217] Elizaveta Zimina. Fitting a round peg in a square hole: Japanese resource grammar in GF. Lecture Notes in Computer Science, 7614: 156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33983-7_16/.

Aliwy:2012:ALA

[218] Ahmed H. Aliwy. Arabic language analyzer with lemma extraction and rich tagset. Lecture Notes in Computer Science, 7614:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33983-7_17/.

Flores:2012:TRM

[219] Jorge J. García Flores, Pierre Zweigenbaum, Zhao Yue, and William Turner. Tracking researcher mobility on the Web using snippet semantic analysis. Lecture Notes in Computer Science, 7614:180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33983-7_18/.

Goluchowski:2012:SRL

[220] Konrad Gołuchowski and Adam Przepiórkowski. Semantic role labelling without deep syntactic parsing. Lecture Notes in Computer Science, 7614: 192–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33983-7_19/.

Jarzebowski:2012:TIE

[221] Przemysław Jarzebowski and Adam Przepiórkowski. Temporal information extraction with cross-language projected data. Lecture Notes in Computer Science, 7614:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33983-7_20/.

Anonymous:2012:FMn

[222] Anonymous. Front matter. Lecture Notes in Computer Science, 7614: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33983-7/1.

Kersten:2012:IBL

[223] Thomas P. Kersten and Maren Lindstaedt. Image-based low-cost systems for automatic 3D recording and modelling of archaeological finds and objects. Lecture Notes in Computer Science, 7616:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34234-9_1/.

Fritsch:2012:PPC

[224] Dieter Fritsch, Mohammed Abdel-Wahab, Alessandro Cefalu, and Konrad Wenzel. Photogrammetric point cloud collection with multi-camera systems. Lecture Notes in Computer Science, 7616:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34234-9_2/.

Trinks:2012:PAL

[225] Immo Trinks, Wolfgang Neubauer, and Michael Doneus. Prospecting archaeological landscapes. Lecture Notes in Computer Science, 7616: 21–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34234-9_3/.

Hollaus:2012:MIA

[226] Fabian Hollaus, Melanie Gau, and Robert Sablatnig. Multispectral image acquisition of ancient manuscripts. Lecture Notes in Computer Science, 7616:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34234-9_4/.

Remondino:2012:LCO

[227] Fabio Remondino, Silvio Del Pizzo, Thomas P. Kersten, and Salvatore Troisi. Low-cost and open-source solutions for automated image orientation — a critical overview. Lecture Notes in Computer Science, 7616: 40–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34234-9_5/.

Kersten:2012:AGH

[228] Thomas P. Kersten, Friedrich Keller, Jerome Saenger, and Jochen Schiewe. Automated generation of an historic 4D city model of Hamburg and its visualisation with the GE Engine. Lecture Notes in Computer Science, 7616:55–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34234-9_6/.

Balzani:2012:ABC

[229] Marcello Balzani, Carlo Bughi, Federico Ferrari, Luca Rossato, and Alessandra Tursi. Alberti’s box: The cultural multimedia project on the architectures of Leon Battista Alberti. Lecture Notes in Computer Science, 7616:66–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34234-9_7/.

Moussa:2012:AFD

[230] Wassim Moussa, Mohammed Abdel-Wahab, and Dieter Fritsch. Automatic fusion of digital images and laser scanner data for heritage preservation. Lecture Notes in Computer Science, 7616:76–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34234-9_8/.

Bernikola:2012:MDS

[231] Eirini Bernikola, Elsa Tsiranidou, Vivi Tornari, Vincent Detalle, and Jean Luc Bodnar. Mapping of defect structural micro-morphology in the documentation of conservation approaches. Lecture Notes in Computer Science, 7616:86–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34234-9_9/.

Sidiropoulou-Velidou:2012:ETI

[232] Dafni Sidiropoulou-Velidou, Andreas Georgopoulos, and José Luis Lerma. Exploitation of thermal imagery for the detection of pathologies in monuments. Lecture Notes in Computer Science, 7616:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34234-9_10/.

Doulamis:2012:LSL

[233] Nikolaos Doulamis, Christos Yiakoumettis, and George Miaoulis. On-line spectral learning in exploring 3D large scale geo-referred scenes. Lecture Notes in Computer Science, 7616:109–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34234-9_11/.

Drap:2012:ISM

[234] Pierre Drap, Djamal Merad, Julien Seinturier, Jean-Marc Boï, and Daniela Peloso. An information system for medieval archaeology based on photogrammetry and archaeological database: The Shawbak Castle Project. Lecture Notes in Computer Science, 7616:119–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34234-9_12/.

Agapiou:2012:TAI

[235] Athos Agapiou, Diofantos G. Hadjimitsis, Andreas Georgopoulos, and Apostolos Sarris. Towards an archaeological index: Identification of the spectral regions of stress vegetation due to buried archaeological remains. Lecture Notes in Computer Science, 7616:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34234-9_13/.

Gigliarelli:2012:HSH

[236] Elena Gigliarelli, Donato Carlea, Angela Corcella, and Heleni Porfyriou. Historical social housing: Smart analysis and design for conservation and energy regeneration. Lecture Notes in Computer Science, 7616: 139–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34234-9_14/.

Echavarria:2012:HMD

[237] Karina Rodriguez Echavarria, Jaime Kaminski, and David Arnold. 3D heritage on mobile devices: Scenarios and opportunities. Lecture Notes in Computer Science, 7616:149–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34234-9_15/.

Fleury:2012:RAR

[238] Philippe Fleury and Sophie Madeleine. Reviving Ancient Rome: Virtual reality at the service of cultural heritage. Lecture Notes in Computer Science, 7616:159–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34234-9_16/.

Zotti:2012:VRD

[239] Georg Zotti and Wolfgang Neubauer. Virtual reconstructions in a desktop planetarium for demonstrations in cultural astronomy. Lecture Notes in Computer Science, 7616:170–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34234-9_17/.

Fernandez-Palacios:2012:ARA

[240] Belén Jiménez Fernández-Palacios, Alessandro Rizzi, and Francesco Nex. Augmented reality for archaeological finds. Lecture Notes in Computer Science, 7616:181–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34234-9_18/.

Rahaman:2012:VHE

[241] Hafizur Rahaman, Rana Das, and Shehzad Zahir. Virtual heritage: Exploring photorealism. Lecture Notes in Computer Science, 7616:191– 200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34234-9_19/.

Kersten:2012:VAM

[242] Thomas P. Kersten and Maren Lindstaedt. Virtual architectural 3D model of the Imperial Cathedral (Kaiserdom) of Königslutter, Germany through terrestrial laser scanning. Lecture Notes in Computer Science, 7616: 201–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34234-9_20/.

Anonymous:2012:FMo

[243] Anonymous. Front matter. Lecture Notes in Computer Science, 7616: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34234-9/1.

Blom:2012:DFU

[244] Arjan Blom, Gerhard de Koning Gans, Erik Poll, Joeri de Ruiter, and Roel Verdult. Designed to fail: a USB-connected reader for online banking. Lecture Notes in Computer Science, 7617:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34210-3_1/.

Adolphi:2012:SAO

[245] Benjamin Adolphi and Hanno Langweg. Security add-ons for mobile platforms. Lecture Notes in Computer Science, 7617:17–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34210-3_2/.

Jensen:2012:TAV

[246] Torben Jensen, Heine Pedersen, Mads Chr. Olesen, and René Rydhof Hansen. THAPS: Automated vulnerability scanning of PHP applications. Lecture Notes in Computer Science, 7617:31–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34210-3_3/.

Sommestad:2012:CSE

[247] Teodor Sommestad and Jonas Hallberg. Cyber security exercises and competitions as a platform for cyber security experiments. Lecture Notes in Computer Science, 7617:47–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34210-3_4/.

Proctor:2012:DCS

[248] Tony Proctor. The development of cyber security warning, advice and report points. Lecture Notes in Computer Science, 7617:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34210-3_5/.

Groner:2012:TEE

[249] Ramona Groner and Philipp Brune. Towards an empirical examination of IT security infrastructures in SME. Lecture Notes in Computer Science, 7617:73–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34210-3_6/.

Massacci:2012:HSS

[250] Fabio Massacci and Federica Paci. How to select a security requirements method? A comparative study with students and practitioners. Lecture Notes in Computer Science, 7617:89–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34210-3_7/.

Noorman:2012:TSN

[251] Job Noorman, Nick Nikiforakis, and Frank Piessens. There is safety in numbers: Preventing control-flow hijacking by duplication. Lecture Notes in Computer Science, 7617:105–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34210-3_8/.

Milushev:2012:CUS

[252] Dimiter Milushev and Dave Clarke. Coinductive unwinding of securityrelevant hyperproperties. Lecture Notes in Computer Science, 7617: 121–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34210-3_9/.

Staahl:2012:RSS

[253] Björn Ståhl and Per Mellstrand. Retooling and securing systemic debugging. Lecture Notes in Computer Science, 7617:137–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34210-3_10/.

Helkala:2012:CAP

[254] Kirsi Helkala, Nils Kalstad Svendsen, Per Thorsheim, and Anders Wiehe. Cracking associative passwords. Lecture Notes in Computer Science, 7617: 153–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34210-3_11/.

Zibuschka:2012:HAH

[255] Jan Zibuschka and Lothar Fritsch. A hybrid approach for highly available and secure storage of pseudo-SSO credentials. Lecture Notes in Computer Science, 7617:169–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34210-3_12/.

Sangchoolie:2012:AQP

[256] Behrooz Sangchoolie, Mazdak Rajabi Nasab, Tomas Olovsson, and Wolfgang John. Assessing the quality of packet-level traces collected on Internet backbone links. Lecture Notes in Computer Science, 7617: 184–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34210-3_13/.

Clark:2012:EKS

[257] Jason W. Clark. Everything but the kitchen sink: Determining the effect of multiple attacks on privacy preserving technology users. Lecture Notes in Computer Science, 7617:199–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34210-3_14/.

Duan:2012:CWI

[258] Huiying Duan and Cäcilia Zirn. Can we identify manipulative behavior and the corresponding suspects on review Websites using supervised learning? Lecture Notes in Computer Science, 7617:215–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34210-3_15/.

Pulls:2012:PFC

[259] Tobias Pulls. Privacy-friendly cloud storage for the data track. Lecture Notes in Computer Science, 7617:231–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34210-3_16/.

Anonymous:2012:BMc

[260] Anonymous. Back matter. Lecture Notes in Computer Science, 7617: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34210-3/1.

Anonymous:2012:FMp

[261] Anonymous. Front matter. Lecture Notes in Computer Science, 7617: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34210-3/1.

Djebbar:2012:ASB

[262] Fatiha Djebbar and Beghdad Ayad. Audio steganalysis based on lossless data-compression techniques. Lecture Notes in Computer Science, 7618:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34129-8_1/.

Chow:2012:EPV

[263] Yang-Wai Chow, Willy Susilo, and Duncan S. Wong. Enhancing the perceived visual quality of a size invariant visual cryptography scheme. Lecture Notes in Computer Science, 7618:10–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34129-8_2/.

Ganan:2012:IRS

[264] Carlos Gañán, Jose L. Muñoz, Oscar Esparza, Jorge Mata-Díaz, and Juanjo Alins. Impact of the revocation service in PKI prices. Lecture Notes in Computer Science, 7618:22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34129-8_3/.

Bahig:2012:CMP

[265] Hatem M. Bahig, Ashraf Bhery, and Dieaa I. Nassr. Cryptanalysis of multi-prime RSA with small prime difference. Lecture Notes in Computer Science, 7618:33–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34129-8_4/.

Xu:2012:IPR

[266] Jun Xu, Lei Hu, and Siwei Sun. Implicit polynomial recovery and cryptanalysis of a combinatorial key cryptosystem. Lecture Notes in Computer Science, 7618:45–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34129-8_5/.

Liu:2012:IRK

[267] Shusheng Liu, Zheng Gong, and Libin Wang. Improved related-key differential attacks on reduced-round LBlock. Lecture Notes in Computer Science, 7618:58–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34129-8_6/.

Tang:2012:CAL

[268] Yajuan Tang. Countermeasures on application level low-rate denialof-service attack. Lecture Notes in Computer Science, 7618:70–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34129-8_7/.

Trabelsi:2012:FPF

[269] Zouheir Trabelsi, Liren Zhang, and Safaa Zeidan. Firewall packet filtering optimization using statistical traffic awareness test. Lecture Notes in Computer Science, 7618:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34129-8_8/.

Felix:2012:GBM

[270] John Felix, Charles Joseph, and Ali A. Ghorbani. Group behavior metrics for P2P botnet detection. Lecture Notes in Computer Science, 7618: 93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34129-8_9/.

Ma:2012:HPO

[271] Yuan Ma, Luning Xia, Jingqiang Lin, Jiwu Jing, Zongbin Liu, and Xingjie Yu. Hardware performance optimization and evaluation of SM3 hash algorithm on FPGA. Lecture Notes in Computer Science, 7618: 105–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34129-8_10/.

Xiong:2012:CLR

[272] Hao Xiong, Cong Zhang, Tsz Hon Yuen, Echo P. Zhang, Siu Ming Yiu, and Sihan Qing. Continual leakage-resilient dynamic secret sharing in the split-state model. Lecture Notes in Computer Science, 7618: 119–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34129-8_11/.

Henecka:2012:CRN

[273] Wilko Henecka, Nigel Bean, and Matthew Roughan. Conversion of realnumbered privacy-preserving problems into the integer domain. Lecture Notes in Computer Science, 7618:131–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34129-8_12/.

Wang:2012:PAO

[274] Yang Wang, Man Ho Au, and Willy Susilo. Perfect ambiguous optimistic fair exchange. Lecture Notes in Computer Science, 7618:142– 153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34129-8_13/.

Pang:2012:PPN

[275] Xiaoqiong Pang, Bo Yang, and Qiong Huang. Privacy-preserving noisy keyword search in cloud computing. Lecture Notes in Computer Science, 7618:154–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34129-8_14/.

Yuen:2012:FSA

[276] Tsz Hon Yuen, Joseph K. Liu, Xinyi Huang, Man Ho Au, Willy Susilo, and Jianying Zhou. Forward secure attribute-based signatures. Lecture Notes in Computer Science, 7618:167–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34129-8_15/.

Ding:2012:CRP

[277] Ning Ding and Dawu Gu. On constant-round precise zero-knowledge. Lecture Notes in Computer Science, 7618:178–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34129-8_16/.

Li:2012:OEA

[278] Jingwei Li, Chunfu Jia, Jin Li, and Xiaofeng Chen. Outsourcing encryption of attribute-based encryption with MapReduce. Lecture Notes in Computer Science, 7618:191–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34129-8_17/.

Fujioka:2012:SEI

[279] Atsushi Fujioka, Taiichi Saito, and Keita Xagawa. Security enhancement of identity-based identification with reversibility. Lecture Notes in Computer Science, 7618:202–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34129-8_18/.

Akram:2012:CAS

[280] Raja Naeem Akram, Konstantinos Markantonakis, and Keith Mayes. Coopetitive architecture to support a dynamic and scalable NFC based mobile services architecture. Lecture Notes in Computer Science, 7618: 214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34129-8_19/.

Zhu:2012:PBA

[281] Jiawei Zhu, Zhi Guan, Yang Yang, Liangwen Yu, Huiping Sun, and Zhong Chen. Permission-based abnormal application detection for Android. Lecture Notes in Computer Science, 7618:228–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34129-8_20/.

Anonymous:2012:FMq

[282] Anonymous. Front matter. Lecture Notes in Computer Science, 7618: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34129-8/1.

Cawley:2012:FMS

[283] Gavin C. Cawley. Over-fitting in model selection and its avoidance. Lecture Notes in Computer Science, 7619:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-34156-4_1.

Sebastiani:2012:IDA

[284] Paola Sebastiani. Intelligent data analysis of human genetic data. Lecture Notes in Computer Science, 7619:2–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34156-4_2/.

Siebes:2012:QDA

[285] Arno Siebes. Queries for data analysis. Lecture Notes in Computer Science, 7619:7–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34156-4_3/.

Akbar:2012:PDM

[286] Zaenal Akbar, Violeta N. Ivanova, and Michael R. Berthold. Parallel data mining revisited. Better, not faster. Lecture Notes in Computer Science, 7619:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34156-4_4/.

deAmorim:2012:WFP

[287] Renato Cordeiro de Amorim and Trevor Fenner. Weighting features for partition around medoids using the Minkowski metric. Lecture Notes in Computer Science, 7619:35–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34156-4_5/.

deAmorim:2012:IMW

[288] Renato Cordeiro de Amorim and Peter Komisarczuk. On initializations for the Minkowski weighted K-means. Lecture Notes in Computer Science, 7619:45–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34156-4_6/.

Berk:2012:SNM

[289] Maurice Berk and Giovanni Montana. A skew-t-normal multi-level reduced-rank functional PCA model for the analysis of replicated genomics time course data. Lecture Notes in Computer Science, 7619: 56–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34156-4_7/.

Boulares:2012:MFS

[290] Mehrez Boulares and Mohamed Jemni. Methodological foundation for sign language 3D motion trajectory analysis. Lecture Notes in Computer Science, 7619:67–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34156-4_8/.

Braune:2012:ADC

[291] Christian Braune, Christian Borgelt, and Sonja Grün. Assembly detection in continuous neural Spike train data. Lecture Notes in Computer Science, 7619:78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34156-4_9/.

Carmona:2012:OTD

[292] Josep Carmona and Ricard Gavaldà. Online techniques for dealing with concept drift in process mining. Lecture Notes in Computer Science, 7619: 90–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34156-4_10/.

Chen:2012:EBC

[293] Jun Chen, Mahdi Mahfouf, Chris Bingham, Yu Zhang, Zhijing Yang, and Michael Gallimore. An evolutionary based clustering algorithm applied to dada compression for industrial systems. Lecture Notes in Computer Science, 7619:103–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34156-4_11/.

Duivesteijn:2012:MLL

[294] Wouter Duivesteijn, Eneldo Loza Mencía, Johannes Fürnkranz, and Arno Knobbe. Multi-label LeGo — enhancing multi-label classifiers with local patterns. Lecture Notes in Computer Science, 7619:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34156-4_12/.

Gisbrecht:2012:DDR

[295] Andrej Gisbrecht, Daniela Hofmann, and Barbara Hammer. Discriminative dimensionality reduction mappings. Lecture Notes in Computer Science, 7619:126–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34156-4_13/.

Jorge:2012:FIC

[296] Alípio M. Jorge, João Mendes-Moreira, Jorge Freire de Sousa, and Carlos Soares. Finding interesting contexts for explaining deviations in bus trip duration using distribution rules. Lecture Notes in Computer Science, 7619:139–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34156-4_14/.

Klawonn:2012:CFS

[297] Frank Klawonn, Nada Abidi, Evelin Berger, and Lothar Jänsch. Curve fitting for short time series data from high throughput experiments with correction for biological variation. Lecture Notes in Computer Science, 7619:150–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34156-4_15/.

Kontonasios:2012:FCP

[298] Kleanthis-Nikolaos Kontonasios and Tijl DeBie. Formalizing complex prior information to quantify subjective interestingness of frequent pattern sets. Lecture Notes in Computer Science, 7619:161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34156-4_16/.

Largeron:2012:MTS

[299] Christine Largeron, Christophe Moulin, and Mathias Géry. MCut: a thresholding strategy for multi-label classification. Lecture Notes in Computer Science, 7619:172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34156-4_17/.

vanLeeuwen:2012:ITR

[300] Matthijs van Leeuwen and Diyah Puspitaningrum. Improving tag recommendation using few associations. Lecture Notes in Computer Science, 7619:184–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34156-4_18/.

Malmi:2012:IAS

[301] Eric Malmi, Juha Raitio, Oskar Kohonen, Krista Lagus, and Timo Honkela. Identifying anomalous social contexts from mobile proximity data using binomial mixture models. Lecture Notes in Computer Science, 7619:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34156-4_19/.

Metivier:2012:CCU

[302] Jean-Philippe Métivier, Patrice Boizumault, Bruno Crémilleux, and Mehdi Khiari. Constrained clustering using SAT. Lecture Notes in Computer Science, 7619:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34156-4_20/.

Anonymous:2012:FMr

[303] Anonymous. Front matter. Lecture Notes in Computer Science, 7619: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34156-4/1.

Kitto:2012:QIM

[304] Kirsty Kitto, Fabio Boschetti, and Peter Bruza. The quantum inspired modelling of changing attitudes and self-organising societies. Lecture Notes in Computer Science, 7620:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35659-9_1/.

Dubois:2012:LAP

[305] François Dubois, Isabelle Greff, and Thomas Hélie. On least action principles for discrete quantum scales. Lecture Notes in Computer Science, 7620:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35659-9_2/.

Widdows:2012:RCB

[306] Dominic Widdows and Trevor Cohen. Real, complex, and binary semantic vectors. Lecture Notes in Computer Science, 7620:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35659-9_3/.

Aerts:2012:GEI

[307] Diederik Aerts, Jan Broekaert, Liane Gabora, and Tomas Veloz. The guppy effect as interference. Lecture Notes in Computer Science, 7620: 36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35659-9_4/.

Aerts:2012:QME

[308] Diederik Aerts, Sandro Sozzo, and Jocelyn Tapia. A quantum model for the Ellsberg and Machina paradoxes. Lecture Notes in Computer Science, 7620:48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35659-9_5/.

Asano:2012:QLM

[309] Masanari Asano, Irina Basieva, Andrei Khrennikov, and Masanori Ohya. A quantum-like model of Escherichia coli’s metabolism based on adaptive dynamics. Lecture Notes in Computer Science, 7620:60–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35659-9_6/.

Vitiello:2012:FDC

[310] Giuseppe Vitiello. Fractals, dissipation and coherent states. Lecture Notes in Computer Science, 7620:68–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35659-9_7/.

Busemeyer:2012:HBE

[311] Jerome R. Busemeyer, Zheng Wang, and Jennifer S. Trueblood. Hierarchical Bayesian estimation of quantum decision model parameters. Lecture Notes in Computer Science, 7620:80–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35659-9_8/.

Cohen:2012:MPL

[312] Trevor Cohen, Dominic Widdows, Lance De Vine, and Roger Schvaneveldt. Many paths lead to discovery: Analogical retrieval of cancer therapies. Lecture Notes in Computer Science, 7620:90–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35659-9_9/.

Lambert-Mogiliansky:2012:EII

[313] Ariane Lambert-Mogiliansky and Jerome R. Busemeyer. Emergence and instability of individual identity. Lecture Notes in Computer Science, 7620: 102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35659-9_10/.

Aerts:2012:ECE

[314] Diederik Aerts and Sandro Sozzo. Entanglement of conceptual entities in quantum model theory (QMod). Lecture Notes in Computer Science, 7620:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35659-9_11/.

Aerts:2012:QMT

[315] Diederik Aerts and Sandro Sozzo. Quantum model theory (QMod): Modeling contextual emergent entangled interfering entities. Lecture Notes in Computer Science, 7620:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35659-9_12/.

Asano:2012:QLR

[316] Masanari Asano, Irina Basieva, Andrei Khrennikov, and Masanori Ohya. Quantum-like representation of irrational inference. Lecture Notes in Computer Science, 7620:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35659-9_13/.

Flender:2012:TIP

[317] Christian Flender and Günter Müller. Type indeterminacy in privacy decisions: The privacy paradox revisited. Lecture Notes in Computer Science, 7620:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35659-9_14/.

Asano:2012:ADA

[318] Masanari Asano, Irina Basieva, Andrei Khrennikov, and Masanori Ohya. Adaptive dynamics and its application to context dependent systems breaking the classical probability law. Lecture Notes in Computer Science, 7620:160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35659-9_15/.

Galea:2012:MWA

[319] David Galea, Peter Bruza, Kirsty Kitto, and Douglas Nelson. Modelling word activation in semantic networks: Three scaled entanglement models compared. Lecture Notes in Computer Science, 7620: 172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35659-9_16/.

Dzhafarov:2012:QEI

[320] Ehtibar N. Dzhafarov and Janne V. Kujala. Quantum entanglement and the issue of selective influences in psychology: An overview. Lecture Notes in Computer Science, 7620:184–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35659-9_17/.

Filk:2012:QLB

[321] Thomas Filk. Quantum-like behavior of classical systems. Lecture Notes in Computer Science, 7620:196–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35659-9_18/.

Daranyi:2012:CDM

[322] Sándor Darányi and Peter Wittek. Connecting the dots: Mass, energy, word meaning, and particle-wave duality. Lecture Notes in Computer Science, 7620:207–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35659-9_19/.

Gondran:2012:IMF

[323] Michel Gondran and Sébastien Lepaul. Indiscernability and mean field, a base of quantum interaction. Lecture Notes in Computer Science, 7620: 218–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35659-9_20/.

Anonymous:2012:BMd

[324] Anonymous. Back matter. Lecture Notes in Computer Science, 7620: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35659-9/1.

Anonymous:2012:FMs

[325] Anonymous. Front matter. Lecture Notes in Computer Science, 7620: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35659-9/1.

Gray:2012:MMS

[326] Jesse Gray and Cynthia Breazeal. Manipulating mental states through physical action. Lecture Notes in Computer Science, 7621:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34103-8_1/.

Ham:2012:ASB

[327] Jaap Ham, Mirjam van Esch, Yvonne Limpens, Jente de Pee, and John-John Cabibihan. The automaticity of social behavior towards robots: The influence of cognitive load on interpersonal distance to approachable versus less approachable robots. Lecture Notes in Computer Science, 7621:15–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_2/.

Liu:2012:HMR

[328] Caixia Liu, Jaap Ham, Eric Postma, Cees Midden, Bart Joosten, and Martijn Goudbeek. How to make a robot smile? Perception of emotional expressions from digitally-extracted facial landmark configurations. Lecture Notes in Computer Science, 7621:26–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34103-8_3/.

Trovato:2012:CCS

[329] Gabriele Trovato, Tatsuhiro Kishi, Nobutsuna Endo, Kenji Hashimoto, and Atsuo Takanishi. A cross-cultural study on generation of culture dependent facial expressions of humanoid social robot. Lecture Notes in Computer Science, 7621:35–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34103-8_4/.

Williams:2012:RSI

[330] Mary-Anne Williams. Robot social intelligence. Lecture Notes in Computer Science, 7621:45–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34103-8_5/.

Vlachos:2012:AER

[331] Evgenios Vlachos and Henrik Schärfe. Android emotions revealed. Lecture Notes in Computer Science, 7621:56–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34103-8_6/.

Robins:2012:ECL

[332] Ben Robins, Kerstin Dautenhahn, and Paul Dickerson. Embodiment and cognitive learning — can a humanoid robot help children with autism to learn about tactile social behaviour? Lecture Notes in Computer Science, 7621:66–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_7/.

Wong:2012:UTS

[333] Alvin Wong, Yeow Kee Tan, Adrian Tay, Anthony Wong, Dilip Kumar Limbu, and Tran Anh Dung. A user trial study to understand play behaviors of autistic children using a social robot. Lecture Notes in Computer Science, 7621:76–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34103-8_8/.

Boccanfuso:2012:CHR

[334] Laura Boccanfuso, Eva Juarez Perez, Myra Robinson, and Jason M. O’Kane. Collecting heart rate using a high precision, non-contact, singlepoint infrared temperature sensor. Lecture Notes in Computer Science, 7621:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_9/.

Shimada:2012:HCS

[335] Michihiro Shimada, Takayuki Kanda, and Satoshi Koizumi. How can a social robot facilitate children’s collaboration? Lecture Notes in Computer Science, 7621:98–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34103-8_10/.

Klein:2012:ERE

[336] Barbara Klein and Glenda Cook. Emotional robotics in elder care — a comparison of findings in the UK and Germany. Lecture Notes in Computer Science, 7621:108–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34103-8_11/.

Yan:2012:DRR

[337] Haibin Yan, Marcelo H. Ang Jr., and Aun Neow Poo. Dorothy Robotubby: a robotic nanny. Lecture Notes in Computer Science, 7621: 118–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_12/.

Frennert:2012:WGL

[338] Susanne Frennert, Britt Östlund, and Håkan Eftring. Would Granny let an assistive robot into her home? Lecture Notes in Computer Science, 7621: 128–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_13/.

Kortner:2012:HSR

[339] Tobias Körtner, Alexandra Schmid, Daliah Batko-Klein, Christoph Gisinger, and Andreas Huber. How social robots make older users really feel well — a method to assess users’ concepts of a social robotic assistant. Lecture Notes in Computer Science, 7621:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34103-8_14/.

Stafford:2012:ISR

[340] Rebecca Q. Stafford, Bruce A. MacDonald, and Elizabeth Broadbent. Identifying specific reasons behind unmet needs may inform more specific eldercare robot design. Lecture Notes in Computer Science, 7621: 148–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_15/.

Yamazaki:2012:VFH

[341] Akira Yamazaki and Ryosuke Masuda. Various foods handling movement of chopstick-equipped meal assistant robot and there evaluation. Lecture Notes in Computer Science, 7621:158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34103-8_16/.

Yumakulov:2012:IDP

[342] Sophya Yumakulov, Dean Yergens, and Gregor Wolbring. Imagery of disabled people within social robotics research. Lecture Notes in Computer Science, 7621:168–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34103-8_17/.

Kuo:2012:HEH

[343] I-Han Kuo, Chandimal Jayawardena, Elizabeth Broadbent, and Rebecca Q. Stafford. HRI evaluation of a healthcare service robot. Lecture Notes in Computer Science, 7621:178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34103-8_18/.

Rebola:2012:ABM

[344] Claudia B. Rebola, Gbolabo Ogunmakin, and Patricio A. Vela. Automated behavioral mapping for monitoring social interactions among older adults. Lecture Notes in Computer Science, 7621:188–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34103-8_19/.

Fink:2012:AHL

[345] Julia Fink. Anthropomorphism and human likeness in the design of robots and human-robot interaction. Lecture Notes in Computer Science, 7621: 199–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34103-8_20/.

Anonymous:2012:FMt

[346] Anonymous. Front matter. Lecture Notes in Computer Science, 7621: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34103-8/1.

Christianson:2012:IBP

[347] Bruce Christianson. Introduction: Bringing protocols to life (transcript of discussion). Lecture Notes in Computer Science, 7622:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-35694-0_1.

Andel:2012:SIVa

[348] Todd R. Andel and Alec Yasinsac. Secure Internet voting protocol for overseas military voters. Lecture Notes in Computer Science, 7622: 3–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_2/.

Andel:2012:SIVb

[349] Todd R. Andel. Secure Internet voting protocol for overseas military voters (transcript of discussion). Lecture Notes in Computer Science, 7622:15–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_3/.

Hao:2012:SEEa

[350] Feng Hao, Brian Randell, and Dylan Clarke. Self-enforcing electronic voting. Lecture Notes in Computer Science, 7622:23–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35694-0_4/.

Hao:2012:SEEb

[351] Feng Hao. Self-enforcing electronic voting (transcript of discussion). Lecture Notes in Computer Science, 7622:32–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35694-0_5/.

Brooke:2012:AMSa

[352] Phillip J. Brooke, Richard F. Paige, and Christopher Power. Approaches to modelling security scenarios with domain-specific languages. Lecture Notes in Computer Science, 7622:41–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35694-0_6/.

Brooke:2012:AMSb

[353] Phillip J. Brooke. Approaches to modelling security scenarios with domain-specific languages(transcript of discussion). Lecture Notes in Computer Science, 7622:55–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35694-0_7/.

Clark:2012:COL

[354] Sandy Clark, Matt Blaze, and Jonathan M. Smith. The casino and the OODA loop. Lecture Notes in Computer Science, 7622:60–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35694-0_8/.

Blaze:2012:COL

[355] Matt Blaze. The casino and the OODA loop: Why our protocols always eventually fail (transcript of discussion). Lecture Notes in Computer Science, 7622:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35694-0_9/.

Bonneau:2012:SMIa

[356] Joseph Bonneau. Statistical metrics for individual password strength. Lecture Notes in Computer Science, 7622:76–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35694-0_10/.

Bonneau:2012:SMIb

[357] Joseph Bonneau. Statistical metrics for individual password strength (transcript of discussion). Lecture Notes in Computer Science, 7622: 87–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_11/.

Kim:2012:SLT

[358] Tiffany Hyun-Jin Kim, Virgil Gligor, and Adrian Perrig. Street-level trust semantics for attribute authentication. Lecture Notes in Computer Science, 7622:96–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35694-0_12/.

Gligor:2012:SLT

[359] Virgil Gligor. Street-level trust semantics for attribute authentication (transcript of discussion). Lecture Notes in Computer Science, 7622: 116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_13/.

Clarke:2012:AICa

[360] Dylan Clarke, Feng Hao, and Brian Randell. Analysis of issues and challenges of E-voting in the UK. Lecture Notes in Computer Science, 7622: 126–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_14/.

Clarke:2012:AICb

[361] Dylan Clarke. Analysis of issues and challenges of E-voting in the UK (transcript of discussion). Lecture Notes in Computer Science, 7622: 136–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_15/.

Anderson:2012:PGEa

[362] Ross Anderson. Protocol governance: The elite, or the mob? Lecture Notes in Computer Science, 7622:145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-35694-0_16.

Anderson:2012:PGEb

[363] Ross Anderson. Protocol governance: The elite, or the mob? (transcript of discussion). Lecture Notes in Computer Science, 7622:146– 160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_17/.

Murayama:2012:UISa

[364] Yuko Murayama, Yasuhiro Fujihara, Yoshia Saito, and Dai Nishioka. Usability issues in security. Lecture Notes in Computer Science, 7622: 161–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_18/.

Murayama:2012:UISb

[365] Yuko Murayama. Usability issues in security (transcript of discussion). Lecture Notes in Computer Science, 7622:172–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35694-0_19/.

Mahmood:2012:UPV

[366] Shah Mahmood and Yvo Desmedt. Usable privacy by visual and interactive control of information flow. Lecture Notes in Computer Science, 7622:

181–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35694-0_20/.

Anonymous:2012:FMu

[367] Anonymous. Front matter. Lecture Notes in Computer Science, 7622: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35694-0/1.

Stanton:2012:HFE

[368] Neville A. Stanton. Human factors engineering as the methodological Babel fish: Translating user needs into software design. Lecture Notes in Computer Science, 7623:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34347-6_1/.

Mendes:2012:ISE

[369] Emilia Mendes. Improving software effort estimation using an expertcentred approach. Lecture Notes in Computer Science, 7623:18–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34347-6_2/.

Spano:2012:CMG

[370] Lucio Davide Spano, Antonio Cisternino, and Fabio Paternò. A compositional model for gesture definition. Lecture Notes in Computer Science, 7623:34–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_3/.

Masip:2012:DPE

[371] Llúcia Masip, Célia Martinie, Marco Winckler, Philippe Palanque, and Toni Granollers. A design process for exhibiting design choices and tradeoffs in (potentially) conflicting user interface guidelines. Lecture Notes in Computer Science, 7623:53–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34347-6_4/.

Martinie:2012:DPU

[372] Célia Martinie, Philippe Palanque, David Navarre, and Eric Barboni. A development process for usable large scale interactive critical systems: Application to satellite ground segments. Lecture Notes in Computer Science, 7623:72–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_5/.

Kuusinen:2012:AUE

[373] Kati Kuusinen, Tommi Mikkonen, and Santtu Pakarinen. Agile user experience development in a large software organization: Good expertise but limited impact. Lecture Notes in Computer Science, 7623: 94–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_6/.

Silva:2012:CGI

[374] Carlos Eduardo Silva and José Creissac Campos. Can GUI implementation markup languages be used for modelling? Lecture Notes in Computer Science, 7623:112–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34347-6_7/.

Roast:2012:CCS

[375] Chris Roast. Constraining and creating solutions — reflections on the analysis of early design. Lecture Notes in Computer Science, 7623: 130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_8/.

Kronbauer:2012:SAU

[376] Artur H. Kronbauer, Celso A. S. Santos, and Vaninha Vieira. Smartphone applications usability evaluation: a hybrid model and its implementation. Lecture Notes in Computer Science, 7623:146–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34347-6_9/.

Grill:2012:MTA

[377] Thomas Grill, Ondrej Polacek, and Manfred Tscheligi. Methods towards API usability: a structural analysis of usability problem categories. Lecture Notes in Computer Science, 7623:164–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34347-6_10/.

Vogt:2012:PDC

[378] Joël Vogt, Kris Luyten, Jan Van den Bergh, Karin Coninx, and Andreas Meier. Putting dementia into context. Lecture Notes in Computer Science, 7623:181–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34347-6_11/.

Danado:2012:PVB

[379] Jose Danado and Fabio Paternò. Puzzle: a visual-based environment for end user development in touch-based mobile phones. Lecture Notes in Computer Science, 7623:199–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34347-6_12/.

Faily:2012:RSU

[380] Shamal Faily, John Lyle, Andre Paul, Andrea Atzeni, Dieter Blomme, and Heiko Desruelle. Requirements sensemaking using concept maps. Lecture Notes in Computer Science, 7623:217–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34347-6_13/.

Dery-Pinna:2012:TCM

[381] Anne-Marie Déry-Pinna, Audrey Occello, and Michel Riveill. Towards conflict management in user interface composition driven by business needs. Lecture Notes in Computer Science, 7623:233–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34347-6_14/.

Chouseinoglou:2012:MAO

[382] Oumout Chouseinoglou and Semih Bilgen. A model for assessing organizational learning in software development organizations. Lecture Notes in Computer Science, 7623:251–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34347-6_15/.

Behrenbruch:2012:PBD

[383] Kay Behrenbruch, Martin Atzmüller, Christoph Evers, Ludger Schmidt, and Gerd Stumme. A personality based design approach using subgroup discovery. Lecture Notes in Computer Science, 7623:259–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34347-6_16/.

vanLoggem:2012:AUC

[384] Brigit van Loggem. Assessing use complexity of software: a tool for documentation designers. Lecture Notes in Computer Science, 7623:

267–274, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_17/.

Rasooli:2012:CUP

[385] Amin Rasooli, Peter Forbrig, and Fattaneh Tagivareh. Collecting users profiles for Web applications. Lecture Notes in Computer Science, 7623: 275–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_18/.

Vieira:2012:CPG

[386] Elton R. Vieira, Carina Alves, and Letícia Duboc. Creativity patterns guide: Support for the application of creativity techniques in requirements engineering. Lecture Notes in Computer Science, 7623: 283–290, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34347-6_19/.

Ressin:2012:ELC

[387] Malte Ressin, Cecilia Oyugi, José Abdelnour-Nocera, David Lee, and Dharam Panesar. Exploring local cultural perspectives in user interface development in an Indian offshoring context: a view from the UK. Lecture Notes in Computer Science, 7623:291–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34347-6_20/.

Anonymous:2012:FMv

[388] Anonymous. Front matter. Lecture Notes in Computer Science, 7623: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34347-6/1.

Centieiro:2012:AHF

[389] Pedro Centieiro, Teresa Romão, and A. Eduardo Dias. Applaud having fun: a mobile game to cheer your favourite sports team. Lecture Notes in Computer Science, 7624:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34292-9_1/.

Fujita:2012:PIF

[390] Kazuyuki Fujita, Hiroyuki Kidokoro, and Yuichi Itoh. Paranga: An interactive flipbook. Lecture Notes in Computer Science, 7624:17–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34292-9_2/.

Hachisu:2012:ATM

[391] Taku Hachisu and Hiroyuki Kajimoto. Augmentation of toothbrush by modulating sounds resulting from brushing. Lecture Notes in Computer Science, 7624:31–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34292-9_3/.

Hirai:2012:BTS

[392] Shigeyuki Hirai, Yoshinobu Sakakibara, and Seiho Hayakawa. Bathcratch: Touch and sound-based DJ controller implemented on a bathtub. Lecture Notes in Computer Science, 7624:44–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34292-9_4/.

Kanke:2012:ADD

[393] Hiroyuki Kanke, Yoshinari Takegawa, Tsutomu Terada, and Masahiko Tsukamoto. Airstic drum: a drumstick for integration of real and virtual drums. Lecture Notes in Computer Science, 7624:57–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34292-9_5/.

Mourato:2012:ELD

[394] Fausto Mourato, Fernando Birra, and Manuel Próspero dos Santos. Enhancing level difficulty and additional content in platform videogames through graph analysis. Lecture Notes in Computer Science, 7624: 70–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34292-9_6/.

Okazaki:2012:SSP

[395] Tatsuhiko Okazaki, Tsutomu Terada, and Masahiko Tsukamoto. A system for supporting performers in stuffed suits. Lecture Notes in Computer Science, 7624:85–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34292-9_7/.

Pereira:2012:SPB

[396] André Pereira, Rui Prada, and Ana Paiva. Socially present board game opponents. Lecture Notes in Computer Science, 7624:101–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34292-9_8/.

Shin:2012:LGG

[397] Kiyoshi Shin, Kosuke Kaneko, Yu Matsui, Koji Mikami, and Masaru Nagaku. Localizing global game jam: Designing game development for collaborative learning in the social context. Lecture Notes in Computer Science, 7624:117–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34292-9_9/.

Vyas:2012:PWC

[398] Dhaval Vyas, Yanqing Cui, Jarno Ojala, and Guido Grassel. Producing while consuming: Social interaction around photos shared within private group. Lecture Notes in Computer Science, 7624:133–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34292-9_10/.

Chan:2012:ESD

[399] Shih-Han Chan, Stéphane Natkin, Guillaume Tiger, and Alexandre Topol. Extensible sound description in COLLADA: a unique file for a rich sound design. Lecture Notes in Computer Science, 7624:151–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34292-9_11/.

Chen:2012:ART

[400] Tai-Yun Chen, Hung-Wei Hsu, Wen-Kai Tai, and Chin-Chen Chang. An automatic race track generating system. Lecture Notes in Computer Science, 7624:167–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34292-9_12/.

Choi:2012:LPD

[401] Yongsoon Choi, Rahul Parsani, Xavier Roman, Anshul Vikram Pandey, and Adrian David Cheok. Light perfume: Designing a wearable lighting and olfactory accessory for empathic interactions. Lecture Notes in Computer Science, 7624:182–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34292-9_13/.

Deaker:2012:SPO

[402] Chris Deaker, Masood Masoodian, and Bill Rogers. A survey of players’ opinions on interface customization in world of warcraft. Lecture Notes in Computer Science, 7624:198–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34292-9_14/.

Grammenos:2012:VRR

[403] Dimitris Grammenos, George Margetis, Panagiotis Koutlemanis, and Xenophon Zabulis. 53.090 virtual rusks = 510 real smiles using a fun exergame installation for advertising traditional food products. Lecture Notes in Computer Science, 7624:214–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34292-9_15/.

Gronbaek:2012:DPI

[404] Kaj Grønbæk, Karen Johanne Kortbek, Claus Møller, Jesper Nielsen, and Liselott Stenfeldt. Designing playful interactive installations for urban environments — the SwingScape experience. Lecture Notes in Computer Science, 7624:230–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34292-9_16/.

Guy:2012:FIS

[405] Olivier Guy and Ronan Champagnat. Flashback in interactive storytelling. Lecture Notes in Computer Science, 7624:246–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34292-9_17/.

Howland:2012:SCA

[406] Robert Howland, Sachi Urano, and Junichi Hoshino. SanjigenJiten: Computer assisted language learning system within a 3D game environment. Lecture Notes in Computer Science, 7624:262–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34292-9_18/.

Konya:2012:CPS

[407] Yuko Konya and Itiro Siio. A caption presentation system for the hearing impaired people attending theatrical performances. Lecture Notes in Computer Science, 7624:274–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34292-9_19/.

Kou:2012:EGE

[408] Kinyo Kou and Yoichiro Kawaguchi. Emergent gait evolution of quadruped artificial life. Lecture Notes in Computer Science, 7624:

287–296, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34292-9_20/.

Anonymous:2012:FMw

[409] Anonymous. Front matter. Lecture Notes in Computer Science, 7624: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34292-9/1.

Andrei:2012:TBA

[410] Oana Andrei and Muffy Calder. Trend-based analysis of a population model of the AKAP scaffold protein. Lecture Notes in Computer Science, 7625:1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35524-0_1/.

Angius:2012:QPF

[411] Alessio Angius, András Horváth, and Verena Wolf. Quasi product form approximation for Markov models of reaction networks. Lecture Notes in Computer Science, 7625:26–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35524-0_2/.

Bartocci:2012:MVC

[412] Ezio Bartocci, Pietro Liò, Emanuela Merelli, and Nicola Paoletti. Multiple verification in complex biological systems: The bone remodelling case study. Lecture Notes in Computer Science, 7625:53–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35524-0_3/.

Brim:2012:ARA

[413] L. Brim, J. Fabriková, S. Dražan, and D. Šafránek. On approximative reachability analysis of biochemical dynamical systems. Lecture Notes in Computer Science, 7625:77–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35524-0_4/.

Ehrenfeucht:2012:MRS

[414] Andrzej Ehrenfeucht, Jetty Kleijn, Maciej Koutny, and Grzegorz Rozenberg. Minimal reaction systems. Lecture Notes in Computer Science, 7625: 102–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35524-0_5/.

Versari:2012:CFR

[415] Cristian Versari and Gianluigi Zavattaro. Complex functional rates in rulebased languages for biochemistry. Lecture Notes in Computer Science, 7625:123–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35524-0_6/.

Yuan:2012:PMC

[416] Qixia Yuan, Panuwat Trairatphisan, Jun Pang, Sjouke Mauw, and Monique Wiesinger. Probabilistic model checking of the PDGF signaling pathway. Lecture Notes in Computer Science, 7625:151–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35524-0_7/.

Mizera:2012:SAM

[417] Andrzej Mizera, Eugen Czeizler, and Ion Petre. Self-assembly models of variable resolution. Lecture Notes in Computer Science, 7625: 181–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35524-0_8/.

Wallace:2012:MCE

[418] Rodrick Wallace. Metabolic constraints on the evolution of genetic codes: Did multiple preaerobic’ ecosystem transitions entrain richer dialects via serial endosymbiosis? Lecture Notes in Computer Science, 7625: 204–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35524-0_9/.

Anonymous:2012:BMe

[419] Anonymous. Back matter. Lecture Notes in Computer Science, 7625: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35524-0/1.

Anonymous:2012:FMx

[420] Anonymous. Front matter. Lecture Notes in Computer Science, 7625: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35524-0/1.

Nagy:2012:ELA

[421] George Nagy. Estimation, learning, and adaptation: Systems that improve with use. Lecture Notes in Computer Science, 7626:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34166-3_1/.

Kanatani:2012:OTG

[422] Kenichi Kanatani. Optimization techniques for geometric estimation: Beyond minimization. Lecture Notes in Computer Science, 7626:11– 30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34166-3_2/.

Leonardis:2012:HCR

[423] Aleš Leonardis. Hierarchical compositional representations of object structure. Lecture Notes in Computer Science, 7626:31–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34166-3_3.

Han:2012:ITP

[424] Lin Han, Luca Rossi, Andrea Torsello, Richard C. Wilson, and Edwin R. Hancock. Information theoretic prototype selection for unattributed graphs. Lecture Notes in Computer Science, 7626:33–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34166-3_4/.

Gauzere:2012:GKC

[425] Benoit Gaüzère, Luc Brun, and Didier Villemin. Graph kernels: Crossing information from different patterns using graph edit distance. Lecture Notes in Computer Science, 7626:42–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34166-3_5/.

Duin:2012:MSC

[426] Robert P. W. Duin, Ana L. N. Fred, Marco Loog, and Elżbieta Pekalska. Mode seeking clustering by KNN and mean shift evaluated. Lecture Notes in Computer Science, 7626:51–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34166-3_6/.

Fu:2012:LSK

[427] Zhouyu Fu, Guojun Lu, Kai-Ming Ting, and Dengsheng Zhang. Learning sparse kernel classifiers in the primal. Lecture Notes in Computer Science, 7626:60–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34166-3_7/.

Franek:2012:EWM

[428] Lucas Franek and Xiaoyi Jiang. Evolutionary weighted mean based framework for generalized median computation with application to strings. Lecture Notes in Computer Science, 7626:70–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34166-3_8/.

Bai:2012:GCJ

[429] Lu Bai and Edwin R. Hancock. Graph complexity from the Jensen– Shannon divergence. Lecture Notes in Computer Science, 7626:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34166-3_9/.

Feragen:2012:CCD

[430] Aasa Feragen. Complexity of computing distances between geometric trees. Lecture Notes in Computer Science, 7626:89–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34166-3_10/.

Cortes:2012:AGM

[431] Xavier Cortés, Francesc Serratosa, and Albert Solé-Ribalta. Active graph matching based on pairwise probabilities between nodes. Lecture Notes in Computer Science, 7626:98–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34166-3_11/.

Rebagliati:2012:RBC

[432] Nicola Rebagliati, Albert Solé-Ribalta, Marcello Pelillo, and Francesc Serratosa. On the relation between the common labelling and the median graph. Lecture Notes in Computer Science, 7626:107–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34166-3_12/.

Guimaraes:2012:HIS

[433] Silvio Jamil F. Guimarães, Jean Cousty, Yukiko Kenmochi, and Laurent Najman. A hierarchical image segmentation algorithm based on an observation scale. Lecture Notes in Computer Science, 7626:116– 125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34166-3_13/.

Tschirsich:2012:DSS

[434] Martin Tschirsich and Arjan Kuijper. A discrete scale space neighborhood for robust deep structure extraction. Lecture Notes in Computer Science, 7626:126–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34166-3_14/.

Gibert:2012:CGE

[435] Jaume Gibert, Ernest Valveny, Horst Bunke, and Alicia Fornés. On the correlation of graph edit distance and L_{1} distance in the attribute statistics embedding space. Lecture Notes in Computer Science, 7626: 135–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34166-3_15/.

Rossi:2012:AAS

[436] Luca Rossi, Andrea Torsello, and Edwin R. Hancock. Approximate axial symmetries from continuous time quantum walks. Lecture Notes in Computer Science, 7626:144–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34166-3_16/.

Lewin:2012:CBE

[437] Sergej Lewin, Xiaoyi Jiang, and Achim Clausing. A clustering-based ensemble technique for shape decomposition. Lecture Notes in Computer Science, 7626:153–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34166-3_17/.

Tschirsich:2012:LED

[438] Martin Tschirsich and Arjan Kuijper. Laplacian eigenimages in discrete scale space. Lecture Notes in Computer Science, 7626:162–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34166-3_18/.

Antanas:2012:RKB

[439] Laura Antanas, Paolo Frasconi, Fabrizio Costa, Tinne Tuytelaars, and Luc De Raedt. A relational kernel-based framework for hierarchical image understanding. Lecture Notes in Computer Science, 7626:171–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34166-3_19/.

Bai:2012:JSK

[440] Lu Bai, Edwin R. Hancock, and Peng Ren. A Jensen–Shannon kernel for hypergraphs. Lecture Notes in Computer Science, 7626:181– 189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34166-3_20/.

Anonymous:2012:FMy

[441] Anonymous. Front matter. Lecture Notes in Computer Science, 7626: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34166-3/1.

Laumond:2012:GPA

[442] Jean-Paul Laumond. A geometric perspective of anthropomorphic embodied actions. Lecture Notes in Computer Science, 7628:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34327-8_1.

Sankai:2012:CFH

[443] Yoshiyuki Sankai. Cybernics: Fusion of human, machine and information. Lecture Notes in Computer Science, 7628:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-34327-8_2.

Beetz:2012:IAA

[444] Michael Beetz. If abstraction is the answer, what is the question?— reasoning for everyday manipulation tasks. Lecture Notes in Computer Science, 7628:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-34327-8_3.

Anzalone:2012:TPP

[445] Salvatore M. Anzalone, Yuichiro Yoshikawa, Hiroshi Ishiguro, and Emanuele Menegatti. Towards partners profiling in human robot interaction contexts. Lecture Notes in Computer Science, 7628:4–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34327-8_4/.

Vladimirov:2012:MBA

[446] Blagovest Vladimirov, Hyun Kim, and Namshik Park. Motivation-based autonomous behavior control of robotic computer. Lecture Notes in Computer Science, 7628:16–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34327-8_5/.

Wada:2012:EMS

[447] Kazuyoshi Wada, Keisuke Takayama, Yusuke Suganuma, and Toshihiko Suzuki. An evaluation method for smart variable space in living space. Lecture Notes in Computer Science, 7628:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34327-8_6/.

Kulakowski:2012:MRB

[448] Konrad Kułakowski and Tomasz Szmuc. Modeling robot behavior with CCL. Lecture Notes in Computer Science, 7628:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34327-8_7/.

Coleman:2012:VTS

[449] Robert Coleman, Vladimir Estivill-Castro, René Hexel, and Carl Lusty. Visual-trace simulation of concurrent finite-state machines for validation and model-checking of complex behaviour. Lecture Notes in Computer Science, 7628:52–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34327-8_8/.

Drumwright:2012:FDS

[450] Evan Drumwright. Fast dynamic simulation of highly articulated robots with contact via Θ(n^{2}) time dense generalized inertia matrix inversion. Lecture Notes in Computer Science, 7628:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34327-8_9/.

Xiong:2012:DAM

[451] Xiaogang Xiong, Ryo Kikuuwe, and Motoji Yamamoto. A differentialalgebraic multistate friction model. Lecture Notes in Computer Science, 7628:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34327-8_10/.

Fugl:2012:SFO

[452] Andreas Rune Fugl, Henrik Gordon Petersen, and Morten Willatzen. Simulation of flexible objects in robotics. Lecture Notes in Computer Science, 7628:89–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34327-8_11/.

Lier:2012:CII

[453] Florian Lier, Simon Schulz, and Ingo Lütkebohle. Continuous integration for iterative validation of simulated robot models. Lecture Notes in Computer Science, 7628:101–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34327-8_12/.

Nordmann:2012:SAS

[454] Arne Nordmann, Matthias Rolf, and Sebastian Wrede. Software abstractions for simulation and control of a continuum robot. Lecture Notes in Computer Science, 7628:113–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34327-8_13/.

Kilgo:2012:VML

[455] Paul Kilgo, Eugene Syriani, and Monica Anderson. A visual modeling language for RDIS and ROS nodes using AToM^{3}. Lecture Notes in Computer Science, 7628:125–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34327-8_14/.

Kimmel:2012:PEA

[456] Andrew Kimmel, Andrew Dobson, Zakary Littlefield, Athanasios Krontiris, and James Marble. PRACSYS: An extensible architecture for composing motion controllers and planners. Lecture Notes in Computer Science, 7628:137–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34327-8_15/.

Dhouib:2012:RDS

[457] Saadia Dhouib, Selma Kchir, Serge Stinckwich, Tewfik Ziadi, and Mikal Ziane. RobotML, a domain-specific language to design, simulate and deploy robotic applications. Lecture Notes in Computer Science, 7628: 149–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34327-8_16/.

Gherardi:2012:JVC

[458] Luca Gherardi, Davide Brugali, and Daniele Comotti. A Java vs. C++ performance evaluation: a 3D modeling benchmark. Lecture Notes in Computer Science, 7628:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34327-8_17/.

Klaus:2012:CSS

[459] Gordon Klaus, Kyrre Glette, and Jim Tørresen. A comparison of sampling strategies for parameter estimation of a robot simulator. Lecture Notes in Computer Science, 7628:173–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34327-8_18/.

Shiomi:2012:FPS

[460] Masahiro Shiomi, Francesco Zanlungo, Kotaro Hayashi, and Takayuki Kanda. A framework with a pedestrian simulator for deploying robots into a real environment. Lecture Notes in Computer Science, 7628: 185–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34327-8_19/.

Echeverria:2012:SCR

[461] Gilberto Echeverria, Séverin Lemaignan, Arnaud Degroote, Simon Lacroix, and Michael Karg. Simulating complex robotic scenarios with MORSE. Lecture Notes in Computer Science, 7628:197–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34327-8_20/.

Anonymous:2012:FMz

[462] Anonymous. Front matter. Lecture Notes in Computer Science, 7628: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34327-8/1.

Jing:2012:MBC

[463] Yiming Jing, Gail-Joon Ahn, and Hongxin Hu. Model-based conformance testing for Android. Lecture Notes in Computer Science, 7631: 1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34117-5_1/.

Yasuda:2012:ASM

[464] Takanori Yasuda, Tsuyoshi Takagi, and Kouichi Sakurai. Application of scalar multiplication of Edwards curves to pairing-based cryptography. Lecture Notes in Computer Science, 7631:19–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34117-5_2/.

Chen:2012:SSA

[465] Chien-Ning Chen, Nisha Jacob, Sebastian Kutzner, San Ling, and Axel Poschmann. Standardized signature algorithms on ultra-constrained 4-bit MCU. Lecture Notes in Computer Science, 7631:37–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34117-5_3/.

Nekado:2012:VSC

[466] Kenta Nekado, Yasuyuki Nogami, and Kengo Iokibe. Very short critical path implementation of AES with direct logic gates. Lecture Notes in Computer Science, 7631:51–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34117-5_4/.

Yoneyama:2012:ORA

[467] Kazuki Yoneyama. One-round authenticated key exchange with strong forward secrecy in the standard model against constrained adversary. Lecture Notes in Computer Science, 7631:69–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34117-5_5/.

Vivek:2012:CSE

[468] S. Sree Vivek, S. Sharmila Deva Selvi, and C. Pandu Rangan. Compact stateful encryption schemes with ciphertext verifiability. Lecture Notes in Computer Science, 7631:87–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34117-5_6/.

Poh:2012:SEC

[469] Geong Sen Poh, Moesfa Soeheila Mohamad, and Muhammad Reza Z’aba. Structured encryption for conceptual graphs. Lecture Notes in Computer Science, 7631:105–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34117-5_7/.

Kawachi:2012:SKE

[470] Akinori Kawachi, Hirotoshi Takebe, and Keisuke Tanaka. Symmetric-key encryption scheme with multi-ciphertext non-malleability. Lecture Notes in Computer Science, 7631:123–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34117-5_8/.

Isobe:2012:SCL

[471] Takanori Isobe, Toshihiro Ohigashi, and Masakatu Morii. Slide cryptanalysis of lightweight stream cipher RAKAPOSHI. Lecture Notes in Computer Science, 7631:138–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34117-5_9/.

Sasaki:2012:BDF

[472] Yu Sasaki, Lei Wang, Yasuhiro Takasaki, Kazuo Sakiyama, and Kazuo Ohta. Boomerang distinguishers for full HAS-160 compression function. Lecture Notes in Computer Science, 7631:156–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34117-5_10/.

Wang:2012:PAC

[473] Lei Wang, Yu Sasaki, Kazuo Sakiyama, and Kazuo Ohta. Polynomialadvantage cryptanalysis of 3D cipher and 3D-based hash function. Lecture Notes in Computer Science, 7631:170–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34117-5_11/.

Wang:2012:AFD

[474] Jingjing Wang, Kefei Chen, and Shixiong Zhu. Annihilators of fast discrete Fourier spectra attacks. Lecture Notes in Computer Science, 7631: 182–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34117-5_12/.

Lu:2012:MMA

[475] Jiqiang Lu, Yongzhuang Wei, Enes Pasalic, and Pierre-Alain Fouque. Meet-in-the-middle attack on reduced versions of the Camellia block cipher. Lecture Notes in Computer Science, 7631:197–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34117-5_13/.

Kiyoshima:2012:ECO

[476] Susumu Kiyoshima, Yoshifumi Manabe, and Tatsuaki Okamoto. Efficient concurrent oblivious transfer in super-polynomial-simulation security. Lecture Notes in Computer Science, 7631:216–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34117-5_14/.

Zhu:2012:ESP

[477] Youwen Zhu, Tsuyoshi Takagi, and Liusheng Huang. Efficient secure primitive for privacy preserving distributed computations. Lecture Notes in Computer Science, 7631:233–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34117-5_15/.

Suzuki:2012:GCG

[478] Itsuki Suzuki, Maki Yoshida, and Toru Fujiwara. Generic construction of GUC secure commitment in the KRK model. Lecture Notes in Computer Science, 7631:244–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34117-5_16/.

Anonymous:2012:BMf

[479] Anonymous. Back matter. Lecture Notes in Computer Science, 7631: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34117-5/1.

Anonymous:2012:FMba

[480] Anonymous. Front matter. Lecture Notes in Computer Science, 7631: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34117-5/1.

vanLaarhoven:2012:RCD

[481] Twan van Laarhoven and Elena Marchiori. Robust community detection methods with resolution parameter for complex detection in protein protein interaction networks. Lecture Notes in Computer Science, 7632:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_1/.

Ballester:2012:MLS

[482] Pedro J. Ballester. Machine learning scoring functions based on random forest and support vector regression. Lecture Notes in Computer Science, 7632:14–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_2/.

Laroum:2012:GAS

[483] Sami Laroum, Béatrice Duval, Dominique Tessier, and Jin-Kao Hao. A genetic algorithm for scale-based translocon simulation. Lecture Notes in Computer Science, 7632:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34123-6_3/.

Li:2012:FGS

[484] Yifeng Li, Alioune Ngom, and Luis Rueda. A framework of gene subset selection using multiobjective evolutionary algorithm. Lecture Notes in Computer Science, 7632:38–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34123-6_4/.

Hosoda:2012:MTA

[485] Masae Hosoda, Yukie Akune, and Kiyoko F. Aoki-Kinoshita. Multiple tree alignment with weights applied to carbohydrates to extract binding recognition patterns. Lecture Notes in Computer Science, 7632: 49–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_5/.

Zhang:2012:UAC

[486] Shuzhong Zhang, Kun Wang, Cody Ashby, Bilian Chen, and Xiuzhen Huang. A unified adaptive co-identification framework for high-D expression data. Lecture Notes in Computer Science, 7632:59–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34123-6_6/.

Suryanto:2012:PCG

[487] Chendra Hadi Suryanto, Hiroto Saigo, and Kazuhiro Fukui. Protein clustering on a Grassmann manifold. Lecture Notes in Computer Science, 7632:71–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_7/.

Shabash:2012:IPP

[488] Boris Shabash, Kay Wiese, and Edward Glen. Improving the portability and performance of jViz.RNA — a dynamic RNA visualization software. Lecture Notes in Computer Science, 7632:82–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34123-6_8/.

Singh:2012:NML

[489] Lavneet Singh, Girija Chetty, and Dharmendra Sharma. A novel machine learning approach for detecting the brain abnormalities from MRI structural images. Lecture Notes in Computer Science, 7632:94–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34123-6_9/.

Keeble-Gagnere:2012:OFE

[490] Gabriel Keeble-Gagnère, Johan Nyström-Persson, Matthew I. Bellgard, and Kenji Mizuguchi. An open framework for extensible multi-stage bioinformatics software. Lecture Notes in Computer Science, 7632:106– 117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_10/.

Cardoso:2012:AAG

[491] João Cardoso, Paulo Vilaça, Simão Soares, and Miguel Rocha. An algorithm to assemble gene-protein-reaction associations for genome-scale metabolic model reconstruction. Lecture Notes in Computer Science, 7632: 118–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_11/.

Maraschin:2012:MLC

[492] Marcelo Maraschin, Amélia Somensi-Zeggio, Simone K. Oliveira, and Shirley Kuhnen. A machine learning and chemometrics assisted interpretation of spectroscopic data — a NMR-based metabolomics platform for the assessment of Brazilian Propolis. Lecture Notes in Computer Science, 7632:129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34123-6_12/.

Taguchi:2012:PCA

[493] Y.-H. Taguchi and Akira Okamoto. Principal component analysis for bacterial proteomic analysis. Lecture Notes in Computer Science, 7632: 141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34123-6_13/.

Razin:2012:AMM

[494] Nikolay Razin, Dmitry Sungurov, Vadim Mottl, Ivan Torshin, and Valentina Sulimova. Application of the multi-modal relevance vector machine to the problem of protein secondary structure prediction. Lecture Notes in Computer Science, 7632:153–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34123-6_14/.

Thomarat:2012:CDG

[495] Fabienne Thomarat, Fabien Lauer, and Yann Guermeur. Cascading discriminant and generative models for protein secondary structure prediction. Lecture Notes in Computer Science, 7632:166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34123-6_15/.

Ohue:2012:IPP

[496] Masahito Ohue, Yuri Matsuzaki, Takashi Ishida, and Yutaka Akiyama. Improvement of the protein–protein docking prediction by introducing a simple hydrophobic interaction model: An application to interaction pathway analysis. Lecture Notes in Computer Science, 7632:178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34123-6_16/.

Meydan:2012:RPS

[497] Cem Meydan and Osman Ugur Sezerman. Representation of protein secondary structure using bond-orientational order parameters. Lecture Notes in Computer Science, 7632:188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34123-6_17/.

Li:2012:DPP

[498] Yifeng Li and Alioune Ngom. Diagnose the premalignant pancreatic cancer using high dimensional linear machine. Lecture Notes in Computer Science, 7632:198–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34123-6_18/.

Malhotra:2012:PVD

[499] Raunaq Malhotra, Shruthi Prabhakara, and Raj Acharya. Predicting V(D)J recombination using conditional random fields. Lecture Notes in Computer Science, 7632:210–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34123-6_19/.

Popovic:2012:SGA

[500] Dusan Popovic, Alejandro Sifrim, Georgios A. Pavlopoulos, Yves Moreau, and Bart De Moor. A simple genetic algorithm for biomarker mining. Lecture Notes in Computer Science, 7632:222–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34123-6_20/.

Anonymous:2012:FMbb

[501] Anonymous. Front matter. Lecture Notes in Computer Science, 7632: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34123-6/1.

Qu:2012:ISD

[502] Zhenhua Qu, Weiqi Luo, and Jiwu Huang. Identifying shifted double JPEG compression artifacts for non-intrusive digital image forensics. Lecture Notes in Computer Science, 7633:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34263-9_1/.

Ding:2012:NCR

[503] Shouhong Ding, Feiyue Huang, Zhifeng Xie, Yongjian Wu, and Lizhuang Ma. A novel customized recompression framework for massive Internet images. Lecture Notes in Computer Science, 7633:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34263-9_2/.

Qin:2012:DEB

[504] Xueying Qin, Rui Zhang, Lili Lin, Fan Zhong, Guanyu Xing, and Qunsheng Peng. Decomposition equation of basis images with consideration of global illumination. Lecture Notes in Computer Science, 7633: 17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_3/.

Wang:2012:IID

[505] Zhongqiang Wang and Li Zhu. Intrinsic image decomposition with local smooth assumption and global color assumption. Lecture Notes in Computer Science, 7633:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34263-9_4/.

Li:2012:GTA

[506] Jing Li, Gang Zeng, Rui Gan, Hongbin Zha, and Long Wang. A gametheoretical approach to image segmentation. Lecture Notes in Computer Science, 7633:33–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34263-9_5/.

Guo:2012:CNH

[507] Yu Guo, Xiaowu Chen, Bin Zhou, and Qinping Zhao. Clothed and naked human shapes estimation from a single image. Lecture Notes in Computer Science, 7633:43–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34263-9_6/.

Wang:2012:ICA

[508] Xiaohui Wang, Jia Jia, Hanyu Liao, and Lianhong Cai. Image colorization with an affective word. Lecture Notes in Computer Science, 7633: 51–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_7/.

Chen:2012:SIC

[509] Na Chen and Viktor K. Prasanna. Semantic image clustering using object relation network. Lecture Notes in Computer Science, 7633: 59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_8/.

Zhang:2012:EST

[510] Guo-Xin Zhang, Yu-Kun Lai, and Shi-Min Hu. Efficient solid texture synthesis using gradient solids. Lecture Notes in Computer Science, 7633: 67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_9/.

Fan:2012:RAD

[511] Hanqi Fan, Qunsheng Peng, and Yizhou Yu. A robust algorithm for denoising meshes with high-resolution details. Lecture Notes in Computer Science, 7633:75–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34263-9_10/.

Zhao:2012:MSP

[512] Jieyi Zhao, Min Tang, and Ruofeng Tong. Mesh segmentation for parallel decompression on GPU. Lecture Notes in Computer Science, 7633: 83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_11/.

Zhou:2012:MRU

[513] Qian-Yi Zhou and Ulrich Neumann. Modeling residential urban areas from dense aerial LiDAR point clouds. Lecture Notes in Computer Science, 7633:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_12/.

Weng:2012:CTM

[514] Yanlin Weng, Dongping Li, and Yiying Tong. Constrained texture mapping on subdivision surfaces. Lecture Notes in Computer Science, 7633: 99–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_13/.

Fan:2012:SRC

[515] Qiang Fan and Chun Qi. Similar region contrast based salient object detection. Lecture Notes in Computer Science, 7633:107–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34263-9_14/.

Miao:2012:SET

[516] Yongwei Miao, Jieqing Feng, Jinrong Wang, and Renato Pajarola. A shape enhancement technique based on multi-channel salience measure. Lecture Notes in Computer Science, 7633:115–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34263-9_15/.

Yang:2012:MSS

[517] Yong-Liang Yang and Chao-Hui Shen. Multi-scale salient feature extraction on mesh models. Lecture Notes in Computer Science, 7633: 122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_16/.

Wang:2012:GCS

[518] Jie Wang, Caiming Zhang, Yuanfeng Zhou, Yu Wei, and Yi Liu. Global contrast of superpixels based salient region detection. Lecture Notes in Computer Science, 7633:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34263-9_17/.

Zhang:2012:ISS

[519] Lei Zhang, Yao Zhao, and Zhenfeng Zhu. Incremental shared subspace learning for multi-label classification. Lecture Notes in Computer Science, 7633:138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34263-9_18/.

Liu:2012:LDB

[520] Yong-Jin Liu, Qiu-Fang Fu, Ye Liu, and Xiao-Lan Fu. 2D-line-drawing-based 3D object recognition. Lecture Notes in Computer Science, 7633: 146–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_19/.

Xiao:2012:GRI

[521] Yanhui Xiao, Zhenfeng Zhu, and Yao Zhao. Graph regularized ICA for over-complete feature learning. Lecture Notes in Computer Science, 7633: 154–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34263-9_20/.

Anonymous:2012:FMbc

[522] Anonymous. Front matter. Lecture Notes in Computer Science, 7633: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34263-9/1.

Feng:2012:USD

[523] XiaoXiao Feng, Koichi Matsumoto, and Shigeo Sugimoto. Uncovering the secrets of Daoism Fus using digital Dao-Fa Hui-Yuan. Lecture Notes in Computer Science, 7634:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34752-8_1/.

Baumann:2012:APC

[524] Ryan Baumann, Christopher Blackwell, and W. Brent Seales. Automatic perspective correction of manuscript images. Lecture Notes in Computer Science, 7634:11–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34752-8_2/.

Wang:2012:OBF

[525] Chao Wang, Ling Zhuang, Jiangqin Wu, and Feng Zhou. An ontologybased fuzzy matching approach to semantic retrieval of historical place names. Lecture Notes in Computer Science, 7634:19–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34752-8_3/.

Usman:2012:ESS

[526] Amjad Usman and Sharifullah Khan. Exploiting semantics in subjectbased searching of DSpace. Lecture Notes in Computer Science, 7634: 29–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34752-8_4/.

Fujikawa:2012:CCA

[527] Kazuya Fujikawa, Hideo Joho, and Shin’ichi Nakayama. Constraint can affect human perception, behaviour, and performance of search. Lecture Notes in Computer Science, 7634:39–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34752-8_5/.

Phiri:2012:BGD

[528] Lighton Phiri, Kyle Williams, Miles Robinson, Stuart Hammar, and Hussein Suleman. Bonolo: a general digital library system for file-based collections. Lecture Notes in Computer Science, 7634:49–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34752-8_6/.

Ishita:2012:RCP

[529] Emi Ishita, Yukiko Watanabe, Ai Yamaguchi, Takako Oda, Yuiko Higa, and Miki Horiuchi. The ratio of conference papers in citations of engineering dissertations at Kyushu University. Lecture Notes in Computer Science, 7634:59–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34752-8_7/.

Mercun:2012:PEC

[530] Tanja Merčun, Maja Žumer, and Trond Aalberg. Presenting and exploring the complexity of bibliographic relationships. Lecture Notes in Computer Science, 7634:63–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34752-8_8/.

Guruprasad:2012:UPA

[531] R. Guruprasad, Khaiser Nikam, and P. Marimuthu. Use patterns of aerospace E-indexing, abstracting and citation services: a research study of aerospace scientists and engineers of Bangalore. Lecture Notes in Computer Science, 7634:67–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34752-8_9/.

Yang:2012:TAD

[532] Christopher C. Yang, Xuning Tang, Min Song, and Suyeon Kim. A trend analysis of domain-specific literatures with content and co-author network similarity. Lecture Notes in Computer Science, 7634:73–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34752-8_10/.

Shams:2012:ISK

[533] Rushdi Shams and Robert E. Mercer. Improving supervised keyphrase indexer classification of keyphrases with text denoising. Lecture Notes in Computer Science, 7634:77–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34752-8_11/.

Mihara:2012:MCA

[534] Tetsuya Mihara, Mitsuharu Nagamori, and Shigeo Sugimoto. A metadatacentric approach to a production and browsing platform of Manga. Lecture Notes in Computer Science, 7634:87–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34752-8_12/.

Cai:2012:CMS

[535] Yundong Cai, Yin-Leng Theng, Qimeng Cai, Zhi Ling, Yangbo Ou, and Gladys Theng. Crowdsourcing metadata schema generation for Chinesestyle costume digital library. Lecture Notes in Computer Science, 7634: 97–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34752-8_13/.

Tan:2012:SIP

[536] Esther Meng-Yoke Tan, Dion Hoe-Lian Goh, Yin-Leng Theng, and Schubert Foo. Social interaction patterns during mobile information seeking. Lecture Notes in Computer Science, 7634:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34752-8_14/.

Zheng:2012:SAS

[537] Liangju Zheng, Weiming Lu, Meng Zhang, Baofeng Zhang, and Baogang Wei. A scalable, adaptive, self-management and fault-tolerant architecture for digital library. Lecture Notes in Computer Science, 7634: 116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34752-8_15/.

Pe-Than:2012:EMI

[538] Ei Pa Pa Pe-Than, Dion Hoe-Lian Goh, and Chei Sian Lee. Enjoyment of a mobile information sharing game: Perspectives from needs satisfaction and information quality. Lecture Notes in Computer Science, 7634: 126–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34752-8_16/.

Kanyacome:2012:NCD

[539] Suparp Kanyacome, Smarn Loipha, and Somchai Numprasertchai. Needs of collaborative digital library for secondary school students in Thailand. Lecture Notes in Computer Science, 7634:136–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34752-8_17/.

Nel:2012:ELI

[540] Marius Nel, Kyle Williams, and Hussein Suleman. Evaluating large image support for DSpace. Lecture Notes in Computer Science, 7634: 140–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34752-8_18/.

Pu:2012:UNO

[541] Hsiao-Tieh Pu and Ying-Han Chang. User navigation for online video leisure: An exploratory study. Lecture Notes in Computer Science, 7634: 144–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34752-8_19/.

Zumer:2012:UVF

[542] Maja Žumer, Athena Salaba, and Yin Zhang. User verification of the FRBR conceptual model and testing of FRBR prototypes. Lecture Notes in Computer Science, 7634:148–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34752-8_20/.

Anonymous:2012:FMbd

[543] Anonymous. Front matter. Lecture Notes in Computer Science, 7634: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34752-8/1.

Tokoro:2012:TPA

[544] Mario Tokoro. Toward practical application of formal methods in software lifecycle processes. Lecture Notes in Computer Science, 7635:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34281-3_1.

Cofer:2012:FMA

[545] Darren Cofer. Formal methods in the aerospace industry: Follow the money. Lecture Notes in Computer Science, 7635:2–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34281-3_2.

Shostak:2012:ATR

[546] Robert E. Shostak. Applying term rewriting to speech recognition of numbers. Lecture Notes in Computer Science, 7635:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-34281-3_3.

Le:2012:VPC

[547] Duy-Khanh Le, Wei-Ngan Chin, and Yong-Meng Teo. Variable permissions for concurrency verification. Lecture Notes in Computer Science, 7635:5–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34281-3_4/.

Yang:2012:CTP

[548] Xiaoxiao Yang, Yu Zhang, Ming Fu, and Xinyu Feng. A concurrent temporal programming model with atomic blocks. Lecture Notes in Computer Science, 7635:22–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34281-3_5/.

Barnett:2012:CMM

[549] Granville Barnett and Shengchao Qin. A composable mixed mode concurrency control semantics for transactional programs. Lecture Notes in Computer Science, 7635:38–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34281-3_6/.

Gjondrekaj:2012:TFV

[550] Edmond Gjondrekaj, Michele Loreti, Rosario Pugliese, Francesco Tiezzi, and Carlo Pinciroli. Towards a formal verification methodology for collective robotic systems. Lecture Notes in Computer Science, 7635: 54–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34281-3_7/.

Johnsen:2012:MRA

[551] Einar Broch Johnsen, Rudolf Schlatte, and Silvia Lizeth Tapia Tarifa. Modeling resource-aware virtualized applications for the cloud in real-time ABS. Lecture Notes in Computer Science, 7635:71–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34281-3_8/.

Ogata:2012:SMC

[552] Kazuhiro Ogata and Phan Thi Thanh Huyen. Specification and model checking of the chandy and Lamport distributed snapshot algorithm in rewriting logic. Lecture Notes in Computer Science, 7635:87–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34281-3_9/.

Mu:2012:QPD

[553] Chunyan Mu. Quantitative program dependence graphs. Lecture Notes in Computer Science, 7635:103–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34281-3_10/.

Mhamdi:2012:QAI

[554] Tarek Mhamdi, Osman Hasan, and Sofiène Tahar. Quantitative analysis of information flow using theorem proving. Lecture Notes in Computer Science, 7635:119–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34281-3_11/.

Varshosaz:2012:MVP

[555] Mahsa Varshosaz and Ramtin Khosravi. Modeling and verification of probabilistic actor systems using pRebeca. Lecture Notes in Computer Science, 7635:135–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34281-3_12/.

Zongyan:2012:MVO

[556] Qiu Zongyan, Hong Ali, and Liu Yijing. Modular verification of OO programs with interfaces. Lecture Notes in Computer Science, 7635: 151–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34281-3_13/.

Bobot:2012:SPT

[557] François Bobot and Jean-Christophe Filliâtre. Separation predicates: a taste of separation logic in first-order logic. Lecture Notes in Computer Science, 7635:167–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34281-3_14/.

Harrison:2012:CPP

[558] William L. Harrison, Adam Procter, and Gerard Allwein. The confinement problem in the presence of faults. Lecture Notes in Computer Science, 7635:182–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34281-3_15/.

Buttner:2012:VAT

[559] Fabian Büttner, Marina Egea, Jordi Cabot, and Martin Gogolla. Verification of ATL transformations using transformation models and model finders. Lecture Notes in Computer Science, 7635:198–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34281-3_16/.

Lin:2012:AGP

[560] Shang-Wei Lin, Yang Liu, Pao-Ann Hsiung, Jun Sun, and Jin Song Dong. Automatic generation of provably correct embedded systems. Lecture Notes in Computer Science, 7635:214–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34281-3_17/.

Su:2012:CMD

[561] Wen Su, Jean-Raymond Abrial, and Huibiao Zhu. Complementary methodologies for developing hybrid systems with event-B. Lecture Notes in Computer Science, 7635:230–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34281-3_18/.

Tomita:2012:TLM

[562] Takashi Tomita, Shin Hiura, Shigeki Hagihara, and Naoki Yonezaki. A temporal logic with mean-payoff constraints. Lecture Notes in Computer Science, 7635:249–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34281-3_19/.

Han:2012:TCT

[563] Meng Han, Zhenhua Duan, and Xiaobing Wang. Time constraints with temporal logic programming. Lecture Notes in Computer Science, 7635: 266–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34281-3_20/.

Anonymous:2012:FMbe

[564] Anonymous. Front matter. Lecture Notes in Computer Science, 7635: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34281-3/1.

Moha:2012:SDS

[565] Naouel Moha, Francis Palma, Mathieu Nayrolles, and Benjamin Joyen Conseil. Specification and detection of SOA antipatterns. Lecture Notes in Computer Science, 7636:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34321-6_1/.

Belardinelli:2012:VGB

[566] Francesco Belardinelli, Alessio Lomuscio, and Fabio Patrizi. Verification of GSM-based artifact-centric systems through finite abstraction. Lecture Notes in Computer Science, 7636:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34321-6_2/.

Americo:2012:SCA

[567] João Claudio Américo and Didier Donsez. Service component architecture extensions for dynamic systems. Lecture Notes in Computer Science, 7636:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34321-6_3/.

Sengupta:2012:WDY

[568] Bikram Sengupta, Anshu Jain, Kamal Bhattacharya, Hong-Linh Truong, and Schahram Dustdar. Who do you call? Problem resolution through social compute units. Lecture Notes in Computer Science, 7636:48–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34321-6_4/.

Kurniawan:2012:RPC

[569] Tri A. Kurniawan, Aditya K. Ghose, Hoa Khanh Dam, and Lam-Son Lê. Relationship-preserving change propagation in process ecosystems. Lecture Notes in Computer Science, 7636:63–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34321-6_5/.

Gupta:2012:SST

[570] Hari S. Gupta and Bikram Sengupta. Scheduling service tickets in shared delivery. Lecture Notes in Computer Science, 7636:79–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34321-6_6/.

Truong:2012:PHS

[571] Hong-Linh Truong, Schahram Dustdar, and Kamal Bhattacharya. Programming hybrid services in the cloud. Lecture Notes in Computer Science, 7636:96–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_7/.

Ye:2012:QAC

[572] Zhen Ye, Athman Bouguettaya, and Xiaofang Zhou. QoS-aware cloud service composition based on economic models. Lecture Notes in Computer Science, 7636:111–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_8/.

Wittern:2012:CSS

[573] Erik Wittern, Jörn Kuhlenkamp, and Michael Menzel. Cloud service selection based on variability modeling. Lecture Notes in Computer Science, 7636:127–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_9/.

Roy:2012:EES

[574] Marcus Roy, Ingo Weber, and Boualem Benatallah. Extending enterprise service design knowledge using clustering. Lecture Notes in Computer Science, 7636:142–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_10/.

Wittern:2012:PSD

[575] Erik Wittern, Nelly Schuster, Jörn Kuhlenkamp, and Stefan Tai. Participatory service design through composed and coordinated service feature models. Lecture Notes in Computer Science, 7636:158–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34321-6_11/.

Yu:2012:PAE

[576] Jian Yu, Jun Han, Quan Z. Sheng, and Steven O. Gunarso. Per-CAS: An approach to enabling dynamic and personalized adaptation for context-aware services. Lecture Notes in Computer Science, 7636: 173–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34321-6_12/.

Shrinivasan:2012:MAI

[577] Yedendra Babu Shrinivasan, Gargi Banerjee Dasgupta, Nirmit Desai, and Jayan Nallacherry. A method for assessing influence relationships among KPIs of service systems. Lecture Notes in Computer Science, 7636: 191–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34321-6_13/.

Patikirikorala:2012:DPM

[578] Tharindu Patikirikorala, Indika Kumara, Alan Colman, Jun Han, and Liuping Wang. Dynamic performance management in multi-tenanted business process servers using nonlinear control. Lecture Notes in Computer Science, 7636:206–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_14/.

Baouab:2012:ODE

[579] Aymen Baouab, Olivier Perrin, and Claude Godart. An optimized derivation of event queries to monitor choreography violations. Lecture Notes in Computer Science, 7636:222–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34321-6_15/.

Ramacher:2012:DSS

[580] Rene Ramacher and Lars Mönch. Dynamic service selection with endto-end constrained uncertain QoS attributes. Lecture Notes in Computer Science, 7636:237–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_16/.

Ivanovic:2012:CBA

[581] Dragan Ivanović, Manuel Carro, and Manuel V. Hermenegildo. A constraint-based approach to quality assurance in service choreographies. Lecture Notes in Computer Science, 7636:252–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34321-6_17/.

Lu:2012:SOR

[582] Kuan Lu, Ramin Yahyapour, Edwin Yaqub, and Constantinos Kotsokalis. Structural optimization of reduced ordered binary decision diagrams for SLA negotiation in IaaS of cloud computing. Lecture Notes in Computer Science, 7636:268–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34321-6_18/.

Oster:2012:SCF

[583] Zachary J. Oster, Syed Adeel Ali, Ganesh Ram Santhanam, Samik Basu, and Partha S. Roop. A service composition framework based on goal-oriented requirements engineering, model checking, and qualitative preference analysis. Lecture Notes in Computer Science, 7636: 283–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34321-6_19/.

Wang:2012:WNW

[584] Hongbing Wang, Jie Zhang, Wenlong Sun, Hongye Song, Guibing Guo, and Xiang Zhou. WCP-Nets: a weighted extension to CP-Nets for Web service selection. Lecture Notes in Computer Science, 7636:298– 312, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34321-6_20/.

Anonymous:2012:FMbf

[585] Anonymous. Front matter. Lecture Notes in Computer Science, 7636: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34321-6/1.

Oliverio:2012:CDO

[586] Vinicius Oliverio and Estevam R. Hruschka, Jr. Contradiction detection and ontology extension in a never-ending learning system. Lecture Notes in Computer Science, 7637:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34654-5_1/.

Yee:2012:PRM

[587] Arturo Yee and Matías Alvarado. Pattern recognition and Monte-Carlo Tree Search for Go gaming better automation. Lecture Notes in Computer Science, 7637:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34654-5_2/.

Chacon:2012:ERS

[588] José Luis Chacón and Ramón Pino Pérez. Exploring the rationality of some syntactic merging operators. Lecture Notes in Computer Science, 7637:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_3/.

Grau:2012:FCM

[589] Isel Grau, Gonzalo Nápoles, Maikel León, and Ricardo Grau. Fuzzy cognitive maps for modelling, predicting and interpreting HIV drug resistance. Lecture Notes in Computer Science, 7637:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34654-5_4/.

Tomaz:2012:UMO

[590] Hilário Tomaz, Rinaldo Lima, João Emanoel, and Fred Freitas. An unsupervised method for ontology population from the Web. Lecture Notes in Computer Science, 7637:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34654-5_5/.

Gomez:2012:MCS

[591] Jonatan Gomez, Elizabeth León, and Olfa Nasraoui. Minimum cluster size estimation and cluster refinement for the randomized gravitational clustering algorithm. Lecture Notes in Computer Science, 7637: 51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_6/.

Leon:2012:OCP

[592] Elizabeth León, Jonatan Gómez, and Fabián Giraldo. Online cluster prototype generation for the gravitational clustering algorithm. Lecture Notes in Computer Science, 7637:61–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34654-5_7/.

Fernandes:2012:ARV

[593] Leandro A. F. Fernandes and Ana Cristina Bicharra García. Association rule visualization and pruning through response-style data organization and clustering. Lecture Notes in Computer Science, 7637: 71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_8/.

Rodrigues:2012:IRT

[594] Melissa Rodrigues, João Gama, and Carlos Abreu Ferreira. Identifying relationships in transactional data. Lecture Notes in Computer Science, 7637:81–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_9/.

Zalewski:2012:TSD

[595] Willian Zalewski, Fabiano Silva, Huei Diana Lee, Andre Gustavo Maletzke, and Feng Chung Wu. Time series discretization based on the approximation of the local slope information. Lecture Notes in Computer Science, 7637:91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34654-5_10/.

Santiago:2012:EPS

[596] José Hernández Santiago, Jair Cervantes, Asdrúbal López-Chau, and Farid García Lamont. Enhancing the performance of SVM on skewed data sets by exciting support vectors. Lecture Notes in Computer Science, 7637:101–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34654-5_11/.

Skrobanek:2012:SIA

[597] Paweł Skrobanek, Olgierd Unold, Ewa Walkowicz, Henryk Maciejewski, and Maciej Dobrowolski. Study on the impact of affinity on the results of data mining in biological populations. Lecture Notes in Computer Science, 7637:111–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34654-5_12/.

Romao:2012:HCG

[598] Luiz Melo Romão and Julio César Nievola. Hierarchical classification of gene ontology with learning classifier systems. Lecture Notes in Computer Science, 7637:120–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34654-5_13/.

Pereira:2012:DSP

[599] Ricardo Timarán Pereira, Maria Clara Yepez Chamorro, and Andrés Calderón Romero. Detecting survival patterns in women with invasive cervical cancer with decision trees. Lecture Notes in Computer Science, 7637: 130–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_14/.

Pasa:2012:USM

[600] Leandro A. Pasa, José Alfredo F. Costa, Marcelo C. Tosin, and Fábio A. Procópio de Paiva. Using SOM maps for clustering and visualization of diamond films deposited by HFCVD process. Lecture Notes in Computer Science, 7637:140–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34654-5_15/.

deLima:2012:SFN

[601] Tiago P. F. de Lima, Adenilton J. da Silva, and Teresa B. Ludermir. Selection and fusion of neural networks via differential evolution. Lecture Notes in Computer Science, 7637:149–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34654-5_16/.

Villuendas-Rey:2012:PSC

[602] Yenny Villuendas-Rey, Yailé Caballero-Mota, and María Matilde García-Lorenzo. Prototype selection with compact sets and extended rough sets. Lecture Notes in Computer Science, 7637:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34654-5_17/.

Verbiest:2012:ISF

[603] Nele Verbiest, Enislay Ramentol, Chris Cornelis, and Francisco Herrera. Improving SMOTE with fuzzy rough prototype selection to detect noise in imbalanced classification data. Lecture Notes in Computer Science, 7637: 169–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_18/.

Cobos:2012:FFO

[604] Carlos Cobos, Leydy Muñoz, Martha Mendoza, Elizabeth León, and Enrique Herrera-Viedma. Fitness function obtained from a genetic programming approach for Web document clustering using evolutionary algorithms. Lecture Notes in Computer Science, 7637:179–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34654-5_19/.

Metz:2012:EPE

[605] Jean Metz, Luís F. D. de Abreu, Everton A. Cherman, and Maria C. Monard. On the estimation of predictive evaluation measure baselines for multi-label learning. Lecture Notes in Computer Science, 7637:

189–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34654-5_20/.

Anonymous:2012:FMbg

[606] Anonymous. Front matter. Lecture Notes in Computer Science, 7637: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34654-5/1.

Johnson:2012:WHB

[607] Benjamin Johnson, Pascal Schöttle, and Rainer Böhme. Where to hide the bits? Lecture Notes in Computer Science, 7638:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34266-0_1/.

Nojoumian:2012:SRS

[608] Mehrdad Nojoumian and Douglas R. Stinson. Socio-rational secret sharing as a new direction in rational cryptography. Lecture Notes in Computer Science, 7638:18–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34266-0_2/.

Blocki:2012:AMP

[609] Jeremiah Blocki, Nicolas Christin, Anupam Datta, and Arunesh Sinha. Audit mechanisms for provable risk management and accountable data governance. Lecture Notes in Computer Science, 7638:38–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34266-0_3/.

Hoe:2012:GTA

[610] SingRu (Celine) Hoe, Murat Kantarcioglu, and Alain Bensoussan. A game theoretical analysis of lemonizing cybercriminal black markets. Lecture Notes in Computer Science, 7638:60–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34266-0_4/.

Dritsoula:2012:CNE

[611] Lemonia Dritsoula, Patrick Loiseau, and John Musacchio. Computing the Nash equilibria of intruder classification games. Lecture Notes in Computer Science, 7638:78–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34266-0_5/.

Buldas:2012:UBA

[612] Ahto Buldas and Roman Stepanenko. Upper bounds for adversaries’ utility in attack trees. Lecture Notes in Computer Science, 7638:98– 117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34266-0_6/.

Lin:2012:USG

[613] Jingqiang Lin, Peng Liu, and Jiwu Jing. Using signaling games to model the multi-step attack-defense scenarios on confidentiality. Lecture Notes in Computer Science, 7638:118–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34266-0_7/.

Nochenson:2012:SGT

[614] Alan Nochenson and C. F. Larry Heimann. Simulation and game-theoretic analysis of an attacker-defender game. Lecture Notes in Computer Science, 7638:138–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34266-0_8/.

Laszka:2012:LLF

[615] Aron Laszka, Dávid Szeszlér, and Levente Buttyán. Linear loss function for the network blocking game: An efficient model for measuring network robustness and link criticality. Lecture Notes in Computer Science, 7638: 152–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34266-0_9/.

Clark:2012:DRR

[616] Andrew Clark, Quanyan Zhu, Radha Poovendran, and Tamer Başar. Deceptive routing in relay networks. Lecture Notes in Computer Science, 7638:171–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34266-0_10/.

Gueye:2012:GTF

[617] Assane Gueye and Vladimir Marbukh. A game-theoretic framework for network security vulnerability assessment and mitigation. Lecture Notes in Computer Science, 7638:186–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34266-0_11/.

Pibil:2012:GTM

[618] Radek Píbil, Viliam Lisý, Christopher Kiekintveld, Branislav Bošanský, and Michal Pěchouček. Game theoretic model of strategic honeypot selection in computer networks. Lecture Notes in Computer Science, 7638: 201–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34266-0_12/.

Collins:2012:CBM

[619] M. Patrick Collins. A cost-based mechanism for evaluating the effectiveness of moving target defenses. Lecture Notes in Computer Science, 7638: 221–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34266-0_13/.

Pham:2012:WCM

[620] Viet Pham and Carlos Cid. Are we compromised? Modelling security assessment games. Lecture Notes in Computer Science, 7638: 234–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34266-0_14/.

Bowers:2012:DAU

[621] Kevin D. Bowers, Marten van Dijk, Robert Griffin, Ari Juels, and Alina Oprea. Defending against the unknown enemy: Applying FlipIt to system security. Lecture Notes in Computer Science, 7638:248–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34266-0_15/.

Amin:2012:ISE

[622] Saurabh Amin, Galina A. Schwartz, and Hamidou Tembine. Incentives and security in electricity distribution networks. Lecture Notes in Computer Science, 7638:264–280, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34266-0_16/.

Law:2012:SGR

[623] Yee Wei Law, Tansu Alpcan, Marimuthu Palaniswami, and Subhrakanti Dey. Security games and risk minimization for automatic generation control in smart grid. Lecture Notes in Computer Science, 7638:281– 295, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34266-0_17/.

Nix:2012:CAD

[624] Robert Nix and Murat Kantarcioglu. Contractual agreement design for enforcing honesty in cloud outsourcing. Lecture Notes in Computer Science, 7638:296–308, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34266-0_18/.

Anonymous:2012:BMg

[625] Anonymous. Back matter. Lecture Notes in Computer Science, 7638: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34266-0/1.

Anonymous:2012:FMbh

[626] Anonymous. Front matter. Lecture Notes in Computer Science, 7638: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34266-0/1.

Sartor:2012:CNS

[627] Giovanni Sartor. Compliance with normative systems. Lecture Notes in Computer Science, 7639:1–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35731-2_1/.

Araszkiewicz:2012:CBA

[628] Michał Araszkiewicz. Coherence-based account of the doctrine of consistent interpretation. Lecture Notes in Computer Science, 7639:33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35731-2_2/.

Pagallo:2012:TRC

[629] Ugo Pagallo. Three roads to complexity, AI and the law of robots: On crimes, contracts, and torts. Lecture Notes in Computer Science, 7639: 48–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35731-2_3/.

Weng:2012:LCN

[630] Yueh-Hsuan Weng and Sophie Ting Hong Zhao. The legal challenges of networked robotics: From the safety intelligence perspective. Lecture Notes in Computer Science, 7639:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35731-2_4/.

Bourcier:2012:CCN

[631] Daniele Bourcier and Primavera De Filippi. Cloud computting: New research perspectives for computers and law. Lecture Notes in Computer Science, 7639:73–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35731-2_5/.

Agnoloni:2012:BRV

[632] Tommaso Agnoloni, Maria-Teresa Sagri, and Daniela Tiscornia. Balancing rights and values in the Italian courts: a benchmark for a quantitative analysis. Lecture Notes in Computer Science, 7639:93–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35731-2_6/.

Winkels:2012:SFN

[633] Radboud Winkels and Jelle de Ruyter. Survival of the fittest: Network analysis of Dutch Supreme Court cases. Lecture Notes in Computer Science, 7639:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35731-2_7/.

Ceci:2012:OFJ

[634] Marcello Ceci and Monica Palmirani. Ontology framework for judgment modelling. Lecture Notes in Computer Science, 7639:116–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35731-2_8/.

Boella:2012:ELD

[635] Guido Boella, Llio Humphreys, Marco Martin, Piercarlo Rossi, and Leendert van der Torre. Eunomos, a legal document and knowledge management system to build legal services. Lecture Notes in Computer Science, 7639:131–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35731-2_9/.

Francesconi:2012:ASM

[636] Enrico Francesconi. Axioms on a semantic model for legislation for accessing and reasoning over normative provisions. Lecture Notes in Computer Science, 7639:147–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35731-2_10/.

Francesconi:2012:OAP

[637] Enrico Francesconi and Ginevra Peruginelli. An open access policy for legal informatics dissemination and sharing. Lecture Notes in Computer Science, 7639:162–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35731-2_11/.

Casanovas:2012:AOA

[638] Pompeu Casanovas and Enric Plaza. Advancing an open access publication model for legal information institutes. Lecture Notes in Computer Science, 7639:171–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35731-2_12/.

Smith:2012:CNN

[639] Clara Smith, Agustín Ambrossio, Leandro Mendoza, and Antonino Rotolo. Combinations of normal and non-normal modal logics for modeling collective trust in normative MAS. Lecture Notes in Computer Science, 7639:189–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35731-2_13/.

Laukyte:2012:SAB

[640] Migle Laukyte. Software agents as boundary objects. Lecture Notes in Computer Science, 7639:204–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35731-2_14/.

Vincent:2012:AID

[641] Andrew Vincent. Argumentation and intuitive decision making: Criminal sentencing and sentence indication. Lecture Notes in Computer Science, 7639:217–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35731-2_15/.

Boer:2012:AMB

[642] Alexander Boer and Tom van Engers. Application of model-based diagnosis to multi-agent systems representing public administration. Lecture Notes in Computer Science, 7639:235–244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35731-2_16/.

Ceci:2012:SAL

[643] Marcello Ceci, Leonardo Lesmo, Alessandro Mazzei, and Monica Palmirani. Semantic annotation of legal texts through a FrameNet-based approach. Lecture Notes in Computer Science, 7639:245–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35731-2_17/.

Bertoldi:2012:DFB

[644] Anderson Bertoldi and Rove Luiza de Oliveira Chishman. Developing a frame-based lexicon for the Brazilian legal language: The case of the Criminal Process Frame. Lecture Notes in Computer Science, 7639: 256–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35731-2_18/.

Myska:2012:CCG

[645] Matěj Myška, Terezie Smejkalová, Jaromír Šavelka, and Martin Škop. Creative Commons and grand challenge to make legal language simple. Lecture Notes in Computer Science, 7639:271–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35731-2_19/.

Fernandez-Barrera:2012:UNE

[646] Meritxell Fernández-Barrera and Pompeu Casanovas. From user needs to expert knowledge: Mapping laymen queries with ontologies in the domain of consumer mediation. Lecture Notes in Computer Science, 7639: 286–308, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35731-2_20/.

Anonymous:2012:BMh

[647] Anonymous. Back matter. Lecture Notes in Computer Science, 7639: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35731-2/1.

Anonymous:2012:FMbi

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35731-2/1.

Meng:2012:ELB

[649] Yuxin Meng and Lam for Kwok. Enhancing list-based packet filter using IP verification mechanism against IP spoofing attack in network intrusion detection. Lecture Notes in Computer Science, 7645:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34601-9_1/.

Ranise:2012:AAS

[650] Silvio Ranise and Alessandro Armando. On the automated analysis of safety in usage control: a new decidability result. Lecture Notes in Computer Science, 7645:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34601-9_2/.

Wagner:2012:AMB

[651] Steffen Wagner, Sascha Wessel, and Frederic Stumpf. Attestation of mobile baseband stacks. Lecture Notes in Computer Science, 7645: 29–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34601-9_3/.

Lee:2012:SLM

[652] Tae-Gyu Lee and Gi-Soo Chung. A scalable link model for local optical wireless networks. Lecture Notes in Computer Science, 7645:44– 57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34601-9_4/.

Alcaraz:2012:ASA

[653] Cristina Alcaraz and Javier Lopez. Addressing situational awareness in critical domains of a smart grid. Lecture Notes in Computer Science, 7645:58–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34601-9_5/.

Ibrahim:2012:IKO

[654] Amani S. Ibrahim, James Hamlyn-Harris, John Grundy, and Mohamed Almorsy. Identifying OS kernel objects for run-time security analysis. Lecture Notes in Computer Science, 7645:72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34601-9_6/.

Lee:2012:BTM

[655] Tae-Gyu Lee and Gi-Soo Chung. Background transfer method for ubiquitous computing. Lecture Notes in Computer Science, 7645:86–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34601-9_7/.

Liu:2012:SOCa

[656] Shengli Liu, Fangguo Zhang, and Kefei Chen. Selective opening chosen ciphertext security directly from the DDH assumption. Lecture Notes in Computer Science, 7645:100–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34601-9_8/.

Tang:2012:PSS

[657] Shaohua Tang and Lingling Xu. Proxy signature scheme based on isomorphisms of polynomials. Lecture Notes in Computer Science, 7645: 113–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34601-9_9/.

Tang:2012:UDV

[658] Fei Tang, Changlu Lin, and Pinhui Ke. Universal designated verifier signcryption. Lecture Notes in Computer Science, 7645:126–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34601-9_10/.

Timpanaro:2012:BEV

[659] Juan Pablo Timpanaro, Isabelle Chrisment, and Olivier Festor. A Bird’s eye view on the I2P anonymous file-sharing environment. Lecture Notes in Computer Science, 7645:135–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34601-9_11/.

Mahdavifar:2012:CBA

[660] Samaneh Mahdavifar, Mahdi Abadi, Mohsen Kahani, and Hassan Mahdikhani. A clustering-based approach for personalized privacy preserving publication of moving object trajectory data. Lecture Notes in Computer Science, 7645:149–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34601-9_12/.

Wahid:2012:ENH

[661] Alif Wahid, Christopher Leckie, and Chenfeng Zhou. Estimating the number of hosts corresponding to an address while preserving anonymity. Lecture Notes in Computer Science, 7645:166–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34601-9_13/.

Han:2012:ERI

[662] Qi Han, Yinghui Zhang, Xiaofeng Chen, Hui Li, and Jiaxiang Quan. Efficient and robust identity-based handoff authentication in wireless networks. Lecture Notes in Computer Science, 7645:180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34601-9_14/.

Zhao:2012:IAS

[663] Yifan Zhao, Swee-Won Lo, Robert H. Deng, and Xuhua Ding. An improved authentication scheme for H.264/SVC and its performance evaluation over non-stationary wireless mobile networks. Lecture Notes in Computer Science, 7645:192–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34601-9_15/.

Xiong:2012:PPK

[664] Kaiqi Xiong. The performance of public key-based authentication protocols. Lecture Notes in Computer Science, 7645:206–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34601-9_16/.

Li:2012:BVS

[665] Long-Hai Li, Cheng-Qiang Huang, and Shao-Feng Fu. Boardroom voting scheme with unconditionally secret ballots based on DC-Net. Lecture Notes in Computer Science, 7645:220–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34601-9_17/.

Yu:2012:RSN

[666] Yue Yu, Michael Fry, Bernhard Plattner, Paul Smith, and Alberto Schaeffer-Filho. Resilience strategies for networked malware detection and remediation. Lecture Notes in Computer Science, 7645:233–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34601-9_18/.

Menahem:2012:DSA

[667] Eitan Menahem, Rami Pusiz, and Yuval Elovici. Detecting spammers via aggregated historical data set. Lecture Notes in Computer Science, 7645: 248–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34601-9_19/.

Ibrahim:2012:OSK

[668] Amani S. Ibrahim, John Grundy, James Hamlyn-Harris, and Mohamed Almorsy. Operating system kernel data disambiguation to support security analysis. Lecture Notes in Computer Science, 7645:263–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34601-9_20/.

Anonymous:2012:FMbj

[669] Anonymous. Front matter. Lecture Notes in Computer Science, 7645: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34601-9/1.

Hameurlain:2012:TLS

[670] Abdelkader Hameurlain, Josef Küng, Roland Wagner, Stephen W. Liddle, Klaus-Dieter Schewe, and Xiaofang Zhou, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems VI: Special Issue on Database- and Expert-Systems Applications, volume 7600 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34178-0 (print), 3-642-34179-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-34179-3.

Yoshida:2012:AIC

[671] Hiroyuki Yoshida, David Hawkes, and Michael W. Vannier, editors. Abdominal Imaging. Computational and Clinical Applications: 4th International Workshop, Held in Conjunction with MICCAI 2012, Nice, France, October 1, 2012. Proceedings, volume 7601 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33611-6 (print), 3-642-33612-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33612-6.

Abello:2012:MDE

[672] Alberto Abelló, Ladjel Bellatreche, and Boualem Benatallah, editors. Model and Data Engineering: 2nd International Conference, MEDI 2012, Poitiers, France, October 3–5, 2012. Proceedings, volume 7602 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33608-6 (print), 3-642-33609-4 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-33609-6.

tenTeije:2012:KEK

[673] Annette ten Teije, Johanna Völker, Siegfried Handschuh, Heiner Stuckenschmidt, Mathieu d’Acquin, Andriy Nikolov, Nathalie Aussenac-Gilles, and Nathalie Hernandez, editors. Knowledge Engineering and Knowledge Management: 18th International Conference, EKAW 2012, Galway City, Ireland, October 8–12, 2012. Proceedings, volume 7603 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33875-5 (print), 3-642-33876-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33876-2.

Baeten:2012:TCS

[674] Jos C. M. Baeten, Tom Ball, and Frank S. de Boer, editors. Theoretical Computer Science: 7th IFIP TC 1/WG 2.2 International Conference, TCS 2012, Amsterdam, The Netherlands, September 26–28, 2012. Proceedings, volume 7604 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33474-1 (print), 3-642-33475-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33475-7.

Calderon-Benavides:2012:SPI

[675] Liliana Calderón-Benavides, Cristina González-Caro, Edgar Chávez, and Nivio Ziviani, editors. String Processing and Information Retrieval: 19th International Symposium, SPIRE 2012, Cartagena de Indias, Colombia, October 21–25, 2012. Proceedings, volume 7608 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34108-X (print), 3-642-34109-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-34109-0.

Margaria:2012:LAF

[676] Tiziana Margaria and Bernhard Steffen, editors. Leveraging Applications of Formal Methods, Verification and Validation. Applications and Case Studies: 5th International Symposium, ISoLA 2012, Heraklion, Crete, Greece, October 15–18, 2012, Proceedings, Part II, volume 7610 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34031-8 (print), 3-642-34032-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-34032-1.

Aguilera:2012:DCI

[677] Marcos K. Aguilera, editor. Distributed Computing: 26th International Symposium, DISC 2012, Salvador, Brazil, October 16–18, 2012. Proceedings, volume 7611 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33650-7 (print), 3-642-33651-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33651-5.

Ortmeier:2012:CSRa

[678] Frank Ortmeier and Peter Daniel, editors. Computer Safety, Reliability, and Security: 31st International Conference, SAFECOMP 2012, Magdeburg, Germany, September 25–28, 2012. Proceedings, volume 7612 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33677-9 (print), 3-642-33678-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33678-2.

Ortmeier:2012:CSRb

[679] Frank Ortmeier and Peter Daniel, editors. Computer Safety, Reliability, and Security: SAFECOMP 2012 Workshops: Sassur, ASCoMS, DESEC4LCCI, ERCIM/EWICS, IWDE, Magdeburg, Germany, September 25–28, 2012. Proceedings, volume 7613 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33674-4 (print), 3-642-33675-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33675-1.

Isahara:2012:ANL

[680] Hitoshi Isahara and Kyoko Kanzaki, editors. Advances in Natural Language Processing: 8th International Conference on NLP, JapTAL 2012, Kanazawa, Japan, October 22–24, 2012. Proceedings, volume 7614 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33982-4 (print), 3-642-33983-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33983-7.

Ioannides:2012:PCH

[681] Marinos Ioannides, Dieter Fritsch, Johanna Leissner, Rob Davies, Fabio Remondino, and Rossella Caffo, editors. Progress in Cultural Heritage Preservation: 4th International Conference, EuroMed 2012, Limassol, Cyprus, October 29–November 3, 2012. Proceedings, volume 7616 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34233-7 (print), 3-642-34234-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34234-9.

Josang:2012:SIS

[682] Audun Jøsang and Bengt Carlsson, editors. Secure IT Systems: 17th Nordic Conference, NordSec 2012, Karlskrona, Sweden, October 31– November 2, 2012. Proceedings, volume 7617 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34209-4 (print), 3-642-34210-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34210-3.

Chim:2012:ICS

[683] Tat Wing Chim and Tsz Hon Yuen, editors. Information and Communications Security: 14th International Conference, ICICS 2012, Hong Kong, China, October 29–31, 2012. Proceedings, volume 7618 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34128-4 (print), 3-642-34129-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-34129-8.

Hollmen:2012:AID

[684] Jaakko Hollmén, Frank Klawonn, and Allan Tucker, editors. Advances in Intelligent Data Analysis XI: 11th International Symposium, IDA 2012, Helsinki, Finland, October 25–27, 2012. Proceedings, volume 7619 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34155-1 (print), 3-642-34156-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34156-4.

Busemeyer:2012:QII

[685] Jerome R. Busemeyer, François Dubois, Ariane Lambert-Mogiliansky, and Massimo Melucci, editors. Quantum Interaction: 6th International Symposium, QI 2012, Paris, France, June 27–29, 2012, Revised Selected Papers, volume 7620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35658-3 (print), 3-642-35659-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-35659-9.

Ge:2012:SRI

[686] Shuzhi Sam Ge, Oussama Khatib, John-John Cabibihan, Reid Simmons, and Mary-Anne Williams, editors. Social Robotics: 4th International Conference, ICSR 2012, Chengdu, China, October 29–31, 2012. Proceedings, volume 7621 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34102-0 (print), 3-642-34103-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34103-8.

Christianson:2012:SPX

[687] Bruce Christianson, James Malcolm, Frank Stajano, and Jonathan Anderson, editors. Security Protocols XX: 20th International Workshop, Cambridge, UK, April 12–13, 2012, Revised Selected Papers, volume 7622 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35693-1 (print), 3-642-35694-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-35694-0.

Winckler:2012:HCS

[688] Marco Winckler, Peter Forbrig, and Regina Bernhaupt, editors. Human-Centered Software Engineering: 4th International Conference, HCSE 2012, Toulouse, France, October 29–31, 2012. Proceedings, volume 7623 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34346-5 (print), 3-642-34347-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34347-6.

Nijholt:2012:ACE

[689] Anton Nijholt, Teresa Romão, and Dennis Reidsma, editors. Advances in Computer Entertainment: 9th International Conference, ACE 2012, Kathmandu, Nepal, November 3–5, 2012. Proceedings, volume 7624 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34291-4 (print), 3-642-34292-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-34292-9.

Priami:2012:TCS

[690] Corrado Priami, Ion Petre, and Erik de Vink, editors. Transactions on Computational Systems Biology XIV: Special Issue on Computational Models for Cell Processes, volume 7625 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35523-4 (print), 3-642-35524-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-35524-0.

Gimelfarb:2012:SSS

[691] Georgy Gimel’farb, Edwin Hancock, Atsushi Imiya, Arjan Kuijper, Mineichi Kudo, Shinichiro Omachi, Terry Windeatt, and Keiji Yamada, editors. Structural, Syntactic, and Statistical Pattern Recognition: Joint IAPR International Workshop, SSPR&SPR 2012, Hiroshima, Japan, November 7– 9, 2012. Proceedings, volume 7626 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34165-9 (print), 3-642-34166-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34166-3.

Noda:2012:SMP

[692] Itsuki Noda, Noriaki Ando, Davide Brugali, and James J. Kuffner, editors. Simulation, Modeling, and Programming for Autonomous Robots: Third International Conference, SIMPAR 2012, Tsukuba, Japan, November 5– 8, 2012. Proceedings, volume 7628 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34326-0 (print), 3-642-34327-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34327-8.

Hanaoka:2012:AIC

[693] Goichiro Hanaoka and Toshihiro Yamauchi, editors. Advances in Information and Computer Security: 7th International Workshop on Security, IWSEC 2012, Fukuoka, Japan, November 7–9, 2012. Proceedings, volume 7631 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34116-0 (print), 3-642-34117-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
34117-5.

Shibuya:2012:PRB

[694] Tetsuo Shibuya, Hisashi Kashima, Jun Sese, and Shandar Ahmad, editors. Pattern Recognition in Bioinformatics: 7th IAPR International Conference, PRIB 2012, Tokyo, Japan, November 8–10, 2012. Proceedings, volume 7632 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34122-5 (print), 3-642-34123-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34123-6.

Hu:2012:CVM

[695] Shi-Min Hu and Ralph R. Martin, editors. Computational Visual Media: First International Conference, CVM 2012, Beijing, China, November 8– 10, 2012. Proceedings, volume 7633 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34262-0 (print), 3-642-34263-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34263-9.

Chen:2012:ODL

[696] Hsin-Hsi Chen and Gobinda Chowdhury, editors. The Outreach of Digital Libraries: A Globalized Resource Network: 14th International Conference on Asia-Pacific Digital Libraries, ICADL 2012, Taipei, Taiwan, November 12–15, 2012, Proceedings, volume 7634 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-34751-7 (print), 3-642-34752-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-34752-8.

Aoki:2012:FMS

[697] Toshiaki Aoki and Kenji Taguchi, editors. Formal Methods and Software Engineering: 14th International Conference on Formal Engineering Methods, ICFEM 2012, Kyoto, Japan, November 12–16, 2012. Proceedings, volume 7635 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34280-9 (print), 3-642-34281-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34281-3.

Liu:2012:SOCb

[698] Chengfei Liu, Heiko Ludwig, Farouk Toumani, and Qi Yu, editors. Service-Oriented Computing: 10th International Conference, ICSOC 2012, Shanghai, China, November 12–15, 2012. Proceedings, volume 7636 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34320-1 (print), 3-642-34321-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
34321-6.

Pavon:2012:AAI

[699] Juan Pavón, Néstor D. Duque-Méndez, and Rubén Fuentes-Fernández, editors. Advances in Artificial Intelligence — IBERAMIA 2012: 13th Ibero– American Conference on AI, Cartagena de Indias, Colombia, November 13–16, 2012. Proceedings, volume 7637 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-34653-7 (print), 3-642-34654-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-34654-5.

Grossklags:2012:DGT

[700] Jens Grossklags and Jean Walrand, editors. Decision and Game Theory for Security: Third International Conference, GameSec 2012, Budapest, Hungary, November 5–6, 2012. Proceedings, volume 7638 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34265-5 (print), 3-642-34266-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-34266-0.

Palmirani:2012:AAC

[701] Monica Palmirani, Ugo Pagallo, Pompeu Casanovas, and Giovanni Sartor, editors. AI Approaches to the Complexity of Legal Systems. Models and Ethical Challenges for Legal Systems, Legal Language and Legal Ontologies, Argumentation and Software Agents: International Workshop AICOL-III, Held as Part of the 25th IVR Congress, Frankfurt am Main, Germany, August 15–16, 2011. Revised Selected Papers, volume 7639 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35730-X (print), 3-642-35731-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-35731-2.

Xu:2012:NSS

[702] Li Xu, Elisa Bertino, and Yi Mu, editors. Network and System Security: 6th International Conference, NSS 2012, Wuyishan, Fujian, China, November 21–23, 2012. Proceedings, volume 7645 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-34600-6 (print), 3-642-34601-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-34601-9.